### v1

# Supersymmetry and Localization on Three-Dimensional Orbifolds

Abstract
We consider three-dimensional ${\mathcal{N}}=2$ supersymmetric field theories defined on general complex-valued backgrounds of Euclidean new minimal supergravity admitting two Killing spinors of opposite R-charges. We focus on compact spaces admitting orbifold singularities, including $\mathbb{\Sigma}\times S^{1}$, where $\mathbb{\Sigma}$ is a spindle, as well as branched and squashed lens spaces. In each of these supersymmetric backgrounds we compute the corresponding partition function, thus obtaining novel observables for three-dimensional supersymmetric gauge theories. In particular, our results cover a wide class of circle bundles over a spindle. For theories compactified on $\mathbb{\Sigma}\times S^{1}$ we discuss both the twist and anti-twist and show that the partition functions on these backgrounds may be expressed in terms of a single formula, which we refer to as the spindle index, unifying and generalising the refined superconformal and topologically twisted indices, in the limits wherein the orbifold singularities are removed. We test our new index by means of non-perturbative dualities. The main results of this paper were announced in [1].
1 Introduction
The investigation of supersymmetric quantum field theories (SQFTs) stands as a longstanding pursuit in theoretical physics as it represents an invaluable avenue for exploring non-trivial phenomena, particularly in the regime of strong interactions, where perturbation theory fails. From a physics standpoint, the exact computations of SQFTs’ partition functions on curved backgrounds, made possible through localization techniques [2], have produced profound insights on non-perturbative aspects of gauge theories, including the discovery of dualities among seemingly completely different models. From a mathematical point of view, such exact partition functions may serve as generating functions for a number of interesting invariants characterizing the manifolds where the theory is defined (for a recent review see e.g. [3] and references therein).
Partition functions of SQFTs defined on the direct product of a circle and a compact manifold play a special role as they correspond to supersymmetric indices, which may be interpreted in terms of “counting” of various quantities. The study of indices has revealed unforeseen correlations across different areas of physics and mathematics, establishing connections between SQFTs, the geometry of the moduli space in the context of string theory, as well as in purely mathematical fields such as symplectic and algebraic geometry and representation theory, e.g. Gromov-Witten invariants and characters of infinite-dimensional algebras. Recent advancements in localization techniques have yielded significant breakthroughs, facilitating explicit computations of supersymmetric indices as well as other partition functions corresponding to manifolds with the different topology and dimension.
Supersymmetric partition functions play also an important role in the AdS/CFT correspondence, where they provide precision checks of dualities between gravitational backgrounds and holographically dual field theories. The most striking instance of this aspect is the use of supersymmetric indices to perform a precise counting of black hole microstates, thereby offering a glimpse into the fundamental tenets of quantum gravity. In particular, through the study of large-$N$ limits, supersymmetric indices have been used to obtain a microscopic derivation of the entropy of supersymmetric black holes [4, 5, 6, 7]. Several other works have studied different asymptotic limits and determined sub-leading corrections, matching these to higher-derivative or quantum corrections in supergravity [8, 9, 10, 11]. Another set of precision tests in holography involves matching the large-$N$ limit of the partition of function of three-dimensional ${\cal N}=2$ on manifold with $S^{3}$ topology [12, 13, 14] to various supergravity solutions [15, 16, 17, 18, 19].
In the past two years, starting with [20], a substantial body of literature has demonstrated that supergravity solutions displaying orbifold singularities should not be discarded. On the contrary, such solutions are expected to play a crucial role as non-trivial saddle points in the quantum gravity path integral, or in its semi-classical limit. In this extended landscape of solutions supersymmetry can be realized in novel fashions, raising the intriguing question of how typical these supergravity backgrounds are. One of the simplest and perhaps most interesting set-ups is that four-dimensional supersymmetric [^1] See also [21] for a discussion of supersymmetry of the solutions of [22, 23]. and accelerating black holes [22, 23], in which the horizon is a spindle $\mathbb{\Sigma}$ [24], where $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$ is the two-dimensional weighted projective space whose topology is that of a two-sphere with conical singularities $\mathbb{C}/\mathbb{Z}_{n_{\pm}}$ at the poles. The conformal boundary of these geometries comprises a metric on $\mathbb{\Sigma}\times\mathbb{R}_{t}$ together with a gauge field with “anti-twisted” flux on $\mathbb{\Sigma}$; therefore, via holography, their microstates are expected to be captured by the supersymmetric partition function of the dual ${\cal N}=2$ three-dimensional field theories defined on a rigid supersymmetric background on $\mathbb{\Sigma}\times S^{1}$ [25]. In this work we will discuss in detail the calculation of this new observable, that we named spindle index in [1], where the main results of the present paper were anticipated.
More generally, the recent supergravity constructions involving orbifolds raised a number of related questions concerning superconformal field theories (SCFTs). In would be interesting to understand systematically the data that define a SCFT at an orbifold point and what are the additional properties that one can hope to detect, on such backgrounds. For SCFTs in even space-time dimensions some progress has been made by studying the ’t Hooft anomalies of the theories compactified on orbifolds, extending anomaly polynomial techniques to spaces with conical singularities [20, 26, 27]. This approach is somewhat insensitive to the detailed description of the microscopic fields at the orbifold singularities, suggesting that a similar point of view may be fruitful for studying field theories on orbifolds also in odd space-time dimensions. In particular, we expect that orbifold supersymmetric indices and other partition functions of field theories should unveil novel features of SQFTs (in odd as well as in even space-time dimension), beyond those that are immediately relevant for applications in holography.
[…]
Outline.
In Section 2 we characterise a very general class of complex local rigidly supersymmetric backgrounds. We will show that these comprise all the known supersymmetric backgrounds as well as several generalizations, that we discuss in detail. In particular, we describe general backgrounds on $\mathbb{\Sigma}\times S^{1}$, with either twist or anti-twist for the background R-symmetry gauge field. The conformal boundary of the accelerating black hole of [24] as well as its complex deformation [25] are shown to fit into our general canonical form. In Section 3 we compute the full supersymmetric partition function of three-dimensional ${\cal N}=2$ Chern-Simons-matter theories on $\mathbb{\Sigma}\times S^{1}$. The one-loop determinants are obtained employing both the method of unpaired eigenvalues as well as the equivariant orbifold index theorem. We also compute various classical terms and discuss subtleties related to the breaking of gauge invariance and its restoration through the introduction of effecting terms. Our final result for the spindle index is tested with some simple field theory dualities, that hold in a form modified by the presence of the orbifold singularities. We consider both anti-twist and twist and show that the final expression can be written in a unified form, giving the spindle index [1]. We also include flavour fugacities as well as several classical terms, that were not discussed in [1]. As further applications of our formalism, in Section 4 we compute the full supersymmetric partition function of three-dimensional ${\cal N}=2$ Chern-Simons-matter theories on a branched and squashed lens space, again finding a result that unifies and generalises previous constructions regarding non-trivial circle (orbi-)bundles over two-dimensional (orbifold) Riemann surfaces. In particular, our partition function includes the squashed lens space one [31] and the branched sphere one [32]. Features of this observable and its possible applications will be discussed further in Section 6. In Section 5 we show how to recover the one-loop determinants relevant for the spindle index by viewing the three-sphere as an orbi-bundle over the spindle and reducing the three-sphere one-loop determinant, extending a method proposed in [33] to obtain quickly the two-sphere partition function, starting from the three-sphere. In Section 6 we discuss our findings, indicating a number of immediate applications as well as outlining a broader program for continuing the exploration of SQFTs on orbifolds that we initiated here. Further technical material is reported in the appendices.
2 Supersymmetric complex backgrounds
2.1 General analysis
We will employ the well-established framework of the rigid limit of new minimal supergravity in three-dimensions to study three-dimensonal supersymmetric ${\cal N}=2$ theories in curved backgrounds. The bosonic fields in the gravity multiplet are the metric and the auxiliary fields $A_{\mu},V_{\mu},H$, which in the rigid limit become background fields. In Euclidean signature $A_{\mu},V_{\mu},H$ are allowed to take complex values, while traditionally the metric $g_{\mu\nu}$ has been assumed to be real-valued. In view of some recent developments, in particular in the context of BPS black holes this turn out to be too restrictive and in this paper we will therefore allow the metric to take complex values.
The Killing-spinor equations (KSEs) read
${\mathcal{D}}_{\mu}\zeta=(\nabla_{\mu}-{\rm i}A_{\mu})\zeta=-{\frac{H}{2}}\gamma_{\mu}\zeta-{\rm i}V_{\mu}\zeta-{\frac{1}{2}}\varepsilon_{\mu\nu\rho}V^{\nu}\gamma^{\rho}\zeta~{},$
${\mathcal{D}}_{\mu}{\widetilde{\zeta}}=(\nabla_{\mu}+{\rm i}A_{\mu}){\widetilde{\zeta}}=-{\frac{H}{2}}\gamma_{\mu}{\widetilde{\zeta}}+{\rm i}V_{\mu}{\widetilde{\zeta}}+{\frac{1}{2}}\varepsilon_{\mu\nu\rho}V^{\nu}\gamma^{\rho}{\widetilde{\zeta}}~{},$ (2.1)
where in Euclidean the two Killing spinors $(\zeta,{\widetilde{\zeta}})$ of $R$-charges $(+1,-1)$ are independent. On the other hand, in Lorentzian signature, the second equation is simply as the charge conjugate of the first one, with ${\widetilde{\zeta}}=\zeta^{c}$. In previous literature [34, 14] the attention has been restricted to such “real” backgrounds, while here we will relax this assumption.
In order to encompass general backgrounds in which the metric can be complex and the Killing spinors $(\zeta,{\widetilde{\zeta}})$ are not related by charge conjugation, we will work with a “holomorphic” set of spinor bilinears, that are constructed without performing charge conjugation. In this way, we will never have to work with equations that contain the complex conjugate of the metric (or of the other background fields). In particular, we define the following complex-valued bilinears of Killing spinors:
${v}=\zeta{\widetilde{\zeta}}~{},\qquad K^{\mu}=\zeta\gamma^{\mu}{\widetilde{\zeta}}~{},\qquad P^{\mu}=\zeta\gamma^{\mu}\zeta/{v}~{},\qquad\widetilde{P}^{\mu}={\widetilde{\zeta}}\gamma^{\mu}{\widetilde{\zeta}}/{v}~{},$ (2.2)
where $\widetilde{P}^{\mu}\neq(P_{\mu})^{*}$. Using the Fierz identities we obtain
${v}^{-2}K_{(\mu}K_{\nu)}-\widetilde{P}_{(\mu}P_{\nu)}=g_{\mu\nu}~{},$ $\iota_{K}K^{\flat}={v}^{2}~{},\qquad\iota_{\widetilde{P}}P^{\flat}=\iota_{P}\widetilde{P}^{\flat}=-2~{},$ (2.3)
where $X^{\flat}=\iota_{X}g$ is the 1-form dual to the vector $X$ with respect to the three-dimensional metric tensor $g$. as well as the identities
$K_{[\mu}P_{\nu]}={\frac{{\rm i}{v}}{2}}\varepsilon_{\mu\nu\rho}P^{\rho}~{},\qquad K_{[\mu}\widetilde{P}_{\nu]}=-{\frac{{\rm i}{v}}{2}}\varepsilon_{\mu\nu\rho}\widetilde{P}^{\rho}~{},\qquad\widetilde{P}_{[\mu}P_{\nu]}={\frac{{\rm i}}{{v}}}\varepsilon_{\mu\nu\rho}K^{\rho}~{},$ (2.4)
so that the complex one-forms $(K,\widetilde{P},P)$ constitutes a three-dimensional “complex frame”. It is then straightforward to obtain the covariant derivatives of the Killing-spinor bilinears, that read
$\nabla_{\mu}{v}$ $=\varepsilon_{\mu\nu\rho}V^{\nu}K^{\rho}~{},$
$\nabla_{\mu}K_{\nu}$ $=\varepsilon_{\mu\nu\rho}({\rm i}HK^{\rho}-{v}V^{\rho})~{},$
${\left({\nabla_{\mu}-2{\rm i}A_{\mu}}\right)}P_{\nu}$ $=+{\frac{2H}{{v}}}K_{[\mu}P_{\nu]}-2{\rm i}V_{\mu}P_{\nu}+{\frac{{\rm i}}{{v}^{2}}}{\left[{(\iota_{P}V)K_{\mu}-(\iota_{K}V)P_{\mu}}\right]}K_{\nu}~{},$
${\left({\nabla_{\mu}+2{\rm i}A_{\mu}}\right)}\widetilde{P}_{\nu}$ $=-{\frac{2H}{{v}}}K_{[\mu}\widetilde{P}_{\nu]}+2{\rm i}V_{\mu}\widetilde{P}_{\nu}-{\frac{{\rm i}}{{v}^{2}}}{\left[{(\iota_{\widetilde{P}}V)K_{\mu}-(\iota_{K}V)\widetilde{P}_{\mu}}\right]}K_{\nu}~{}.$ (2.5)
In particular, $\mathcal{L}_{K}{v}=0$ and
$\nabla_{(\mu}K_{\nu)}=0~{},\qquad{\left[{\nabla_{\mu}-2{\rm i}{\left({A_{\mu}-{\frac{1}{2}}V_{\mu}}\right)}}\right]}P^{\mu}=0~{},\qquad{\left[{\nabla_{\mu}+2{\rm i}{\left({A_{\mu}-{\frac{1}{2}}V_{\mu}}\right)}}\right]}\widetilde{P}^{\mu}=0~{},$ (2.6)
ensuring that $K^{\mu}$ is a complex Killing vector. The equations (2.1) can be inverted to obtain expressions of the background fields in term of the bilinears. Firstly, we have
$\varepsilon^{\mu\nu\sigma}\nabla_{\mu}K_{\nu}$ $=2{\left({{\rm i}HK^{\sigma}-{v}V^{\sigma}}\right)}~{},$
$\varepsilon^{\mu\nu\sigma}\nabla_{\mu}{v}$ $=2V^{[\nu}K^{\sigma]}~{},$
$\varepsilon^{\mu\nu\sigma}{\left({\nabla_{\mu}-2{\rm i}A^{C}_{\mu}}\right)}P_{\nu}$ $={\frac{(\iota_{P}V)}{{v}}}K^{\sigma}+{\frac{2}{{v}}}{\left[{{\rm i}{v}H-(\iota_{K}V)}\right]}P^{\sigma}~{},$
$\varepsilon^{\mu\nu\sigma}{\left({\nabla_{\mu}+2{\rm i}A^{C}_{\mu}}\right)}\widetilde{P}_{\nu}$ $={\frac{(\iota_{\widetilde{P}}V)}{{v}}}K^{\sigma}+{\frac{2}{{v}}}{\left[{{\rm i}{v}H-(\iota_{K}V)}\right]}\widetilde{P}^{\sigma}~{},$ (2.7)
enjoying the shift symmetry
$H\to H+{\rm i}\,h~{},\qquad V_{\mu}\to V_{\mu}-(h/{v})K_{\mu}~{},\qquad A_{\mu}\to A_{\mu}-{\frac{3}{2}}(h/{v})K_{\mu}~{},$ (2.8)
with $h$ being a complex function satisfying $\mathcal{L}_{K}h=0$. The identities above immediately imply
$V^{\mu}$ $={\frac{(\iota_{K}V)}{{v}^{2}}}K^{\mu}+{\rm i}\widetilde{P}^{[\mu}P^{\nu]}\nabla_{\nu}\log{v}={\frac{{\rm i}}{{v}}}HK^{\mu}-{\frac{1}{2{v}}}\varepsilon^{\mu\nu\rho}\nabla_{\nu}K_{\rho}~{},$ (2.9)
namely
$V^{\mu}$ $={\frac{1}{{v}^{2}}}{\left({{\rm i}{v}H-{\frac{1}{2{v}}}K_{\lambda}\varepsilon^{\lambda\nu\rho}\nabla_{\nu}K_{\rho}}\right)}K^{\mu}+{\rm i}\widetilde{P}^{[\mu}P^{\nu]}\nabla_{\nu}\log{v}~{}.$ (2.10)
Moreover, by using (2.1), we can write down $A_{\mu}$ in terms of the derivatives of $K,P,\widetilde{P}$:
$A_{\mu}$ $=+{\frac{{\rm i}}{4}}\widetilde{P}^{\nu}\nabla_{\mu}P_{\nu}+{\frac{{\rm i}H}{2{v}}}K_{\mu}+V_{\mu}~{},$
$=-{\frac{{\rm i}}{4}}P^{\nu}\nabla_{\mu}\widetilde{P}_{\nu}+{\frac{{\rm i}H}{2{v}}}K_{\mu}+V_{\mu}~{},$
$=+{\frac{{\rm i}}{8}}{\left({\widetilde{P}^{\nu}\nabla_{\mu}P_{\nu}-P^{\nu}\nabla_{\mu}\widetilde{P}_{\nu}}\right)}+{\frac{{\rm i}H}{2{v}}}K_{\mu}+V_{\mu}~{}.$ (2.11)
In particular, the 1-form $A^{C}_{\mu}=A_{\mu}-{\frac{3}{2}}V_{\mu}$ appearing in the conformal Killing spinor equation is independent of the background scalar field $H$, indeed
$A^{C}_{\mu}$ $=A_{\mu}-{\frac{3}{2}}V_{\mu}~{},$
$={\frac{{\rm i}}{4}}\widetilde{P}^{\nu}\nabla_{\mu}P_{\nu}+{\frac{{\rm i}H}{2{v}}}K_{\mu}-{\frac{1}{2}}V_{\mu}~{},$
$={\frac{{\rm i}}{4}}\widetilde{P}^{\nu}\nabla_{\mu}P_{\nu}+{\frac{1}{2{v}^{2}}}{\left({{\frac{1}{2{v}}}K_{\lambda}\varepsilon^{\lambda\nu\rho}\nabla_{\nu}K_{\rho}}\right)}K^{\mu}+{\frac{{\rm i}}{2}}P^{[\mu}\widetilde{P}^{\nu]}\nabla_{\nu}\log{v}~{}.$ (2.12)
The background scalar field $H$ appears in the cohomological complex and, in turn, in the partition function of the theory only through the combination
$h_{R}=\iota_{K}V-{\rm i}{v}H=-{\frac{1}{2{v}}}K_{\mu}\varepsilon^{\mu\nu\rho}\nabla_{\nu}K_{\rho}~{},$ (2.13)
which is completely fixed by the Killing spinors $\zeta,{\widetilde{\zeta}}$ via the 1-form $K$ and the scalar $s$. As an example, the covariant Lie derivative $L_{K}$ can be written as follows:
$L_{K}$ $={\mathcal{L}}_{K}-{\rm i}q_{R}{\left({\iota_{K}A-{\frac{1}{2}}\iota_{K}V}\right)}-{\rm i}q_{Z}(\iota_{K}{\mathcal{C}})+{v}(q_{Z}-q_{R}H)~{},$
$={\mathcal{L}}_{K}-{\rm i}q_{R}{\left({\iota_{K}A^{C}+h_{R}}\right)}-{\rm i}q_{Z}(\iota_{K}{\mathcal{C}}+{\rm i}{v})~{},$ (2.14)
with $h_{R}$ representing the deviation from the background 1-form $A^{C}$ appearing in the CKSE. The field $h_{R}$ is invariant under the shift symmetry reported in (2.8). The linear combination
$\Phi_{R}=\iota_{K}A^{C}+h_{R}=\iota_{K}(A^{C}+V)-{\rm i}{v}H~{},$ (2.15)
is analogous to the field $\Phi_{G}$, built out of vector multiplet fields $\mathcal{A}_{\mu}$ and $\sigma$, as it is apparent by performing the replacements $(A^{C}+V)\to\mathcal{A}$ and $H\to\sigma$. The quantity $\Phi_{R}$ can be seen as a fugacity related to the R-symmetry and appears e.g. in the Lie derivatives along $K$ of the Killing spinors:
$\mathcal{L}_{K}\zeta=+{\rm i}\,\Phi_{R}\,\zeta~{},$ $\mathcal{L}_{K}{\widetilde{\zeta}}=-{\rm i}\,\Phi_{R}\,{\widetilde{\zeta}}~{}.$ (2.16)
Explicitly, we parametrize the Killing vector on the three-dimensional orbifold as
$K=K^{\mu}\partial_{\mu}={k_{0}}{\left({{\omega}\partial_{\varphi_{1}}+\partial_{\varphi_{2}}}\right)}={k_{0}}{\left({{\omega}\partial_{\varphi}+\partial_{\psi}}\right)}~{},$ (2.17)
with ${k_{0}}$ being a normalization constant and ${\omega}$ a fugacity for the angular momentum, while $\varphi_{1}=\varphi$ and $\varphi_{2}=\psi$ are angular coordinates with period $2\pi$. We consider a metric that is manifestly invariant with respect to both $\partial_{\varphi}$ and $\partial_{\psi}$:
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+{\mathrm{c}}_{ij}{\rm d}\varphi_{i}{\rm d}\varphi_{j}=\delta_{mn}e^{m}e^{n}~{},$ $i,j=1,2~{},\qquad m,n=1,2,3~{},$ (2.18)
where the metric functions $f=f(x)$ and ${\mathrm{c}}_{ij}={\mathrm{c}}_{ij}(x)$ are allowed to complex values and are independent of $\varphi_{1,2}$. The dreibein $e^{m}$ are
$e^{1}=-f{\rm d}x~{},\qquad e^{2}=\sqrt{\frac{\det{\mathrm{c}}}{{\mathrm{c}}_{22}}}{\rm d}\varphi~{},\qquad e^{3}=\sqrt{{\mathrm{c}}_{22}}{\left({\frac{{\mathrm{c}}_{12}}{{\mathrm{c}}_{22}}{\rm d}\varphi+{\rm d}\psi}\right)}~{},$ (2.19)
which define a frame regarding the three-dimensional orbifold as a circle fibration over a two-dimensional orbifold parametrized by ${\left({x,\varphi}\right)}$ with line element ${\rm d}s^{2}_{(2)}$, namely
${\rm d}s^{2}={\rm d}s^{2}_{(2)}+{\mathrm{c}}_{22}{\left({{A^{(1)}}+{\rm d}\psi}\right)}^{2}~{},$ ${\rm d}s^{2}_{(2)}=f^{2}{\rm d}x^{2}+\frac{\det{\mathrm{c}}}{{\mathrm{c}}_{22}}{\rm d}\varphi^{2}~{},\qquad{A^{(1)}}=\frac{{\mathrm{c}}_{12}}{{\mathrm{c}}_{22}}{\rm d}\varphi~{}.$ (2.20)
In this setting, the orthogonality relations (2.3) and the algebra (2.4) imply
${v}^{2}$ $={k_{0}}^{2}{\left({{\omega}^{2}c_{11}+2{\omega}c_{12}+c_{22}}\right)}~{},$
$K^{\flat}$ $={k_{0}}{\left[{{\left({{\omega}{\mathrm{c}}_{11}+{\mathrm{c}}_{12}}\right)}{\rm d}\varphi+{\left({{\omega}{\mathrm{c}}_{12}+{\mathrm{c}}_{22}}\right)}{\rm d}\psi}\right]}~{},$
$P^{\flat}$ $={\rm e}^{2{\rm i}\theta}{\left[{{\rm i}f{\rm d}x-{k_{0}}v^{-1}\sqrt{\det{\mathrm{c}}}{\left({{\rm d}\varphi-{\omega}{\rm d}\psi}\right)}}\right]}~{},$
$\widetilde{P}^{\flat}$ $={\rm e}^{-2{\rm i}\theta}{\left[{{\rm i}f{\rm d}x+{k_{0}}v^{-1}\sqrt{\det{\mathrm{c}}}{\left({{\rm d}\varphi-{\omega}{\rm d}\psi}\right)}}\right]}~{},$ (2.21)
where we parametrize the phase factor $\theta=\theta(x,\varphi,\psi)$ as
$\theta=\frac{1}{2}{\left[{\alpha_{1}{\left({x}\right)}+\alpha_{2}\varphi+\alpha_{3}\psi}\right]}~{}.$ (2.22)
The dual vectors $K,P$ and $\widetilde{P}$ can be found by contraction with the metric tensor, for instance
$P$ $={\rm e}^{2{\rm i}\theta}{\left\{{{\rm i}f^{-1}\partial_{x}-\frac{{k_{0}}}{v\sqrt{\det{\mathrm{c}}}}{\left[{{\left({{\omega}{\mathrm{c}}_{12}+{\mathrm{c}}_{22}}\right)}\partial_{\varphi}-{\left({{\omega}{\mathrm{c}}_{11}+{\mathrm{c}}_{12}}\right)}\partial_{\psi}}\right]}}\right\}}~{}.$ (2.23)
In turn, the definitions (2.2) determine the form of the Killing spinors $\zeta,\widetilde{\zeta}$ in a purely algebraic way, with no need to solve any differential equations. Indeed, the spinors compactly written as
$\zeta_{\alpha}={\rm e}^{{\rm i}\theta}\begin{pmatrix}u_{1}\\ -u_{2}\end{pmatrix}_{\alpha}~{},\qquad\widetilde{\zeta}_{\alpha}=-{\rm e}^{-{\rm i}\theta}\begin{pmatrix}u_{2}\\ u_{1}\end{pmatrix}_{\alpha}~{},$ (2.28)
$u_{1,2}$ $=\sqrt{\frac{v}{2}\mp\frac{{k_{0}}\,{\omega}}{2}{e^{2}}_{\varphi}}=\sqrt{\frac{v}{2}\mp\frac{{k_{0}}\,{\omega}}{2}\sqrt{\frac{\det{\mathrm{c}}}{{\mathrm{c}}_{22}}}}~{},$ (2.29)
satisfy the Killing spinor equations (2.1) if the background fields given by (2.10), (2.1) and (2.1) are turned on, namely
$A^{C}$ $={\rm d}\theta-\frac{v^{3}}{4{k_{0}}f\sqrt{\det{\mathrm{c}}}}{\left[{{\left({\frac{{\mathrm{c}}_{11}}{v^{2}}}\right)}^{\prime}{\rm d}\varphi-{\omega}^{-1}{\left({\frac{{\mathrm{c}}_{22}}{v^{2}}}\right)}^{\prime}{\rm d}\psi}\right]}~{},$
$V$ $={\rm d}\theta-A^{C}-\frac{{k_{0}}}{2vf}{\left({\sqrt{\det{\mathrm{c}}}}\right)}^{\prime}{\left({{\rm d}\varphi-{\omega}{\rm d}\psi}\right)}+\frac{{\rm i}H}{v}K^{\flat}~{},$
$=\frac{1}{4{k_{0}}f\sqrt{\det{\mathrm{c}}}}{\left\{{{\left[{v^{3}{\left({\frac{{\mathrm{c}}_{11}}{v^{2}}}\right)}^{\prime}-\frac{{k_{0}}^{2}}{v}{\left({\det{\mathrm{c}}}\right)}^{\prime}}\right]}{\rm d}\varphi-{\left[{\frac{v^{3}}{{\omega}}{\left({\frac{{\mathrm{c}}_{22}}{v^{2}}}\right)}^{\prime}-{\omega}\frac{{k_{0}}^{2}}{v}{\left({\det{\mathrm{c}}}\right)}^{\prime}}\right]}{\rm d}\psi}\right\}}+\frac{{\rm i}H}{v}K^{\flat}~{},$
$A$ $=A^{C}+\frac{3}{2}V=-\frac{1}{2}A^{C}+\frac{3}{2}{\left[{{\rm d}\theta-\frac{{k_{0}}}{2vf}{\left({\sqrt{\det{\mathrm{c}}}}\right)}^{\prime}{\left({{\rm d}\varphi-{\omega}{\rm d}\psi}\right)}+\frac{{\rm i}H}{v}K^{\flat}}\right]}~{},$ (2.30)
where $X^{\prime}$ stands for $\partial_{x}X$. Despite the appearance of ${\omega}^{-1}$ factors, by inspection the background fields above are finite in the limit ${\omega}\to 0$.
2.2 $\mathbb{\Sigma}\times S^{1}$ with twist and anti-twist
On $\mathbb{\Sigma}\times S^{1}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}\times S^{1}$ we adopt the line element
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+{\left({1-x^{2}}\right)}{\left({{\rm d}\varphi-\Omega{\rm d}\psi}\right)}^{2}+\beta^{2}{\rm d}\psi^{2}~{},$ (2.31)
with $x\in{\left[{-1,+1}\right]}$ and $\varphi\in[0,2\pi)$ being coordinates on $\mathbb{\Sigma}$ while $\psi\in[0,2\pi)$ is the compactified Euclidean time parametrizing $S^{1}$. The function $f=f(x)$ satisfies
$\lim_{x\to\pm 1}f\to\frac{n_{\pm}}{\sqrt{2{\left({1\mp x}\right)}}}~{}.$ (2.32)
We require that $f$ is regular on the complementary domain ${\left({-1,+1}\right)}$, but we leave arbitrary the specific profile of $f$. The positive integers $n_{\pm}$ are coprime and encode the conical singularities $\mathbb{C}/\mathbb{Z}_{n_{\pm}}$ at the poles of the spindle. The dimensionless complex parameter $\Omega$ induces a refinement of the partition function by a fugacity for the angular momentum on the spindle, as it happens on the round sphere [35]. The parameter $\beta$ is the ratio between the radius of $S^{1}$ and the radius of the equatorial circle of $\mathbb{\Sigma}$. The integral of the Ricci scalar of the spindle is
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}x{\rm d}\varphi\sqrt{g_{\mathbb{\Sigma}}}R_{\mathbb{\Sigma}}=\frac{1}{{n_{+}}}+\frac{1}{{n_{-}}}~{},$ (2.33)
correctly reproducing the orbifold Euler characteristic $\chi_{\rm orb}$ of $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$.
We use an orthonormal frame explicitly realizing $\mathbb{\Sigma}\times S^{1}$ as a $U(1)$ fibration over the spindle base:
$e^{1}$ $=-f{\rm d}x~{},$
$e^{2}$ $=\beta\sqrt{\frac{1-x^{2}}{\beta^{2}+\Omega^{2}{\left({1-x^{2}}\right)}}}{\rm d}\varphi~{},$
$e^{3}$ $=\sqrt{\beta^{2}+\Omega^{2}{\left({1-x^{2}}\right)}}{\left({{A^{(1)}}+{\rm d}\psi}\right)}~{},\qquad{A^{(1)}}=\frac{-\Omega{\left({1-x^{2}}\right)}}{\beta^{2}+\Omega^{2}{\left({1-x^{2}}\right)}}{\rm d}\varphi~{},$ (2.34)
where the flux of ${A^{(1)}}$ through $\mathbb{\Sigma}$ is vanishing, as it behooves a trivial circle fibration over a spindle. Near the north pole of $\mathbb{\Sigma}$ at $x=+1$ we have
$e^{1}_{{\mathcal{U}_{+}}}=-\frac{{n_{+}}{\rm d}x}{\sqrt{2{\left({1-x}\right)}}}={\rm d}{\rho_{+}}~{},\qquad e^{2}_{{\mathcal{U}_{+}}}=\sqrt{2{\left({1-x}\right)}}{\rm d}\varphi=\frac{{\rho_{+}}}{{n_{+}}}{\rm d}\varphi~{},$ ${\rho_{+}}={n_{+}}\sqrt{2{\left({1-x}\right)}}~{}.$ (2.35)
Hence, an open neighbourhood of $x=+1$ can be parametrized by the complex coordinate $z_{+}$ satisfying
$z_{+}={\rho_{+}}e^{{\rm i}\varphi/{n_{+}}}~{},\qquad z_{+}\sim w_{+}z_{+}~{},$ $w_{+}=e^{2\pi{\rm i}/{n_{+}}}~{},$ (2.36)
where the identification involving the root of unity $w_{+}$ makes manifest that ${\mathcal{U}_{+}}$ is isomorphic to $\mathbb{C}/{n_{+}}$. On the other hand, near the south pole of the spindle at $x=-1$ we have
$e^{1}_{S}=-\frac{{n_{-}}{\rm d}x}{\sqrt{2{\left({1+x}\right)}}}=-{\rm d}{\rho_{+}}~{},\qquad e^{2}_{S}=\sqrt{2{\left({1+x}\right)}}{\rm d}\varphi=\frac{{\rho_{-}}}{{n_{-}}}{\rm d}\varphi~{},$ ${\rho_{-}}={n_{-}}\sqrt{2{\left({1+x}\right)}}~{}.$ (2.37)
Thus, a proper coordinate in an open neighbourhood ${\mathcal{U}_{-}}$ of $x=-1$ is
$z_{-}={\rho_{-}}e^{-{\rm i}\varphi/{n_{-}}}~{},\qquad z_{-}\sim w_{-}^{-1}z_{-}~{},$ $w_{-}=e^{2\pi{\rm i}/{n_{-}}}~{},$ (2.38)
meaning that ${\mathcal{U}_{-}}\cong\mathbb{C}/{n_{-}}$. The minus sign in the exponential of $z_{-}={\rho_{-}}e^{-{\rm i}\varphi/{n_{-}}}$ takes into account the reversed orientation of ${\mathcal{U}_{-}}$ with respect to that of ${\mathcal{U}_{+}}$. A necessary condition for the orbifold $\mathbb{\Sigma}\times S^{1}$ to be supersymmetric is that the magnetic flux of the $U(1)$ R-symmetry field $A^{C}$ through the spindle takes the form
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}A^{C}=\frac{1}{2}{\left({\frac{1}{{n_{-}}}+\frac{{\mathfrak{s}}}{{n_{+}}}}\right)}=\frac{\chi_{\mathfrak{s}}}{2}~{},$ ${\mathfrak{s}}=\pm 1~{}.$ (2.39)
The values of the sign ${\mathfrak{s}}=\pm 1$ represent two inequivalent ways to preserve supersymmetry on $\mathbb{\Sigma}\times S^{1}$. In the case of ${\mathfrak{s}}=+1$, the magnetic flux of the R-symmetry background field $A^{C}$ equals half of the orbifold Euler characteristic of the spindle $\chi_{\rm orb}={n_{-}}^{-1}+{n_{+}}^{-1}$. This configuration is referred to as topological twist as it is the orbifold generalization of the analogous setup found in spheres [35, 36] and Riemann surfaces [37]. On the other hand, when ${\mathfrak{s}}=-1$, the magnetic flux of the R-symmetry background field equals half of $\chi_{-}={\left({{n_{-}}^{-1}-{n_{+}}^{-1}}\right)}$. This specific configuration is dubbed anti-twist to highlight its contrast with the topological twist, primarily due to the presence of a relative minus sign between ${n_{+}}^{-1}$ and ${n_{-}}^{-1}$, resulting in $\chi_{-}$ being radically distinct from $\chi_{\rm orb}=\chi_{+}$. In the special case where ${n_{+}}={n_{-}}=1$ the spindle become a smooth sphere, and the anti-twisted $\mathbb{\Sigma}\times S^{1}$ configuration reduces to the generalized superconformal index background studied in [38, 39], characterized by the absence of magnetic flux for the R-symmetry connection. Eventually, after shrinking to zero the radius of $S^{1}$ in $\mathbb{\Sigma}\times S^{1}$ and keeping ${n_{+}}={n_{-}}=1$, the anti-twisted spindle becomes the supersymmetric sphere with no R-symmetry twist explored in [33].
The Killing spinors corresponding to the twisted $\mathbb{\Sigma}\times S^{1}$ geometry, occurring if ${\mathfrak{s}}=+1$, are
$\zeta_{\alpha}=e^{{\rm i}\theta}\begin{pmatrix}u_{1}\\ -u_{2}\end{pmatrix}_{\alpha}~{},\qquad\widetilde{\zeta}_{\alpha}=-e^{-{\rm i}\theta}\begin{pmatrix}u_{2}\\ u_{1}\end{pmatrix}_{\alpha}~{},$ (2.44)
$u_{1,2}=\sqrt{-\frac{{k_{0}}\beta}{2}{\left({1\pm\Omega\sqrt{\frac{1-x^{2}}{\beta^{2}+\Omega^{2}{\left({1-x^{2}}\right)}}}}\right)}}~{},$ (2.45)
which satisfy the conformal Killing spinor equations provided that the background fields $A^{C},V,A$ take the form
$A^{C}$ $={\rm d}\theta-\frac{x}{2f\sqrt{1-x^{2}}}{\left({{\rm d}\varphi-\Omega{\rm d}\psi}\right)}~{},$
$V$ $=-{\rm i}\beta H{\rm d}\psi~{},$
$A$ $=A^{C}+\frac{3}{2}V~{}.$ (2.46)
Unlike the conventional topological twist, the R-symmetry field $A^{C}$ above is unrelated to the spin connection. Nevertheless, the flux of $A^{C}$ through the spindle $\mathbb{\Sigma}$ precisely equals $\chi_{\rm orb}/2$, as desired. Moreover, we observe that the fluxes of $A^{C}$ and $A$ coincide with each other. The 1-form fields ${\left({A^{C},V,A}\right)}$ and $\zeta,\widetilde{\zeta}$ are regular in the northern patch ${\mathcal{U}_{+}}$ and the southern patch ${\mathcal{U}_{-}}$ of the spindle if $\theta$ takes the form
$\alpha_{2}|_{{\mathcal{U}_{+}}}=\frac{1}{{n_{+}}}~{},\qquad\alpha_{2}|_{{\mathcal{U}_{-}}}=-\frac{1}{{n_{-}}}~{},$ $\alpha_{3}$ $\in\mathbb{Z}~{},$
$\theta=\frac{1}{2}{\left[{\alpha_{1}(x)+\alpha_{2}\varphi+\alpha_{3}\psi}\right]}~{},$ (2.47)
where $\alpha_{2}|_{N,S}$ is the value of $\alpha_{2}$ in the patch covering $x=\pm 1$, respectively.
Conversely, the Killing spinors associated with the anti-twisted $\mathbb{\Sigma}\times S^{1}$ geometry, which occurs when ${\mathfrak{s}}=-1$, exhibit a formal similarity to the Killing spinors corresponding to the twisted background. However, they differ in terms of their components $u_{1,2}$:
$u_{1,2}=\sqrt{\frac{{k_{0}}\beta}{2}{\left[{x\mp{\left({\Omega-{\rm i}\beta}\right)}\sqrt{\frac{1-x^{2}}{\beta^{2}+\Omega^{2}{\left({1-x^{2}}\right)}}}\,}\right]}}~{},$ (2.48)
fulfilling the conformal Killing spinor equations if the background fields $A^{C},V,A$ are
$A^{C}$ $={\rm d}\theta+\frac{1}{2f\sqrt{1-x^{2}}}{\left[{{\rm d}\varphi-{\left({\Omega+{\rm i}\beta}\right)}{\rm d}\psi}\right]}~{},$
$V$ $={\left({x^{-1}+x}\right)}H{\rm d}\varphi+{\rm i}{\left\{{\frac{\beta}{f\sqrt{1-x^{2}}}+x^{-1}{\left[{\beta+{\rm i}\,\Omega{\left({1-x^{2}}\right)}}\right]}H}\right\}}{\rm d}\psi~{},$
$A$ $=A^{C}+\frac{3}{2}V~{}.$ (2.49)
Consistently, the flux of the R-symmetry connection $A^{C}$ through the spindle is ${\left({\chi_{-}/2}\right)}$. To ensure that the 1-form $V$ remains regular over the entire $\mathbb{\Sigma}\times S^{1}$, we require that $H\sim\mathcal{O}(x)$ in the vicinity of the spindle’s equator at $x=0$. The background fields ${\left({A^{C},V,A}\right)}$ and $\zeta,\widetilde{\zeta}$ are non-singular in the patches ${\mathcal{U}_{+}}$ and ${\mathcal{U}_{-}}$ of $\mathbb{\Sigma}$ if $\theta$ is
$\alpha_{2}|_{{\mathcal{U}_{+}}}=-\frac{1}{{n_{+}}}~{},\qquad\alpha_{2}|_{{\mathcal{U}_{-}}}=-\frac{1}{{n_{-}}}~{},$ $\alpha_{3}$ $\in\mathbb{Z}~{},$
$\theta=\frac{1}{2}{\left[{\alpha_{1}(x)+\alpha_{2}\varphi+\alpha_{3}\psi}\right]}~{},$ (2.50)
where again $\alpha_{2}|_{N,S}$ are the values of $\alpha_{2}$ in the patches that respectively include the poles $x=\pm 1$.
2.3 Conformal boundary of the accelerating black hole
In [25] the thermodynamic features of asymptotically ${\rm AdS}_{4}$ black holes carrying electromagnetic charge, angular momentum and acceleration along the axis of rotation were explored in the context of Einstein-Maxwell theory. In particular, it was shown that the conformal boundary of such a class of black holes is the three-dimensional orbifold ×$\mathbb{R}\times\mathbb{\Sigma}$, where $\mathbb{R}$ encodes time while the spindle $\mathbb{\Sigma}$ characterizes the spatial part of the boundary. Specifically, the conformal boundary of this accelerating black holes is described by the Lorentzian line element
${\rm d}\widehat{s}^{2}=-{\left({\widehat{e}^{0}}\right)}^{2}+{\left({\widehat{e}^{1}}\right)}^{2}+{\left({\widehat{e}^{2}}\right)}^{2}~{},$ (2.51)
with
$\widehat{e}^{0}=\sqrt{\widetilde{P}}{\left({\frac{{\rm d}t}{\kappa}-af{\rm d}\phi}\right)}~{},\qquad\widehat{e}^{1}=-\sqrt{F}\,{\rm d}x~{},\qquad\widehat{e}^{2}=\sqrt{G}\,{\rm d}\phi~{},$ (2.52)
being the 1-forms $\widehat{e}^{1,2,3}$ comprising the corresponding Lorentzian frame. The functions appearing in the line element ${\rm d}\widehat{s}^{2}$ read
$P$ $=1-2\alpha mx+{\left[{\alpha^{2}{\left({a^{2}+e^{2}+g^{2}}\right)}-a^{2}}\right]}x^{2}~{},$
$\widetilde{P}$ $=1-\alpha^{2}{\left({1-x^{2}}\right)}P~{},$
$f$ $=\widetilde{P}^{-1}{\left({1-\alpha^{2}P}\right)}{\left({1-x^{2}}\right)}~{},$
$F$ $=\frac{{\left({1+a^{2}\alpha^{2}x^{4}}\right)}^{2}}{P{\left({\widetilde{P}+a^{2}\alpha^{2}x^{4}}\right)}{\left({1-x^{2}}\right)}}~{},$
$G$ $=\widetilde{P}^{-1}P{\left({\widetilde{P}+a^{2}\alpha^{2}x^{4}}\right)}{\left({1-x^{2}}\right)}~{},$ (2.53)
while the parameters ${\left({a,e,g,m,\alpha}\right)}$ can be expressed in terms of fewer constants ${\left({b,s,\mu}\right)}$ according to the following relations [25]:
$\mu$ $=\frac{{n_{-}}+{n_{+}}}{{n_{-}}-{n_{+}}}~{},\qquad g=\alpha m~{},\qquad e=\frac{bs}{\alpha^{2}c}~{},\qquad g=\frac{s}{\alpha^{2}c}~{},\qquad a=\frac{s}{\alpha}~{},$
$c$ $=\frac{2{\left({1+b^{2}}\right)}s\mu}{1-2bs-s^{2}}~{},\qquad\kappa=\frac{{\left({b+s}\right)}{\left({1-bs}\right)}}{\alpha{\left({1+b^{2}}\right)}{\left({1+s^{2}}\right)}}~{},\qquad\alpha^{2}=\frac{4\mu^{2}{\left({1-bs}\right)}^{2}-{\left({1-2bs-s^{2}}\right)}^{2}}{4\mu^{2}{\left({1+b^{2}}\right)}{\left({1+s^{2}}\right)}}~{}.$ (2.54)
The Euclidean version of the Lorentzian conformal boundary above is obtained upon performing the Wick rotation $t=-{\rm i}\tau$ and moving to the dreibein $e^{1,2,3}$ given by
$e^{1}=\widehat{e}^{1}=-\sqrt{F}\,{\rm d}x~{},\qquad e^{2}=\widehat{e}^{2}=\sqrt{G}\,{\rm d}\phi~{},\qquad e^{3}={\rm i}\,\widehat{e}^{0}=\sqrt{\widetilde{P}}{\left({\frac{{\rm d}\tau}{\kappa}-{\rm i}af{\rm d}\phi}\right)}~{},$ (2.55)
which constitute an orthonormal frame yielding in turn the Euclidean line element
${\rm d}s^{2}={\left({e^{1}}\right)}^{2}+{\left({e^{2}}\right)}^{2}+{\left({e^{3}}\right)}^{2}~{}.$ (2.56)
This background is endowed with the Killing vector
$K={\rm i}\kappa{\left({1+a^{2}\alpha^{2}}\right)}\partial_{\tau}+\alpha{\left({\frac{e}{g}+a\alpha}\right)}\partial_{\phi}~{},$ (2.57)
which can be written in the notation of (2.17) if
${k_{0}}={\rm i}\kappa{\left({1+a^{2}\alpha^{2}}\right)}~{},$ ${\omega}=\frac{\alpha{\left({\frac{e}{g}+a\alpha}\right)}}{{\rm i}\kappa{\left({1+a^{2}\alpha^{2}}\right)}}~{}.$ (2.58)
The norm of $K$ is
${v}=e^{{\rm i}\pi/2}{\left[{\frac{1-2bs-s^{2}}{2\mu}{\left({1-x^{2}}\right)}+{\left({1+s^{2}}\right)}x}\right]}~{},$ (2.59)
and plugging $v$ and ${e^{2}}_{\phi}$ into our general formulae for the Killing spinors written in (2.28) indeed yields objects of the form
$\zeta_{\alpha}=e^{{\rm i}\theta}\begin{pmatrix}u_{1}\\ -u_{2}\end{pmatrix}_{\alpha}~{},\qquad\widetilde{\zeta}_{\alpha}=-e^{-{\rm i}\theta}\begin{pmatrix}u_{2}\\ u_{1}\end{pmatrix}_{\alpha}~{},$ (2.64)
with arbitrary $\theta$ and
$u_{1,2}=e^{{\rm i}\pi/4}\sqrt{\frac{1}{2}{\left[{\frac{1-2bs-s^{2}}{2\mu}{\left({1-x^{2}}\right)}+{\left({1+s^{2}}\right)}x\pm{\rm i}\alpha{\left({b+s}\right)}\sqrt{G}}\right]}}~{}.$ (2.65)
Especially, if we choose
$\alpha_{1}(x)$ $=-\frac{\pi}{2}~{},$
$\alpha_{2}$ $=-\frac{4\mu^{2}(1+s^{2})(1-2bs-s^{2})}{4\mu^{2}(1-bs)^{2}-(1-2bs-s^{2})^{2}}~{},$
$\alpha_{3}$ $=\frac{{\rm i}{\left({1+s^{2}}\right)}}{1-bs}~{},$
$\theta$ $=\frac{1}{2}{\left[{\alpha_{1}(x)+\alpha_{2}\phi+\alpha_{3}\tau}\right]}~{},$ (2.66)
our $\zeta_{\alpha}$ and $\widetilde{\zeta}_{\alpha}$ coincide with the boundary Killing spinors found in [25]. Consistently, inserting the values above into our general formulae for the background fields reported in (2.1) gives
$A^{C}$ $=-\frac{x{\left[{g+a^{2}g\alpha^{2}x^{2}-ae\alpha{\left({1-x^{2}}\right)}}\right]}}{1+a^{2}\alpha^{2}x^{4}}{\rm d}\phi+\frac{{\rm i}\,x\alpha{\left({e-ag\alpha x^{2}}\right)}}{\kappa{\left({1+a^{2}\alpha^{2}x^{4}}\right)}}{\rm d}\tau~{},$
$V$ $={\rm d}\theta-A^{C}-\frac{{k_{0}}}{2vf}{\left({\frac{1+a^{2}\alpha^{2}x^{4}}{\kappa\,f}}\right)}^{\prime}{\left({{\rm d}\varphi-{\omega}{\rm d}\psi}\right)}+\frac{{\rm i}H}{v}K^{\flat}~{},$
$A$ $=A^{C}+\frac{3}{2}V~{},$ (2.67)
where, in particular, $A^{C}$ matches the $U(1)_{R}$ gauge field appearing in the (boundary) conformal Killing spinor equations solved in [25].
2.4 The squashed lens space
Given a squashed three-sphere $S^{3}_{b_{1},b_{2}}$ with deformation parameters ${\left({b_{1},b_{2}}\right)}$ and a positive integer $n$, the corresponding squashed lens space $L_{b_{1},b_{2}}{\left({n,1}\right)}$ can be obtained from the free quotient $S^{3}_{b_{1},b_{2}}/\mathbb{Z}_{n}$, where the orbifold projection acts on the Hopf fiber of $S^{3}_{b_{1},b_{2}}$. Topologically, $L_{b_{1},b_{2}}{\left({n,1}\right)}$ is an $\mathcal{O}(-n)$ circle fibration with base a two-dimensional ellipsoid. The non-trivial features of gauge theories on lens spaces were first investigated in [40, 41], the corresponding large-$N$ limit was explored in [42] while their dualities were studied in [31]. Furthermore, lens spaces are important in the context of holomorphic blocks and factorization of supersymmetric partition functions, as shown in [43, 44].
On $L_{b_{1},b_{2}}{\left({n,1}\right)}$ we use the line element
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+b_{1}^{2}\sin^{2}x{\left({{\rm d}\varphi+\frac{{\rm d}\psi}{n}}\right)}^{2}+\frac{b_{2}^{2}}{n^{2}}\cos^{2}x\,{\rm d}\psi^{2}~{},$ (2.68)
with $\varphi,\psi\in{\left[{0,2\pi}\right]}$ being periodic coordinates and $x\in{\left[{0,\pi/2}\right]}$. The function
$f=-\sqrt{b_{1}^{2}\cos^{2}x+b_{2}^{2}\sin^{2}x}~{},$ (2.69)
produces a smooth squashing of the manifold tuned by the complex numbers $b_{1,2}$. The presence of the positive integer $n$ encodes the $\mathbb{Z}_{n}$ orbifold acting on the $S^{1}$ parametrized by the coordinate $\psi$. If $n=1$, the squashed lens space reduces to the squashed three-sphere explored in [13].
We employ the following orthonormal frame:
$e^{1}$ $=-f\,{\rm d}x~{},$
$e^{2}$ $=\frac{b_{1}b_{2}\sin x\cos x}{\sqrt{b_{1}^{2}\sin^{2}x+b_{2}^{2}\cos^{2}x}}{\rm d}\varphi~{},$
$e^{3}$ $=n^{-1}\sqrt{b_{1}^{2}\sin^{2}x+b_{2}^{2}\cos^{2}x}{\left({{A^{(1)}}+{\rm d}\psi}\right)}~{},\qquad{A^{(1)}}=\frac{n\,b_{1}^{2}\sin^{2}x}{b_{1}^{2}\sin^{2}x+b_{2}^{2}\cos^{2}x}{\rm d}\varphi~{},$ (2.70)
which realize the squashed lens space as a non-trivial circle fibration over the base $\mathbb{\Sigma}$, where the latter is an ellipsoid with coordinates ${\left({x,\varphi}\right)}$. Indeed, the Euler characteristic of $\mathbb{\Sigma}$ coincides with that of a two-sphere,
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}x{\rm d}\varphi\sqrt{g_{\mathbb{\Sigma}}}R_{\mathbb{\Sigma}}=2=\chi_{S^{2}}~{},$ (2.71)
while the 1-form ${A^{(1)}}$ is a representative of an $\mathcal{O}(-n)$ bundle over $\mathbb{\Sigma}$:
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{A^{(1)}}=n~{}.$ (2.72)
This geometry admits a Killing vector of the form (2.17):
$K={k_{0}}{\left({-\frac{b_{1}+b_{2}}{n\,b_{1}}\partial_{\varphi}+\partial_{\psi}}\right)}~{},$ (2.73)
with norm
${v}={k_{0}}b_{2}/n~{}.$ (2.74)
In this setting, the Killing spinors (2.28) have components
$u_{1,2}$ $=\sqrt{\frac{{k_{0}}b_{2}}{2n}{\left({1\pm\frac{{\left({b_{1}+b_{2}}\right)}\sin x\cos x}{\sqrt{b_{1}^{2}\sin^{2}x+b_{2}^{2}\cos^{2}x}}}\right)}}~{},$ (2.75)
where the phase $\theta$ can be fixed by requiring that the background fields (2.1), namely
$A^{C}$ $={\rm d}\theta-\frac{b_{1}}{2f}{\rm d}\varphi-\frac{b_{1}-b_{2}}{2nf}{\rm d}\psi~{},$
$V$ $=\frac{b_{1}{\left({1-{\rm i}fH}\right)}\sin^{2}x}{f}{\rm d}\varphi+\frac{{\left[{b_{1}-b_{2}-{\left({b_{1}+b_{2}}\right)}\cos{\left({2x}\right)}}\right]}{\left({1-{\rm i}fH}\right)}}{2nf}{\rm d}\psi~{},$
$A$ $=A^{C}+\frac{3}{2}V~{},$ (2.76)
are non-singular on the whole squashed lens space. By inspection, regularity of both ${\left({A^{C},V,A}\right)}$ and $\zeta,\widetilde{\zeta}$ is achieved if
$\alpha_{2}$ $=-1~{},$
$\alpha_{3}$ $=0~{},$
$\theta$ $=\frac{1}{2}{\left[{\alpha_{1}(x)+\alpha_{2}\varphi+\alpha_{3}\psi}\right]}~{}.$ (2.77)
2.5 The branched, squashed three-sphere
Supersymmetric gauge theories on branched spheres, namely branched coverings of $S^{3}$, were explored in [32] because of their connection to Rényi entropy as well as entanglement entropy. Analogous results for free theories on branched spheres were previously obtained in [45]. Here we consider supersymmetric gauge theories on a more general orbifold, namely a branched squashed three-sphere $S^{3}_{q,p,{n_{+}},{n_{-}}}$, whose line element reads
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+\frac{q^{2}}{{n_{+}}^{2}}\sin^{2}x\,{\rm d}\varphi^{2}+\frac{p^{2}}{{n_{-}}^{2}}\cos^{2}x\,{\rm d}\psi^{2}~{},$ (2.78)
where $\varphi,\psi\in{\left[{0,2\pi}\right]}$ parametrize the two circles $S^{1}$ in $S^{3}_{q,p,{n_{+}},{n_{-}}}$, while $x\in{\left[{0,\pi/2}\right]}$. The (a priori complex) numbers $q,p$ are the squashing parameters also appearing in the function
$f=-\sqrt{p^{2}\sin^{2}x+q^{2}\cos^{2}x}~{}.$ (2.79)
The positive integers ${\left({{n_{+}},{n_{-}}}\right)}$ parametrize the orbifold singularities at $x=0$ and $x=\pi/2$. If ${n_{+}}={n_{-}}=1$, we obtain the squashed sphere studied in [12]. Instead, the metric of the branched sphere considered in [32] is obtained from (2.78) after choosing $q=p=1$ and identifying ${n_{+}}=q^{-1}_{\rm there}$ as well as ${n_{-}}=p^{-1}_{\rm there}$. Further setting ${n_{+}}=1$ yields the line element of the orbifold employed in [45].
We work in the diagonal frame
$e^{1}$ $=-f\,{\rm d}x~{},$
$e^{2}$ $=q\sin x\,{\rm d}\varphi/{n_{+}}~{},$
$e^{3}$ $=p\cos x\,{\rm d}\psi/{n_{-}}~{}.$ (2.80)
The branched,squashed three-sphere is linked to the other orbifold backgrounds explored in this paper as $S^{3}_{q,p,{n_{+}},{n_{-}}}$ is a non-trivial $S^{1}$ fibration with base a spindle $\mathbb{\Sigma}$. Indeed, by applying a $SL{\left({2,\mathbb{Z}}\right)}$ map acting on ${\left({\varphi,\psi}\right)}$ as
$\varphi=t_{11}\widehat{\varphi}_{1}+t_{12}\widehat{\varphi}_{2}~{},\qquad\psi=t_{21}\widehat{\varphi}_{1}+t_{22}\widehat{\varphi}_{2}~{},$ $t_{11}t_{22}-t_{12}t_{21}=1~{},$ (2.81)
with $t_{ij}\in\mathbb{Z}$, the line element (2.78) can be rewritten as
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+c_{22}{\rm d}\widehat{\varphi}_{1}^{2}+h{\left({{\rm d}\widehat{\varphi}_{2}+{A^{(1)}}}\right)}^{2}~{},$ (2.82)
where $\widehat{\varphi}_{2}\in{\left[{0,2\pi}\right]}$ parametrizes $S^{1}$ while $x\in{\left[{0,\pi/2}\right]}$ and $u_{1}\in{\left[{0,2\pi}\right]}$ are coordinates on $\mathbb{\Sigma}$. The objects appearing in the line element immediately above read
$c_{22}$ $=\frac{g_{22}g_{33}}{g_{22}t_{12}^{2}+g_{33}t_{22}^{2}}~{},$
$h$ $=g_{22}t_{12}^{2}+g_{33}t_{22}^{2}~{},$
${A^{(1)}}$ $=\frac{g_{22}t_{11}t_{12}+g_{33}t_{21}t_{22}}{g_{22}t_{12}^{2}+g_{33}t_{22}^{2}}{\rm d}\widehat{\varphi}_{1}~{},$
$g_{22}$ $=q^{2}\sin^{2}x/{n_{+}}^{2}~{},$
$g_{33}$ $=p^{2}\cos^{2}x/{n_{-}}^{2}~{}.$ (2.83)
The Euler characteristic of $\mathbb{\Sigma}$ is
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}x{\rm d}\widehat{\varphi}_{1}\sqrt{g_{\mathbb{\Sigma}}}R_{\mathbb{\Sigma}}=\frac{1}{{n_{-}}t_{12}}+\frac{1}{{n_{+}}t_{22}}=\chi_{\rm orb}~{},$ (2.84)
confirming that the branched squashed three-sphere is a $S^{1}$ fibration over the base encoded by the spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}}t_{22},{n_{-}}t_{12}}\right]}}$. Moreover,the flux of $A_{1}$ through $\mathbb{\Sigma}$ is non-vanishing and fractional:
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{A^{(1)}}=\frac{t_{11}}{t_{12}}-\frac{t_{21}}{t_{22}}=\frac{{n_{+}}{n_{-}}}{{\left({{n_{+}}t_{22}}\right)}{\left({{n_{-}}t_{12}}\right)}}~{},$ (2.85)
verifying that the fibration is non-trivial and that the background field $A_{1}$ is a 1-form representing a $\mathcal{O}(-{n_{+}}{n_{-}})$ orbibundle on $\mathbb{\Sigma}$.
The Killing spinors
$\zeta_{\alpha}$ $=\sqrt{\frac{\mathcal{N}p}{2\,{n_{-}}}}\,e^{\frac{{\rm i}}{2}{\left({\alpha_{2}\varphi+\alpha_{3}\psi}\right)}}\sqrt{1-\mathfrak{s}_{w}\sin x}\begin{pmatrix}1\\ \sec x+\mathfrak{s}_{w}\tan x\end{pmatrix}_{\alpha}~{},$ (2.88)
$\widetilde{\zeta}_{\alpha}$ $=-\sqrt{\frac{\mathcal{N}p}{2\,{n_{-}}}}\,e^{-\frac{{\rm i}}{2}{\left({\alpha_{2}\varphi+\alpha_{3}\psi}\right)}}\sqrt{1-\mathfrak{s}_{w}\sin x}\begin{pmatrix}\sec x+\mathfrak{s}_{w}\tan x\\ 1\end{pmatrix}_{\alpha}~{},$ (2.91)
with $\mathfrak{s}_{w}=\pm 1$ being a sign, solve the three-dimensional $\mathcal{N}=2$ Conformal Killing-spinor equation if the background-field
$A^{C}$ $=\frac{1}{2}{\left({-\frac{q}{{n_{+}}f}+\alpha_{2}}\right)}{\rm d}\varphi+\frac{1}{2}{\left({-\frac{\mathfrak{s}_{w}p}{{n_{-}}f}+\alpha_{3}}\right)}{\rm d}\psi~{},$ (2.92)
is turned on. The integrability conditions for this background are verified for any $\alpha_{2,3}$, even though the 1-form $A^{C}$ vanishes at the conical points of the branched, squashed three-sphere only if
$\alpha_{2}=-1/{n_{+}}~{},$ $\alpha_{3}=-\mathfrak{s}_{w}/{n_{-}}~{}.$ (2.93)
Furthermore, the spinors (2.88) solve the three-dimensional $\mathcal{N}=2$ Killing-spinor equation with background fields
$V$ $={\left({\frac{1}{f}+{\rm i}\,\mathfrak{s}_{w}H}\right)}{\left({\frac{q}{{n_{+}}}\sin^{2}x\,{\rm d}\varphi+\mathfrak{s}_{w}\frac{\,p}{{n_{-}}}\cos^{2}x\,{\rm d}\psi}\right)}~{},$
$A$ $=A^{C}+\frac{3}{2}V~{}.$ (2.94)
2.6 The branched lens space
The three-dimensional orbifold encompassing the squashed three-sphere, the branched sphere and the squashed lens space is the branched lens space ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ with squashing parameters ${\left({b_{1},b_{2}}\right)}$ and conical singularities parametrized by the positive integers ${\left({{n_{+}},{n_{-}}}\right)}$. The branched lens space is a $\mathcal{O}(-n{n_{+}}{n_{-}})$ circle fibration over the spindle given by the weighted projective space $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$. Equivalently, ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ can be viewed as the free quotient $S^{3}_{b_{1},b_{2},{n_{+}},{n_{-}}}/\mathbb{Z}_{n}$ acting upon the Hopf fiber of the branched, squashed three-sphere. A line element on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ is
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+{\left({\frac{b_{1}}{{n_{+}}}}\right)}^{2}\sin^{2}x{\left({{\rm d}\varphi+\frac{{\rm d}\psi}{n}}\right)}^{2}+{\left({\frac{b_{2}}{n\,{n_{-}}}}\right)}^{2}\cos^{2}x\,{\rm d}\psi^{2}~{},$ (2.95)
where $\varphi,\psi\in{\left[{0,2\pi}\right]}$ are periodic and again $x\in{\left[{0,\pi/2}\right]}$. The function is the same $f$ used before, namely
$f=-\sqrt{b_{1}^{2}\cos^{2}x+b_{2}^{2}\sin^{2}x}~{}.$ (2.96)
From the line element above we see that the branched lens space reduces to the smooth (squashed) lens space if ${n_{+}}={n_{-}}=1$; while it becomes the branched-squashed three-sphere if $n=1$. Eventually, if ${n_{+}}={n_{-}}=n=1$, the branches lens space boils down to a squashed three-sphere.
In the orthonormal frame
$e^{1}$ $=-f\,{\rm d}x~{},\qquad e^{2}=b_{1}b_{2}\,\widetilde{f}^{-1}\sin x\cos x\,{\rm d}\varphi~{},\qquad e^{3}=\frac{\widetilde{f}}{n\,{n_{+}}{n_{-}}}{\left({{A^{(1)}}+{\rm d}\psi}\right)}~{},$
$\widetilde{f}$ $=\sqrt{{\left({{n_{-}}b_{1}}\right)}^{2}\sin^{2}x+{\left({{n_{+}}b_{2}}\right)}^{2}\cos^{2}x}~{},\qquad{A^{(1)}}=n{\left({\,{n_{-}}b_{1}\widetilde{f}^{-1}\sin x}\right)}^{2}{\rm d}\varphi~{},$ (2.97)
the branched lens space manifestly is a $\mathcal{O}{\left({-n\,{n_{+}}{n_{-}}}\right)}$ circle fibration over $\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$ parametrized by ${\left({x,\varphi}\right)}$, indeed
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}x{\rm d}\varphi\sqrt{g_{\mathbb{\Sigma}}}R_{\mathbb{\Sigma}}=\frac{1}{{n_{+}}}+\frac{1}{{n_{-}}}=\chi_{\rm orb}~{},$ $\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{A^{(1)}}=n=\frac{n\,{n_{+}}{n_{-}}}{{n_{+}}{n_{-}}}~{}.$ (2.98)
This geometry admits a Killing vector of the type (2.17), that is
$K={k_{0}}{\left({-\frac{{n_{-}}b_{1}+{n_{+}}b_{2}}{n\,{n_{-}}b_{1}}\partial_{\varphi}+\partial_{\psi}}\right)}~{},$ ${\omega}=-\frac{{n_{-}}b_{1}+{n_{+}}b_{2}}{n\,{n_{-}}b_{1}}~{},$ (2.99)
while its norm reads
${v}={k_{0}}b_{2}/{\left({n\,{n_{-}}}\right)}~{}.$ (2.100)
In turn, the Killing spinors (2.28), with components
$u_{1,2}$ $=\sqrt{\frac{{k_{0}}b_{2}}{2\,n\,{n_{-}}}{\left[{1\pm{\left({{n_{-}}b_{1}+{n_{+}}b_{2}}\right)}\widetilde{f}^{-1}\sin x\cos x}\right]}}~{},$ (2.101)
satisfy the Killing spinor equation for an arbitrary phase $\theta$ if the background fields
$A^{C}$ $={\rm d}\theta-\frac{b_{1}}{2{n_{+}}f}{\rm d}\varphi-\frac{{n_{-}}b_{1}-{n_{+}}b_{2}}{2n\,{n_{+}}{n_{-}}f}{\rm d}\psi~{},$
$V$ $=\frac{b_{1}{\left({1-{\rm i}fH}\right)}\sin^{2}x}{{n_{+}}f}{\rm d}\varphi+\frac{{\left[{{n_{-}}b_{1}-{n_{+}}b_{2}-{\left({{n_{-}}b_{1}+{n_{+}}b_{2}}\right)}\cos{\left({2x}\right)}}\right]}{\left({1-{\rm i}fH}\right)}}{2n\,{n_{+}}{n_{-}}f}{\rm d}\psi~{},$
$A$ $=A^{C}+\frac{3}{2}V~{},$ (2.102)
are turned on. Such background fields are everywhere regular on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ if the phase of the Killing spinors $\theta$ is
$\alpha_{2}=-\frac{1}{{n_{+}}}~{},\qquad\alpha_{3}=\frac{{n_{+}}-{n_{-}}}{n\,{n_{+}}{n_{-}}}~{},\qquad\theta=\frac{1}{2}{\left[{\alpha_{1}(x)+\alpha_{2}\varphi+\alpha_{3}\psi}\right]}~{}.$ (2.103)
3 Localization on $\mathbb{\Sigma}\times S^{1}$ with twist and anti-twist
3.1 BPS locus
We parametrize an Abelian gauge or flavour 1-form field on $\mathbb{\Sigma}\times S^{1}$ as
${\mathcal{A}}={\mathcal{A}}_{\varphi}{\left({x}\right)}{\rm d}\varphi+{\mathcal{A}}_{\psi}{\left({x}\right)}{\rm d}\psi~{},$ ${\mathfrak{f}}_{G}=\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{\mathcal{A}}=\frac{{\mathfrak{m}}}{{n_{+}}{n_{-}}}~{},\qquad\exp{\left({{\rm i}\oint_{S^{1}}{\mathcal{A}}}\right)}={\mathfrak{h}}~{},$ (3.1)
with ${\mathfrak{m}}\in\mathbb{Z}$ and ${\mathfrak{h}}\in U(1)$, where we removed the component ${\mathcal{A}}_{x}{\left({x}\right)}$ along ${\rm d}x$ via a gauge transformation. In the non-Abelian case, we consider a Lie group $G=G_{G}\times G_{F}$ that is the direct product of the gauge group $G_{G}$ and the flavour group $G_{F}$. The 1-forms in the vector multiplet are then $\mathfrak{g}$-valued, where $\mathfrak{g}=\mathfrak{g}_{G}\oplus\mathfrak{g}_{G}$ is the Lie algebra of $G$. In general, fields transform in a representation $\mathfrak{R}_{G}$ of $G$ with weight $\rho=\rho_{G}+\rho_{G}$.
The 1-form in (3.1) is the representative of an $\mathcal{O}{\left({-{\mathfrak{m}}}\right)}$ orbibundle on the spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$ and of the $U(1)$ holonomy group of $S^{1}$. In particular, the component ${\mathcal{A}}_{\varphi}{\left({x}\right)}$ evaluated at the north pole at $x=1$ and at the south pole at $x=-1$ of the spindle reads
${\mathcal{A}}_{\varphi}{\left({+1}\right)}=\frac{{\mathfrak{m}_{+}}}{{n_{+}}}~{},\qquad{\mathcal{A}}_{\varphi}{\left({-1}\right)}=\frac{{\mathfrak{m}_{-}}}{{n_{-}}}~{},$ ${n_{+}}{\mathfrak{m}_{-}}-{n_{-}}{\mathfrak{m}_{+}}={\mathfrak{m}}~{},$ (3.2)
where the integers ${\left({{\mathfrak{m}_{+}},{\mathfrak{m}_{-}}}\right)}$ can be expressed in terms of ${\mathfrak{m}}$ and two other integers ${\left({{\mathfrak{a}_{+}},{\mathfrak{a}_{-}}}\right)}$ satisfying
${\mathfrak{m}_{+}}={\mathfrak{m}}\,{\mathfrak{a}_{+}}~{},\qquad{\mathfrak{m}_{-}}={\mathfrak{m}}\,{\mathfrak{a}_{-}}~{},$ ${n_{+}}{\mathfrak{a}_{-}}-{n_{-}}{\mathfrak{a}_{+}}=1~{}.$ (3.3)
Especially, given a pair ${\left({{\mathfrak{a}_{+}},{\mathfrak{a}_{-}}}\right)}$ satisfying the constraint above, any pair ${\left({{\mathfrak{a}_{+}}+{n_{+}}\delta\mathfrak{a},{\mathfrak{a}_{-}}+{n_{-}}\delta\mathfrak{a}}\right)}$ with $\mathfrak{\delta}\mathfrak{a}\in\mathbb{Z}$ fulfils the constraint too. Physical observables are supposed to be independent of $\delta\mathfrak{a}$.
The 1-form (3.1) is consistent with the gauge fields used in [46] provided that
${\mathcal{A}}^{N}={\mathcal{A}}^{N}_{{\left({0}\right)}}+\frac{\mathfrak{m}_{N}}{n_{N}}{\rm d}\varphi={\mathcal{A}}^{S}_{{\left({0}\right)}}+{\left({{\rm p}+\frac{\mathfrak{m}_{S}}{n_{S}}}\right)}{\rm d}\varphi={\mathcal{A}}^{S}+{\rm p}\,{\rm d}\varphi~{},$ (3.4)
with ${\rm p}\in\mathbb{Z}$ modelling a gauge transformation in the overlap of the two patches $(N,S)$, where the connections ${\left({{\mathcal{A}}^{N},{\mathcal{A}}^{S}}\right)}$ are well defined. The dictionary between our notation and that of [46] is
${\mathcal{A}}^{N}_{{\left({0}\right)}}$ $={\mathcal{A}}-\frac{{\mathfrak{m}_{+}}}{{n_{+}}}{\rm d}\varphi~{},\qquad n_{N}={n_{+}}~{},\qquad\mathfrak{m}_{N}={\mathfrak{m}_{+}}~{},$
${\mathcal{A}}^{S}_{{\left({0}\right)}}$ $={\mathcal{A}}-\frac{{\mathfrak{m}_{-}}}{{n_{-}}}{\rm d}\varphi~{},\qquad n_{S}={n_{-}}~{},\qquad\mathfrak{m}_{S}={\mathfrak{m}_{-}}-{n_{-}}{\rm p}~{}.$ (3.5)
In this language, the flux of ${\mathcal{A}}$ reads [46]
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{\mathcal{A}}={\rm p}+\frac{\mathfrak{m}_{S}}{n_{S}}-\frac{\mathfrak{m}_{N}}{n_{N}}=\frac{{\mathfrak{m}}}{{n_{+}}{n_{-}}}~{},$ (3.6)
in agreement with (3.1). If we add a flat connection ${\left({\beta_{2}{\rm d}\varphi}\right)}$ to ${\mathcal{A}}$ in (3.1) as
${\mathcal{A}}\to{\mathcal{A}}^{\prime}={\left({{\mathcal{A}}_{\varphi}+\beta_{2}}\right)}{\rm d}\varphi+{\mathcal{A}}_{\psi}{\rm d}\psi~{},$ (3.7)
we can make ${\mathcal{A}}$ vanish at the poles of $\mathbb{\Sigma}$ at $x=\pm 1$ by setting $\beta_{2}|_{{\mathcal{U}_{+}}}=-{\mathfrak{m}_{+}}/{n_{+}}$ in the northern patch ${\mathcal{U}_{+}}$ including $x=+1$ and $\beta_{2}|_{{\mathcal{U}_{-}}}=-{\mathfrak{m}_{-}}/{n_{-}}$ in the southern patch ${\mathcal{U}_{-}}$ covering $x=-1$.
The vector-multiplet BPS equations are obtained by setting to zero the gauginos ${\left({\lambda,\widetilde{\lambda}}\right)}$ and their variations ${\left({\delta\lambda,\delta\widetilde{\lambda}}\right)}$, giving
${\Phi_{G}}=\iota_{K}{\mathcal{A}}-{\rm i}{v}\sigma={k_{0}}\varphi_{G}={\rm constant}~{},$
$D+{\frac{{\rm i}}{2}}P^{\mu}\widetilde{P}^{\nu}{\mathcal{F}}_{\mu\nu}+{\frac{{\rm i}}{{v}}}\sigma{\left[{\iota_{K}V-{\rm i}{v}H}\right]}=0~{}.$ (3.8)
We do not impose any reality conditions on fields, for the moment. Inserting the gauge field (3.1) into the BPS equations above yields the BPS values of the scalar $\sigma$ and the auxiliary scalar $D$. Specifically, if ${\mathfrak{s}}=+1$, we have topologically twisted $\mathbb{\Sigma}\times S^{1}$ and the BPS locus
$\sigma|_{\rm BPS}={\rm i}\beta^{-1}{\left[{-\varphi_{G}+{\omega}{\mathcal{A}}_{\varphi}{\left({x}\right)}+{\mathcal{A}}_{\psi}{\left({x}\right)}}\right]}~{},$ $D|_{\rm BPS}=-\frac{{\mathcal{A}}_{\varphi}^{\prime}{\left({x}\right)}}{f\sqrt{1-x^{2}}}~{},$ (3.9)
whereas, if ${\mathfrak{s}}=-1$, we have anti-twisted $\mathbb{\Sigma}\times S^{1}$ and
$\sigma|_{\rm BPS}=-{\rm i}{\left[{-\varphi_{G}+{\omega}{\mathcal{A}}_{\varphi}{\left({x}\right)}+{\mathcal{A}}_{\psi}{\left({x}\right)}}\right]}/{\left({\beta x}\right)}~{},$
$D|_{\rm BPS}=\frac{\beta\,\sigma|_{\rm BPS}+{\left[{\beta+{\rm i}\Omega{\left({1-x^{2}}\right)}}\right]}{\mathcal{A}}_{\varphi}^{\prime}{\left({x}\right)}+{\rm i}{\left({1-x^{2}}\right)}{\mathcal{A}}_{\psi}^{\prime}{\left({x}\right)}}{\beta xf\sqrt{1-x^{2}}}~{}.$ (3.10)
Such supersymmetric values of ${\left({\sigma,D}\right)}$ provide a non-trivial classical contribution to the partition function of the corresponding gauge theory thanks to the presence of (possibly mixed) Chern-Simons and Fayet–Iliopoulos terms, for example. Furthermore, the profiles ${\left({\sigma|_{\rm BPS},D|_{\rm BPS}}\right)}$ implicitly affect the one-loop determinant of supersymmetric fluctuations via the gauge fugacity $\varphi_{G}$ and the flux ${\mathfrak{f}}_{G}$.
Analogously, the BPS locus for the chiral multiplet is found by setting to zero the spinor fields ${\left({\psi,\widetilde{\psi}}\right)}$ together with their supersymmetric variations ${\left({\delta\psi,\delta\widetilde{\psi}}\right)}$. The outcome of this procedure is the following set of BPS equations:
${\left({\mathcal{L}_{K}-{\rm i}r\Phi_{R}-{\rm i}q_{G}\Phi_{G}}\right)}\phi=0~{},$ ${\left({\mathcal{L}_{K}+{\rm i}r\Phi_{R}+{\rm i}q_{G}\Phi_{G}}\right)}\widetilde{\phi}=0~{},$
$F+{\rm i}L_{\widetilde{P}}\phi=0~{},$ $\widetilde{F}+{\rm i}L_{P}\widetilde{\phi}=0~{}.$ (3.11)
For arbitrary values of ${\left({\Phi_{R},\Phi_{G}}\right)}$ the unique solution to the chiral-multiplet BPS equations is
$\phi|_{\rm BPS}=\widetilde{\phi}|_{\rm BPS}=F|_{\rm BPS}=\widetilde{F}|_{\rm BPS}=0~{}.$ (3.12)
As a consequence, we do not expect classical contributions to the partition function to come from F-terms or superpotentials, for instance. In this setup, matter fields affect the theory at the quantum level only, via one-loop determinants.
3.2 Chiral-multiplet one-loop determinants
3.2.1 Unpaired eigenvalues
The chiral-multiplet one-loop determinant entering the partition function can be obtained by exploiting the cancellations between bosonic and fermionic degrees of freedom due to supersymmetry. This procedure is implemented by the formula
$Z^{\rm CM}_{\text{1-L}}=\frac{\det_{\text{Ker}L_{P}}\delta^{2}}{\det_{\text{Ker}L_{\widetilde{P}}}\delta^{2}}=\frac{\det_{\text{Ker}L_{P}}{\left({L_{K}+\mathcal{G}_{\Phi_{G}}}\right)}}{\det_{\text{Ker}L_{\widetilde{P}}}{\left({L_{K}+\mathcal{G}_{\Phi_{G}}}\right)}}~{}.$ (3.13)
Indeed, the operators $L_{P}$ and $L_{\widetilde{P}}$ pair bosonic and fermionic fields in the supersymmetry transformations written in cohomological form. The formula above follows from the fact that the squared supersymmetry variation $\delta^{2}\propto{\left({L_{K}+\mathcal{G}_{\Phi_{G}}}\right)}$ commutes with such pairing operators. Thus, the unpaired eigenvalues surviving the cancellations and contributing to $Z^{\rm CM}_{\text{1-L}}$ are those corresponding to the eigenfunctions spanning the kernels of $L_{P}$ and $L_{\widetilde{P}}$. In particular, the eigenfunctions of $\delta^{2}$ that are annihilated by the differential operator $L_{P}$ have the form
$\mathcal{B}_{m_{\varphi},m_{\psi}}=e^{{\rm i}m_{\varphi}\varphi+{\rm i}m_{\psi}\psi}B_{m_{\varphi},m_{\psi}}{\left({x}\right)}~{},$ $m_{\varphi},m_{\psi}\in\mathbb{Z}~{},$ (3.14)
while the eigenfunctions of $\delta^{2}$ that are annihilated by the operator $L_{\widetilde{P}}$ have the form
$\Phi_{n_{\varphi},n_{\psi}}=e^{{\rm i}n_{\varphi}\varphi+{\rm i}n_{\psi}\psi}\phi_{n_{\varphi},n_{\psi}}{\left({x}\right)}~{},$ $n_{\varphi},n_{\psi}\in\mathbb{Z}~{}.$ (3.15)
The eigenfunctions contributing to $Z^{\rm CM}_{\text{1-L}}$ are those that are non-singular on the whole three-dimensional orbifold. For this reason, we now study the regularity of the eigenfunctions ${\left({\mathcal{B}_{m_{\varphi},m_{\psi}},\Phi_{n_{\varphi},n_{\psi}}}\right)}$ on $\mathbb{\Sigma}\times S^{1}$ in presence of both twist and anti-twist for the R-symmetry. As in [1] we keep the notation compact by defining the following quantities:
$\mathfrak{p}_{+}=q_{G}{\mathfrak{m}_{+}}-{\mathfrak{s}}\frac{r}{2}~{},$ $\mathfrak{p}_{-}=q_{G}{\mathfrak{m}_{-}}+\frac{r}{2}~{},$
$\mathfrak{b}=1+{\mathfrak{s}}{\left\lfloor{{\mathfrak{s}}\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rfloor}+{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}~{},$ $\mathfrak{c}=\frac{{\left\llbracket{-\mathfrak{p}_{+}}\right\rrbracket_{{n_{+}}}}}{{n_{+}}}-{\mathfrak{s}}\frac{{\left\llbracket{{\mathfrak{s}}\,\mathfrak{p}_{-}}\right\rrbracket_{{n_{-}}}}}{{n_{-}}}~{},$
${\gamma_{R}}=-\frac{\alpha_{3}}{2}+\frac{{\omega}}{4}\chi_{-{\mathfrak{s}}}~{},$ ${\gamma_{G}}=-\varphi_{G}+\frac{{\omega}}{2}{\left({\frac{{\mathfrak{m}_{-}}}{{n_{-}}}+\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right)}~{},$
$q=e^{2\pi{\rm i}{\omega}}~{},$ $y=q^{\mathfrak{c}/2}e^{2\pi{\rm i}{\left({r\gamma_{R}+q_{G}\gamma_{G}}\right)}}~{},$ (3.16)
where by ${\left\llbracket{\bullet}\right\rrbracket_{\diamond}}$ we indicate the reminder of the integer division of $\bullet$ by $\diamond$. In particular, ${\left({\mathfrak{b}-1}\right)}$ is the degree of the line orbibundle $\mathcal{O}(-\mathfrak{p})=\mathcal{O}(-{n_{+}}\mathfrak{p}_{-}+{n_{-}}\mathfrak{p}_{+})=\mathcal{O}{\left({-{\left({r/2}\right)}{\left({{n_{+}}+{n_{-}}}\right)}-q_{G}\mathfrak{m}}\right)}$, with $\mathfrak{p}={\left({{n_{+}}\mathfrak{p}_{-}-{n_{-}}\mathfrak{p}_{+}}\right)}$. Also, ${\omega}$, $\gamma_{G}$ and $\gamma_{R}$ are effective fugacities for angular momentum, gauge and R-symmetry, respectively. The objects written in (3.2.1) are valid for a theory invariant under a $U(1)$ gauge or flavour group. For a general, possibly non-Abelian, Lie group $G$, the substitutions
$\mathfrak{m}_{\pm}\to\rho{\left({\mathfrak{m}_{\pm}}\right)}~{},$ ${\gamma_{G}}\to\rho{\left({{\gamma_{G}}}\right)}~{},$ (3.17)
are in order, with $\rho$ being the weight of the representation $\mathfrak{R}_{G}$ that we defined previously.
Topological twist.
In the case of the topologically twisted $\mathbb{\Sigma}\times S^{1}$ the behaviour of $B_{m_{\varphi},m_{\psi}}{\left({x}\right)}$ near the north pole of the spindle is
$\lim_{x\to+1}B_{m_{\varphi},m_{\psi}}{\left({x}\right)}\to{\left({1-x}\right)}^{{\left[{2m_{\varphi}{n_{+}}+{\left({r-2}\right)}-2q_{G}{\mathfrak{m}_{+}}}\right]}/4}~{},$ (3.18)
which is non-singular if
$m_{\varphi}\geq{\left\lceil{\frac{1-{\left({r/2}\right)}+q_{G}{\mathfrak{m}_{+}}}{{n_{-}}}}\right\rceil}={\left\lceil{\frac{1+\mathfrak{p}_{+}}{{n_{-}}}}\right\rceil}=1+{\left\lfloor{\frac{\mathfrak{p}_{+}}{{n_{-}}}}\right\rfloor}~{},$ (3.19)
[…]
$\lim_{x\to-1}B_{m_{\varphi},m_{\psi}}{\left({x}\right)}\to{\left({1+x}\right)}^{{\left[{-2m_{\varphi}{n_{-}}+{\left({r-2}\right)}+2q_{G}{\mathfrak{m}_{-}}}\right]}/4}~{},$ (3.20)
which is non-singular if
$m_{\varphi}\leq{\left\lfloor{\frac{{\left({r/2}\right)}-1+q_{G}{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rfloor}={\left\lfloor{\frac{-1+\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}=-1+{\left\lceil{\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rceil}=-1-{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}~{}.$ (3.21)
Analogously, the modes $\phi_{n_{\varphi},n_{\psi}}{\left({x}\right)}$ behave near the north pole of the spindle as
$\lim_{x\to+1}\phi_{n_{\varphi},n_{\psi}}{\left({x}\right)}\to{\left({1-x}\right)}^{{\left({-2n_{\varphi}{n_{+}}-r+2q_{G}{\mathfrak{m}_{+}}}\right)}/4}~{},$ (3.22)
which is non-singular if
$n_{\varphi}\leq{\left\lfloor{\frac{-{\left({r/2}\right)}+q_{G}{\mathfrak{m}_{+}}}{{n_{-}}}}\right\rfloor}={\left\lfloor{\frac{\mathfrak{p}_{+}}{{n_{-}}}}\right\rfloor}~{},$ (3.23)
while at the south pole of the spindle we find
$\lim_{x\to-1}\phi_{n_{\varphi},n_{\psi}}{\left({x}\right)}\to{\left({1+x}\right)}^{{\left({2n_{\varphi}{n_{-}}-r-2q_{G}{\mathfrak{m}_{-}}}\right)}/4}~{},$ (3.24)
which is non-singular if
$n_{\varphi}\geq{\left\lceil{\frac{{\left({r/2}\right)}+q_{G}{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rceil}={\left\lceil{\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rceil}=-{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}~{}.$ (3.25)
The inequalities above imply that there are no regular functions in either ${\rm Ker}L_{\widetilde{P}}$ or ${\rm Ker}L_{P}$ if $\mathfrak{b}=0$; thus, $Z^{\rm CM}_{\text{1-L}}|_{\mathfrak{b}=0}=1$. If $\mathfrak{b}\leq-1$, there are no regular fields in the kernel of $L_{\widetilde{P}}$ and the eigenfunctions $\Phi_{n_{\varphi},n_{\psi}}$ do not contribute to the one-loop determinant of the chiral-multiplet partition function. Nonetheless, the eigenfunctions $\mathcal{B}_{m_{\varphi},m_{\psi}}$ are non-singular if $\mathfrak{b}\leq-1$ and do contribute to $Z^{\rm CM}_{\text{1-L}}$ via
$Z^{\mathcal{B}}_{\text{1-L}}$ $=\prod_{m_{\psi}\in\mathbb{Z}}\prod_{m_{\varphi}=1+{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}^{-1-{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}{\left[{m_{\psi}+{\omega}m_{\varphi}+{\left({2-r}\right)}\frac{\alpha_{3}}{2}-q_{G}\varphi_{G}}\right]}~{},$
$=\prod_{m_{\psi}\in\mathbb{Z}}\prod_{j=0}^{-1-\mathfrak{b}}{\left[{m_{\psi}+{\omega}{\left({j+\frac{1+\mathfrak{b}+\mathfrak{c}}{2}}\right)}+r\gamma_{R}+q_{G}\gamma_{G}}\right]}~{}.$ (3.26)
We can regularize in a single step the infinite product above by writing
$Z^{\mathcal{B}}_{\text{1-L}}$ $=\prod_{m_{\psi}\in\mathbb{Z}}\prod_{j\in\mathbb{N}}\frac{m_{\psi}+{\omega}{\left({j+\frac{1+\mathfrak{b}+\mathfrak{c}}{2}}\right)}+r\gamma_{R}+q_{G}\gamma_{G}}{m_{\psi}+{\omega}{\left({j+\frac{1-\mathfrak{b}+\mathfrak{c}}{2}}\right)}+r\gamma_{R}+q_{G}\gamma_{G}}~{},$
$={\left({-y}\right)}^{\mathfrak{b}/2}\frac{{\left({q^{{\left({1+\mathfrak{b}}\right)}/2}y;q}\right)}_{\infty}}{{\left({q^{{\left({1-\mathfrak{b}}\right)}/2}y;q}\right)}_{\infty}}=\frac{{\left({-y}\right)}^{\mathfrak{b}/2}}{{\left({q^{{\left({1-\mathfrak{b}}\right)}/2}y;q}\right)}_{\mathfrak{b}}}=\frac{{\left({-y^{-1}}\right)}^{\mathfrak{b}/2}}{{\left({q^{{\left({1-\mathfrak{b}}\right)}/2}y^{-1};q}\right)}_{\mathfrak{b}}}~{}.$ (3.27)
On the other hand, If $\mathfrak{b}\geq 1$, the situation is reversed as there are no regular fields in the kernel of $L_{P}$ and $\mathcal{B}_{m_{\varphi},m_{\psi}}$ do not contribute to $Z^{\rm CM}_{\text{1-L}}$; while the eigenfunctions $\Phi_{n_{\varphi},n_{\psi}}$ yield a factor of
$Z^{\Phi}_{\text{1-L}}$ $=\prod_{n_{\psi}\in\mathbb{Z}}\prod_{n_{\varphi}=-{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}^{{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}{\left({n_{\psi}+{\omega}n_{\varphi}-r\frac{\alpha_{3}}{2}-q_{G}\varphi_{G}}\right)}^{-1}~{},$
$=\prod_{n_{\psi}\in\mathbb{Z}}\prod_{j=0}^{\mathfrak{b}-1}{\left[{n_{\psi}+{\omega}{\left({j+\frac{1-\mathfrak{b}+\mathfrak{c}}{2}}\right)}+r\gamma_{R}+q_{G}\gamma_{G}}\right]}^{-1}~{},$
$=\prod_{n_{\psi}\in\mathbb{Z}}\prod_{j\in\mathbb{N}}\frac{n_{\psi}+{\omega}{\left({j+\frac{1+\mathfrak{b}+\mathfrak{c}}{2}}\right)}+r\gamma_{R}+q_{G}\gamma_{G}}{j_{\psi}+{\omega}{\left({j+\frac{1-\mathfrak{b}+\mathfrak{c}}{2}}\right)}+r\gamma_{R}+q_{G}\gamma_{G}}=\frac{{\left({-y^{-1}}\right)}^{\mathfrak{b}/2}}{{\left({q^{{\left({1-\mathfrak{b}}\right)}/2}y^{-1};q}\right)}_{\mathfrak{b}}}~{},$ (3.28)
which is consistent with the result obtained by considering the eigenfunctions $\mathcal{B}_{m_{\varphi},m_{\psi}}$. We then conclude that, on topologically twisted $\mathbb{\Sigma}\times S^{1}$, the one-loop determinant for a chiral multiplet with weight $\rho$ in a representation $\mathfrak{R}_{G}$ of a gauge group $G$ is
$Z^{\rm CM}_{\text{1-L}}$ $=\prod_{\rho\in\mathfrak{R}_{G}}\frac{{\left({-y^{\rho}}\right)}^{\mathfrak{b}/2}}{{\left({q^{{\left({1-\mathfrak{b}}\right)}/2}y^{\rho};q}\right)}_{\mathfrak{b}}}=\prod_{\rho\in\mathfrak{R}_{G}}\frac{{\left({-y^{-\rho}}\right)}^{\mathfrak{b}/2}}{{\left({q^{{\left({1-\mathfrak{b}}\right)}/2}y^{-\rho};q}\right)}_{\mathfrak{b}}}~{},$ (3.29)
as in [1]. The property ${\left({\bullet;q}\right)}_{0}=1$ enjoyed by the finite q-Pochhammer symbol ensures that $Z^{\rm CM}_{\text{1-L}}|_{\mathfrak{b}=0}=1$, as anticipated. In the degenerate case ${n_{+}}={n_{-}}=1$ the spindle becomes a smooth sphere and the simplifications
$\mathfrak{b}=1-q_{G}\mathfrak{m}-r=1-q_{G}{\mathfrak{f}}_{G}-r{\mathfrak{f}}_{R}~{},$ $\mathfrak{c}=0~{},$
$2{\gamma_{R}}=-\alpha_{3}\in\mathbb{Z}~{},$ $y={\left({-1}\right)}^{r\alpha_{3}}e^{2\pi{\rm i}q_{G}{\gamma_{G}}}~{},$ (3.30)
occurr. In turn, (3.29) becomes identical to the one-loop determinant for a chiral multiplet on topologically twisted $S^{2}\times S^{1}$ computed in [35], where $\alpha_{3}=0$ was assumed.
The one-loop determinant (3.29) can also be interpreted as the partition function of a weakly gauged chiral multiplet, namely a chiral multiplet coupled to a non-dynamical background vector multiplet. Such a partition function is not invariant under large gauge transformations shifting the effective fugacity ${\gamma_{G}}$ by an integer $n$:
$y\to e^{2\pi{\rm i}\,n}y\qquad\implies\qquad Z^{\rm CM}_{\text{1-L}}\to{\left({-1}\right)}^{n\,\mathfrak{b}}Z^{\rm CM}_{\text{1-L}}~{}.$ (3.31)
This multi-valuedness of $Z^{\rm CM}_{\text{1-L}}$ is usually associated to a parity anomaly affecting the matter sector of the theory. Such an anomaly can be cured by considering the classical contribution $Z^{\rm CS}_{\rm eff}$ of an effective Chern-Simons term with half-integer level, whose anomaly cancels that of $Z^{\rm CM}_{\text{1-L}}$. In the case of $S^{2}\times S^{1}$ the counterterm needed is $Z^{\rm CS}_{\rm eff}=y^{\pm{\mathfrak{f}}_{G}/2}=y^{\pm\mathfrak{m}/2}$, while for $Z^{\rm CM}_{\text{1-L}}$ on $\mathbb{\Sigma}\times S^{1}$ it is
$Z^{\rm CS}_{\rm eff}=y^{\pm\mathfrak{b}/2}~{}.$ (3.32)
We discuss the differences between effective and canonical Chern-Simons terms in Section 3.3, where we compute the relevant classical contributions to the partition function of a gauge theory on $\mathbb{\Sigma}\times S^{1}$.
Anti-twist.
In the case of the anti-twisted $\mathbb{\Sigma}\times S^{1}$ the eigenfunction $B_{m_{\varphi},m_{\psi}}{\left({x}\right)}$ is non-singular near the north pole of the spindle if
$m_{\varphi}\leq{\left\lfloor{\frac{q_{G}{\mathfrak{m}_{+}}+{\left({r/2}\right)}-1}{{n_{+}}}}\right\rfloor}={\left\lfloor{\frac{\mathfrak{p}_{+}-1}{{n_{+}}}}\right\rfloor}=-1+{\left\lceil{\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rceil}=-1-{\left\lfloor{-\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rfloor}~{},$ (3.33)
while it is non-singular near the south pole of the spindle if
$m_{\varphi}\leq{\left\lfloor{\frac{q_{G}{\mathfrak{m}_{-}}+{\left({r/2}\right)}-1}{{n_{-}}}}\right\rfloor}={\left\lfloor{\frac{\mathfrak{p}_{-}-1}{{n_{-}}}}\right\rfloor}=-1+{\left\lceil{\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rceil}=-1-{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}~{},$ (3.34)
with the two inequalities above being both true if
$m_{\varphi}\leq-1-\frac{1}{2}{\left({{\left\lfloor{-\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rfloor}+{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}}\right)}-\frac{1}{2}\left|{\left\lfloor{-\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rfloor}-{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}\right|~{},$ (3.35)
where we employed the identity
$\min{\left({x_{1},x_{2}}\right)}=\frac{x_{1}+x_{2}}{2}-\left|\frac{x_{1}-x_{2}}{2}\right|~{}.$ (3.36)
Analogously, the fields $\phi_{n_{\varphi},n_{\psi}}{\left({x}\right)}$ are non-singular near the north pole of the spindle if
$n_{\varphi}\geq{\left\lceil{\frac{r+q_{G}{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rceil}={\left\lceil{\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rceil}=-{\left\lfloor{-\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rfloor}~{},$ (3.37)
while they are non-singular at the south pole of the spindle if
$n_{\varphi}\geq{\left\lceil{\frac{r+q_{G}{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rceil}={\left\lceil{\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rceil}=-{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}~{},$ (3.38)
which are constraints that are both satisfied if
$n_{\varphi}\geq-\frac{1}{2}{\left({{\left\lfloor{-\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rfloor}+{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}}\right)}+\frac{1}{2}\left|{\left\lfloor{-\frac{\mathfrak{p}_{+}}{{n_{+}}}}\right\rfloor}-{\left\lfloor{-\frac{\mathfrak{p}_{-}}{{n_{-}}}}\right\rfloor}\right|~{},$ (3.39)
where we used the formula
$\max{\left({x_{1},x_{2}}\right)}=\frac{x_{1}+x_{2}}{2}+\left|\frac{x_{1}-x_{2}}{2}\right|~{}.$ (3.40)
Therefore, the eigenfunctions $\mathcal{B}_{m_{\varphi},m_{\psi}}$ and $\Phi_{n_{\varphi},n_{\psi}}$ both contribute to the one-loop determinant of the chiral-multiplet partition function via the infinite product
$Z^{\mathcal{B},\Phi}_{\text{1-L}}$ $=\prod_{m_{\psi}\in\mathbb{Z}}\prod_{j\in\mathbb{N}}\frac{m_{\psi}+{\omega}{\left({j+1+\frac{|1-\mathfrak{b}|-\mathfrak{c}}{2}}\right)}-r\gamma_{R}-q_{G}\gamma_{G}-\alpha_{3}}{m_{\psi}+{\omega}{\left({j+\frac{|1-\mathfrak{b}|+\mathfrak{c}}{2}}\right)}+r\gamma_{R}+q_{G}\gamma_{G}}~{},$
$=e^{2\pi{\rm i}\Psi_{\rm AT}}\frac{{\left({q^{1+\frac{1}{2}|1-\mathfrak{b}|}y^{-1};q}\right)}_{\infty}}{{\left({q^{\frac{1}{2}|1-\mathfrak{b}|}y;q}\right)}_{\infty}}~{},$ (3.41)
where the overall phase reads
$\Psi_{\rm AT}=\frac{1}{4{\omega}}{\left({1+\alpha_{3}-{\omega}|1-\mathfrak{b}|}\right)}{\left[{\alpha_{3}+2q_{G}\gamma_{G}+r\gamma_{R}+{\omega}{\left({\mathfrak{c}-1}\right)}}\right]}~{}.$ (3.42)
If we choose $\alpha_{3}=-1$, corresponding to anti-periodic Killing spinors on the spindle, then the phase factor simplifies and the one-loop determinant becomes
$Z^{\mathcal{B},\Phi}_{\text{1-L}}={\left({-y^{-1}\sqrt{q}}\right)}^{\frac{1}{2}|1-\mathfrak{b}|}\frac{{\left({q^{1+\frac{1}{2}|1-\mathfrak{b}|}y^{-1};q}\right)}_{\infty}}{{\left({q^{\frac{1}{2}|1-\mathfrak{b}|}y;q}\right)}_{\infty}}~{}.$ (3.43)
Furthermore, if ${n_{+}}={n_{-}}=1$, then
$\mathfrak{b}=1-q_{G}\mathfrak{m}=1-q_{G}{\mathfrak{f}}_{G}~{},$ $\mathfrak{c}=0~{},$
$2{\gamma_{R}}=-\alpha_{3}+{\omega}~{},$ $y={\left({-1}\right)}^{r\alpha_{3}}q^{r/2}\,e^{2\pi{\rm i}q_{G}{\gamma_{G}}}~{},$ (3.44)
and up to overall factors we recover the generalized superconformal index for a chiral multiplet [39],
$Z^{\mathcal{B},\Phi}_{\text{1-L}}|_{n_{\pm}=1}$ $={\left({{\left({-1}\right)}^{1-r}q^{\frac{1-r}{2}}e^{-2\pi{\rm i}q_{G}{\gamma_{G}}}}\right)}^{|q_{G}\mathfrak{m}|/2}\frac{{\left({{\left({-1}\right)}^{r}q^{\frac{1}{2}|q_{G}\mathfrak{m}|+1-\frac{r}{2}}e^{-2\pi{\rm i}q_{G}{\gamma_{G}}};q}\right)}_{\infty}}{{\left({{\left({-1}\right)}^{r}q^{\frac{1}{2}|q_{G}\mathfrak{m}|+\frac{r}{2}}e^{2\pi{\rm i}q_{G}{\gamma_{G}}};q}\right)}_{\infty}}~{}.$ (3.45)
In [47] it was pointed out that the appearence of absolute values of fluxes into the index is spurious, meaning that the replacements
$|q_{G}\mathfrak{m}|\to-q_{G}\mathfrak{m}~{},$ $|1-\mathfrak{b}|\to{\left({\mathfrak{b}-1}\right)}~{},$ (3.46)
should be applied to $Z^{\rm CM}_{\text{1-L}}$. The described substitutions correspond to the absence of regularity constraints on the eigenfunctions $\mathcal{B}_{m\sigma,m_{\psi}}$ and $\Phi_{n_{\sigma},n_{\psi}}$ at the north and south pole of the spindle. This approach becomes apparent when computing $Z^{\rm CM}_{\text{1-L}}$ through the index theorem, wherein calculations are performed separately in patches on $\mathbb{\Sigma}$. Notably, this method avoids generating absolute values in the one-loop determinants as demonstrated e.g. in [48].
In summary, the one-loop determinant for a chiral multiplet with weight $\rho$ in a representation $\mathfrak{R}_{G}$ of a gauge group $G$ on anti-twisted $\mathbb{\Sigma}\times S^{1}$, reads
$Z^{\rm CM}_{\text{1-L}}=\prod_{\rho\in\mathfrak{R}_{G}}{\left({-y^{-\rho}\sqrt{q}}\right)}^{\frac{1}{2}{\left({\mathfrak{b}-1}\right)}}\frac{{\left({q^{\frac{1}{2}{\left({\mathfrak{b}+1}\right)}}y^{-\rho};q}\right)}_{\infty}}{{\left({q^{\frac{1}{2}{\left({\mathfrak{b}-1}\right)}}y^{\rho};q}\right)}_{\infty}}~{}.$ (3.47)
Finally, the identities enjoyed by q-Pochhammer symbols allows for the derivation of a unique expression for $Z^{\rm CM}_{\text{1-L}}$ valid for both topologically twisted and anti-twisted $\mathbb{\Sigma}\times S^{1}$:
$Z^{\rm CM}_{\text{1-L}}=\prod_{\rho\in\mathfrak{R}_{G}}{\left({-y^{-\rho}}\right)}^{\frac{1}{4}{\left({2\mathfrak{b}-1+{\mathfrak{s}}}\right)}}q^{\frac{1}{8}{\left({1-{\mathfrak{s}}}\right)}{\left({\mathfrak{b}-1}\right)}}\frac{{\left({q^{\frac{1}{2}{\left({\mathfrak{b}+1}\right)}}y^{-\rho};q}\right)}_{\infty}}{{\left({q^{\frac{{\mathfrak{s}}}{2}{\left({\mathfrak{b}-1}\right)}}y^{-{\mathfrak{s}}\rho};q}\right)}_{\infty}}~{},$ (3.48)
which for $G=U(1)$ coincides with the expression reported in [1]. The one-loop determinant of a vector multiplet in a gauge group $G$ can be obtained from (3.48) after setting $r=2$ and $\mathfrak{R}_{G}={\rm adj}\,G$.
3.2.2 Equivariant orbifold index theorem
One-loop determinants of supersymmetric partition functions can also be obtained from equivariant index theory [2]. For instance, in the case of a chiral multiplet on a manifold $\mathscr{M}$ and of charge $r$ with respect to the $U(1)_{R}$ R-symmetry and charge $q_{G}$ with respect to a gauge or flavour group $U(1)_{G}$, the eigenvalues contributing to $Z^{\rm CM}_{\text{1-L}}$ are encoded by the formula
${\rm ind}{\left({L_{\widetilde{P}};\widehat{g}}\right)}=\sum_{p\in\mathscr{M}^{\widehat{g}}}\frac{{\rm tr}_{\,\Gamma_{0}}\widehat{g}-{\rm tr}_{\,\Gamma_{1}}\widehat{g}}{\det{\left({1-J_{p}}\right)}}~{},$ (3.49)
which is the index of the operator $L_{\widetilde{P}}$ with respect to the action of the group element $\widehat{g}=\exp{\left({-{\rm i}{\epsilon}\delta^{2}}\right)}$ induced by the square of the supersymmetry variation $\delta^{2}$ and tuned by the equivariant parameter ${\epsilon}$. The summation in (3.49) spans across the fixed-point set $\mathscr{M}^{\widehat{g}}$ of the transformation $\widehat{g}$ on the manifold $\mathscr{M}$. At each fixed point $p\in\mathscr{M}$, the group element $\widehat{g}$ induces an action on the manifold $\mathscr{M}$, thereby effecting a mapping from the coordinate $z_{p}$ to $z^{\prime}_{p}$. This transformation generates a non-trivial Jacobian denoted as $J_{p}=\partial z^{\prime}_{p}/\partial z_{p}$, appearing in the denominator of (3.49). In addition, $\widehat{g}$ acts upon both $\Gamma_{0}$ and $\Gamma_{1}$, where $\Gamma_{0}$ represents the space of sections of the $U(1)_{R}\times U(1)_{G}$-valued line bundle ${\rm L}$ over $\mathscr{M}$, while $\Gamma_{1}$ is the image of $\Gamma_{0}$ under by $L_{\widetilde{P}}$. These combined contributions yield the numerator of (3.49). In the case of our interest $\mathscr{M}=\mathbb{\Sigma}\times S^{1}$ and the neighborhoods $\mathcal{U}_{p}$ covering the fixed points $p\in\mathscr{M}^{\widehat{g}}$ are isomorphic to $\mathbb{C}/\mathbb{Z}_{n_{p}}$. In order for the computation to accomodate such new structures, the orbifold version of (3.49) is required. The fixed point formula in the framework of the equivariant orbifold index theorem is expressed as follows:
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}=\sum_{p\in\mathscr{M}^{\widehat{g}}}\frac{1}{n_{p}}\sum_{\widehat{w}\in\mathbb{Z}_{n_{p}}}\frac{{\rm tr}_{\,\Gamma_{0}}{\left({\widehat{w}\,\widehat{g}}\right)}-{\rm tr}_{\,\Gamma_{1}}{\left({\widehat{w}\,\widehat{g}}\right)}}{\det{\left({1-W_{p}J_{p}}\right)}}~{},$ (3.50)
where $W_{p}$ is the Jacobian resulting from the action of $\widehat{w}$ on the coordinates $z_{p}$ within $\mathcal{U}_{p}$. The elements $\widehat{w}\in\mathbb{Z}_{p}$ act on both sections and coordinates akin to the action of $\widehat{g}$, but their action is weighted by the roots of unity $w_{p}^{\ell}=e^{2\pi{\rm i}\,\ell/n_{p}}$, with $\ell=0,\dots,{\left({n_{p}-1}\right)}$, rather than by the equivariant parameter ${\epsilon}$. Specifically, on $\mathbb{\Sigma}\times S^{1}$ we have
$\delta^{2}$ $=-2{\rm i}{\left({L_{K}+\mathcal{G}_{\Phi_{G}}}\right)}=-2{\rm i}{\left({\mathcal{L}_{K}-{\rm i}q_{R}\Phi_{R}-{\rm i}q_{G}\Phi_{G}}\right)}~{},$
$\mathcal{L}_{K}$ $={k_{0}}{\left({{\omega}\partial_{\varphi}+\partial_{\psi}}\right)}~{},$ (3.51)
where ${\left({\Phi_{R},\Phi_{G}}\right)}$ are to be evaluated on the BPS locus. The operator $\widehat{g}$ factorizes as $\widehat{g}=\widehat{g}_{\mathbb{\Sigma}}\widehat{g}_{S^{1}}$, with
$\widehat{g}_{\mathbb{\Sigma}}$ $=\exp{\left[{-2t{\left({{k_{0}}{\omega}\partial_{\varphi}-{\rm i}q_{R}\Phi_{R}-{\rm i}q_{G}\Phi_{G}}\right)}}\right]}~{},$
$\widehat{g}_{S^{1}}$ $=\exp{\left({-2t{k_{0}}\partial_{\psi}}\right)}~{},$ (3.52)
being group actions corresponding to the spindle $\mathbb{\Sigma}$ and to the circle $S^{1}$, respectively. In particular, the fixed-point set $\mathscr{M}^{\widehat{g}_{S^{1}}}$ is empty as $\widehat{g}_{S^{1}}$ acts freely on $S^{1}$. Instead, $\mathscr{M}^{\widehat{g}_{\mathbb{\Sigma}}}$ is non-trivial and it contains the poles of the spindle at $x=\pm 1$. The actions of $\widehat{g}_{\mathbb{\Sigma}}$ and $\widehat{w}$ upon the complex coordinates defined in (2.36) and (2.38) is
$\widehat{g}_{\mathbb{\Sigma}}z_{+}=q_{+}z_{+}~{},$ $q_{+}=\exp{\left({2{\rm i}{\epsilon}{k_{0}}{\omega}/{n_{+}}}\right)}~{},$
$\widehat{g}_{\mathbb{\Sigma}}z_{-}=q_{-}^{-1}z_{-}~{},$ $q_{-}=\exp{\left({2{\rm i}{\epsilon}{k_{0}}{\omega}/{n_{-}}}\right)}~{},$
$\widehat{w}_{+}z_{+}=w_{+}z_{+}~{},$ $w_{+}=\exp{\left({2\pi{\rm i}/{n_{+}}}\right)}~{},$
$\widehat{w}_{-}z_{-}=w_{-}^{-1}z_{-}~{},$ $w_{-}=\exp{\left({2\pi{\rm i}/{n_{-}}}\right)}~{}.$ (3.53)
The form of $L_{\widetilde{P}}$ as well as the action of $\widehat{g}_{\mathbb{\Sigma}}$ and $\widehat{w}$ on sections in $\Gamma_{0}$ and $\Gamma_{1}$ exhibit slight differences for twisted and anti-twisted spindles. Consequently, we analyze these two cases separately.
Topological twist.
In computing the contribution to the index formula (3.50) coming from the north and south pole of the spindle we work with 1-form fields whose components on $\mathbb{\Sigma}$ vanish at the origin of the patches ${\mathcal{U}_{+}}$ and ${\mathcal{U}_{-}}$. This makes the calculation simple as $L_{\widetilde{P}}$ becomes a pure differential operator, for instance. Especially, the component $\varphi$ of the R-symmetry field vanishes at the origin of ${\mathcal{U}_{+}}$ and ${\mathcal{U}_{-}}$ if
$\alpha_{2}|_{{\mathcal{U}_{+}}}=\frac{1}{{n_{+}}}~{},$ $\alpha_{2}|_{{\mathcal{U}_{-}}}=-\frac{1}{{n_{-}}}~{}.$ (3.54)
Analogously, the component ${\mathcal{A}}_{\varphi}$ of the flavour field (3.7) vanishes at $x=\pm 1$ if
$\beta_{2}|_{{\mathcal{U}_{+}}}=-\frac{{\mathfrak{m}_{+}}}{{n_{+}}}~{},$ $\beta_{2}|_{{\mathcal{U}_{-}}}=-\frac{{\mathfrak{m}_{-}}}{{n_{-}}}~{},$ (3.55)
in the patches ${\mathcal{U}_{+}}$ and ${\mathcal{U}_{-}}$ , respectively. In this setting,
$L_{\widetilde{P}}|_{{\mathcal{U}_{+}}}=-2{\rm i}\partial_{+}~{},$ $L_{\widetilde{P}}|_{{\mathcal{U}_{-}}}=2{\rm i}\partial_{-}~{},$ (3.56)
where $\overline{\partial}_{\pm}$ is the complex conjugate of ±$\partial_{\pm}=\partial_{z_{\pm}}$. In particular, the operators $L_{\widetilde{P}}|_{{\mathcal{U}_{+}}}$ and $L_{\widetilde{P}}|_{{\mathcal{U}_{-}}}$ are transversally elliptic with respect to the free action of $\widehat{g}_{S^{1}}\in U(1)$. Then, according to [49], the index of $L_{\widetilde{P}}|_{{\mathcal{U}_{+}}}$ on $\mathbb{\Sigma}\times S^{1}$ is given by the weighted sum of the index on the quotient ${\left({\mathbb{\Sigma}\times S^{1}}\right)}/S^{1}\cong\mathbb{\Sigma}$ over all irreducible representations of $U(1)$, where the weights are the characters $\chi_{\mathfrak{R}_{U(1)}}{\left({\widehat{g}_{S^{1}}}\right)}=\chi_{n}{\left({\widehat{g}_{S^{1}}}\right)}=e^{-2{\rm i}{\epsilon}{k_{0}}n}$ with $n$ being the label of the representative $\phi_{n}\in\Gamma_{0}$, where the latter is the space of sections of the line orbibundle
${\rm L}=\mathcal{O}(-\mathfrak{p})=\mathcal{O}(-\mathfrak{p})=\mathcal{O}(-{n_{+}}\mathfrak{p}_{-}+{n_{-}}\mathfrak{p}_{+})=\mathcal{O}{\left({-{\left({r/2}\right)}{\left({{n_{+}}+{n_{-}}}\right)}-q_{G}\mathfrak{m}}\right)}~{}.$ (3.57)
The operator $\widehat{g}_{\mathbb{\Sigma}}$ acts as
$\widehat{g}_{\mathbb{\Sigma}}|_{{\mathcal{U}_{+}}}\phi_{n}$ $=\xi\,q_{+}^{-q_{G}{\mathfrak{m}_{+}}+{\left({r/2}\right)}}\phi_{n}=\xi\,q_{+}^{-\mathfrak{p}_{+}}\phi_{n}~{},$
$\widehat{g}_{\mathbb{\Sigma}}|_{{\mathcal{U}_{-}}}\phi_{n}$ $=\xi\,q_{-}^{-q_{G}{\mathfrak{m}_{-}}-{\left({r/2}\right)}}\phi_{n}=\xi\,q_{-}^{-\mathfrak{p}_{-}}\phi_{n}~{},\qquad\xi=e^{2{\rm i}{\epsilon}{k_{0}}{\left[{{\left({r/2}\right)}\alpha_{3}+q_{G}\varphi_{G}}\right]}}~{}.$ (3.58)
Moreover, the action of $\widehat{g}_{\mathbb{\Sigma}}$ on sections in the image $\Gamma_{1}$ reads
$\widehat{g}_{\mathbb{\Sigma}}|_{{\mathcal{U}_{+}}}\partial_{+}\phi_{n}$ $=\xi\,q_{+}^{-\mathfrak{p}_{+}-1}\phi_{n}~{},$
$\widehat{g}_{\mathbb{\Sigma}}|_{{\mathcal{U}_{-}}}\partial_{-}\phi_{n}$ $=\xi\,q_{-}^{-\mathfrak{p}_{-}+1}\phi_{n}~{}.$ (3.59)
Combining all these ingredients in the fixed point formula (3.50) yields
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}$ $=\sum_{n\in\mathbb{Z}}{\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}\chi_{n}{\left({\widehat{g}_{S^{1}}}\right)}\xi~{},$
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}$ $=\frac{1}{{n_{+}}}\sum_{j=0}^{{n_{+}}-1}\frac{w_{+}^{-j\mathfrak{p}_{+}}q_{+}^{-\mathfrak{p}_{+}}}{1-w_{+}^{j}q_{+}}+\frac{1}{{n_{-}}}\sum_{j=0}^{{n_{-}}-1}\frac{w_{-}^{-j\mathfrak{p}_{-}}q_{-}^{-\mathfrak{p}_{-}}}{1-w_{-}^{-j}q_{-}^{-1}}~{},$ (3.60)
where we factored out the flavour fugacity $\xi$. The identities
$\frac{1}{n_{p}}\sum_{j=0}^{n_{p}-1}\frac{w_{p}^{j\nu_{p}}}{1-w_{p}^{j}\,q_{p}}=\frac{1}{n_{p}}\sum_{j=0}^{n_{p}-1}\frac{w_{p}^{-j\nu_{p}}}{1-w_{p}^{-j}\,q_{p}}=\frac{q_{p}^{{\left\llbracket{-\nu_{p}}\right\rrbracket_{n_{p}}}}}{1-q_{p}^{n_{p}}}~{},$ $\frac{\bullet}{n_{p}}={\left\lfloor{\frac{\bullet}{n_{p}}}\right\rfloor}+\frac{{\left\llbracket{\bullet}\right\rrbracket_{n_{p}}}}{n_{p}}~{},$ (3.61)
imply
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}$ $=\sum_{n\in\mathbb{Z}}{\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}\chi_{n}{\left({\widehat{g}_{S^{1}}}\right)}\xi~{},$
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}$ $=\frac{{\rm q}^{-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}}{1-{\rm q}}+\frac{{\rm q}^{{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}}{1-{\rm q}^{-1}}=\frac{{\rm q}^{-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}-{\rm q}^{1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}}{1-{\rm q}}~{},$ (3.62)
with ${\rm q}=q_{+}^{n_{+}}=q_{-}^{n_{-}}$. If
$-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}=1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}\iff\mathfrak{b}=0~{},$ (3.63)
then ${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}=0$, where $\mathfrak{b}={\left({1+\deg{\rm L}}\right)}$ defined in (3.2.1) naturally appeared. Instead, if
$-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}<1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}\iff\mathfrak{b}\geq 1~{},$ (3.64)
then
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}={\rm q}^{-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}\frac{1-{\rm q}^{\mathfrak{b}}}{1-{\rm q}}={\rm q}^{-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}{\left({1+{\rm q}+\dots+{\rm q}^{\mathfrak{b}-1}}\right)}~{},$ (3.65)
where the overall factor ${\rm q}^{-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}$ is the lift of the equivariant action [50] at the north pole of the spindle. In fact, given that $L_{\widetilde{P}}$ equals the Dolbeault operator $\partial$ in the vicinity of any fixed point, ${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}$ represents the equivariant orbifold index of $\partial$ on $\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$ in presence of the line bundle ${\rm L}$. This index assigns a monomial ${\rm q}^{\ell}$ to each element within the basis of anti-holomorphic sections of ${\rm L}$ in the kernel of $\partial$, where $\ell\in\mathbb{N}$, ${\rm q}^{0}=1$ corresponds to the constant section and the polynomial’s degree is ${\left({\mathfrak{b}-1}\right)}=\deg{\rm L}$. This generalizes what occurs in the case of the equivariant index of the Dolbeault operator on manifolds, where, modulo a sign, the degree of a line bundle equals the first Chern class of its representatives. Especially, in the non-equivariant limit ${\rm q}\to 1$ the index becomes
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}}}\right)}={\rm ind}_{\rm orb}{\left({\partial}\right)}=\mathfrak{b}=1+\deg{\rm L}~{},$ (3.66)
reproducing the Riemann-Roch-Kawasaki theorem for orbifolds of genus zero [51]. The anti-holomorphic sections that contribute to a non-trivial ${\rm ind}_{\rm orb}\left(L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}\right)$ for $\mathfrak{b}\geq 1$ exhibit the same charges $\left(q_{G},r\right)$ as those of the eigenfunctions $\Phi_{n_{\sigma},n_{\psi}}\in\ker L_{\widetilde{P}}$ reported in (3.15). We can check that this is no coincidence by rewriting the index as
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}|_{\mathfrak{b}\geq 1}$ $=\sum_{n\in\mathbb{Z}}e^{-2{\rm i}{\epsilon}{k_{0}}n}e^{2{\rm i}{\epsilon}{k_{0}}{\left[{{\left({r/2}\right)}\alpha_{3}+q_{G}\varphi_{G}}\right]}}\sum_{\ell=-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}^{{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}{\rm q}^{\ell}~{},$
$=\sum_{n\in\mathbb{Z}}\sum_{\ell=-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}^{{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}e^{2{\rm i}{\epsilon}{k_{0}}{\left[{-n+{\left({r/2}\right)}\alpha_{3}+q_{G}\varphi_{G}+{\omega}\ell}\right]}}~{},$ (3.67)
and then exploiting the rule
${\rm ind}{\left({\mathscr{D};e^{-{\rm i}{\epsilon}\delta^{2}}}\right)}=\prod_{j}d_{j}\,e^{-{\rm i}{\epsilon}\lambda_{j}}\qquad\to\qquad Z_{\rm ind}=\prod_{j}\lambda_{j}^{-d_{j}}~{},$ (3.68)
relating the index of a differential operator $\mathscr{D}$ to the determinant $Z_{\rm ind}$, with $j$ being a multi-index, $\lambda_{j}$ the $j$-th eigenvalue of $\delta^{2}$ and $d_{j}$ the degeneracy of $\lambda_{j}$. Applying the prescription (3.68) to ${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}|_{\mathfrak{b}\geq 1}$ provides the infinite product
$Z_{\rm ind}|_{\mathfrak{b}\geq 1}$ $=\prod_{n\in\mathbb{Z}}\prod_{\ell=-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}^{{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}{\left[{2{k_{0}}{\left({n-{\left({r/2}\right)}\alpha_{3}-q_{G}\varphi_{G}-{\omega}\ell}\right)}}\right]}^{-1}~{},$
$=\prod_{n\in\mathbb{Z}}\prod_{j=-{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}^{{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}{\left[{2{k_{0}}{\left({n+{\omega}j-\frac{r}{2}\alpha_{3}-q_{G}\varphi_{G}}\right)}}\right]}^{-1}~{},$ (3.69)
exactly matching $Z^{\Phi}_{\text{1-L}}$ in (3.2.1) up to an irrelevant normalization factor. Therefore, the index ${\rm ind}_{\rm orb}\left(L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}\right)$ for $\mathfrak{b}\geq 1$ precisely encodes the eigenvalues associated with $\Phi_{n_{\sigma},n_{\psi}}$.
Finally, if
$-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}>1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}\iff\mathfrak{b}\leq-1~{},$ (3.70)
then
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}$ $={\rm q}^{{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}\frac{{\rm q}^{1-\mathfrak{b}}-{\rm q}}{1-{\rm q}}=-{\rm q}^{1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}{\left({1+{\rm q}^{2}+\dots+{\rm q}^{-\mathfrak{b}-1}}\right)}~{},$
$=-{\rm q}^{-{\left\lfloor{\mathfrak{p}_{-}^{\prime}/{n_{-}}}\right\rfloor}}{\left({1+{\rm q}^{2}+\dots+{\rm q}^{\deg{\rm L}^{\prime}}}\right)}~{},$ (3.71)
with
$\mathfrak{p}_{+}^{\prime}$ $=\mathfrak{p}_{+}+1=q_{G}{\mathfrak{m}_{+}}-\frac{r-2}{2}~{},$
$\mathfrak{p}_{-}^{\prime}$ $=\mathfrak{p}_{-}-1=q_{G}{\mathfrak{m}_{-}}+\frac{r-2}{2}~{},$
${\rm L}^{\prime}$ $=\mathcal{O}{\left({-{n_{+}}\mathfrak{p}_{-}^{\prime}+{n_{-}}\mathfrak{p}_{+}^{\prime}}\right)}=\mathcal{O}{\left[{q_{G}\mathfrak{m}+\frac{r-2}{2}{\left({{n_{+}}+{n_{-}}}\right)}}\right]}~{},$
$-{\left({\mathfrak{b}+1}\right)}$ $={\left\lfloor{-\frac{\mathfrak{p}_{+}^{\prime}}{{n_{+}}}}\right\rfloor}+{\left\lfloor{\frac{\mathfrak{p}_{-}^{\prime}}{{n_{-}}}}\right\rfloor}=\deg{\rm L}^{\prime}~{}.$ (3.72)
The overall factor ${\rm q}^{{\left\lfloor{\mathfrak{p}_{-}^{\prime}/{n_{-}}}\right\rfloor}}$ is the lift of the equivariant action at the south pole of $\mathbb{\Sigma}$. From ${\rm L}^{\prime}$ we see that the sections counted by ${\rm ind}_{\rm orb}\left(L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}\right)$ for $\mathfrak{b}\leq-1$ have charges ${\left({q_{G},r-2}\right)}$, which are those of the Grassmann-odd (fermionic) eigenfunctions $\mathcal{B}_{m_{\sigma},m_{\psi}}\in\ker L_{P}$ written in (3.14), where $L_{P}\sim\overline{\partial}$ at the fixed points of $\widehat{g}_{\mathbb{\Sigma}}$. Indeed, by writing
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}|_{\mathfrak{b}\leq-1}$ $=-\sum_{n\in\mathbb{Z}}e^{-2{\rm i}{\epsilon}{k_{0}}n}e^{2{\rm i}{\epsilon}{k_{0}}{\left[{{\left({r/2}\right)}\alpha_{3}+q_{G}\varphi_{G}}\right]}}\sum_{\ell=1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}^{-1-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}{\rm q}^{\ell}~{},$
$=-\sum_{n\in\mathbb{Z}}\,\sum_{\ell=1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}^{-1-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}e^{2{\rm i}{\epsilon}{k_{0}}{\left[{-n+{\left({r/2}\right)}\alpha_{3}+q_{G}\varphi_{G}+{\omega}\ell}\right]}}~{},$ (3.73)
and applying (3.68), we obtain
$Z_{\rm ind}|_{\mathfrak{b}\leq-1}$ $=\prod_{n\in\mathbb{Z}}\,\prod_{\ell=1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}^{-1-{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}{2{k_{0}}{\left({n-\frac{r}{2}\alpha_{3}-q_{G}\varphi_{G}-{\omega}\ell}\right)}}~{},$
$=\prod_{n^{\prime}\in\mathbb{Z}}\,\prod_{j=1+{\left\lfloor{\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}^{-1-{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}{2{k_{0}}{\left({n^{\prime}-\frac{r-2}{2}\alpha_{3}-q_{G}\varphi_{G}+{\omega}j}\right)}}~{},$ (3.74)
which is identical to $Z^{\mathcal{B}}_{\text{1-L}}$ in (3.2.1), up to a normalization. As the rule (3.68) maps ${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}|_{\mathfrak{b}=0}$ to $Z_{\rm ind}|_{\mathfrak{b}=0}=1$, we can conclude that the method of unpaired eigenvalues and the index theorem lead to the same result for the partition function of a chiral multiplet on topologically twisted $\mathbb{\Sigma}\times S^{1}$.
Anti-twist.
In the case of anti-twisted $\mathbb{\Sigma}\times S^{1}$ the $\varphi$-component of the R-symmetry background field vanishes at the poles of the spindle if
$\alpha_{2}|_{{\mathcal{U}_{+}}}=-\frac{1}{{n_{+}}}~{},$ $\alpha_{2}|_{{\mathcal{U}_{-}}}=-\frac{1}{{n_{-}}}~{},$ (3.75)
while flavour field (3.7) behaves as in the case of the topological twist. The pairing operator $L_{\widetilde{P}}$ is again transversally elliptic with respect to the free action of $\widehat{g}_{S^{1}}\in U(1)$, and near the poles of $\mathbb{\Sigma}$ assumes the value
$L_{\widetilde{P}}|_{{\mathcal{U}_{+}}}=-2{\rm i}\overline{\partial}_{+}~{},$ $L_{\widetilde{P}}|_{{\mathcal{U}_{-}}}=2{\rm i}\partial_{-}~{}.$ (3.76)
Repeating the methodology employed in the case of the topologically twisted $\mathbb{\Sigma}\times S^{1}$ yields
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}$ $=\sum_{n\in\mathbb{Z}}{\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}\chi_{n}{\left({\widehat{g}_{S^{1}}}\right)}\xi~{},$
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}$ $=\frac{{\rm q}^{{\left\lfloor{-\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}}{1-{\rm q}^{-1}}+\frac{{\rm q}^{{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}}{1-{\rm q}^{-1}}=\frac{{\rm q}^{{\left\lfloor{-\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}}{1-{\rm q}^{-1}}-\frac{{\rm q}^{1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}}{1-{\rm q}}~{}.$ (3.77)
Comparing (3.2.2) against (3.2.2) we see that the term in ${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}_{\mathbb{\Sigma}}}\right)}$ involving $\mathfrak{p}_{-}$ is independent of the R-symmetry twist present on $\mathbb{\Sigma}\times S^{1}$. The reason is that all these changes between twist and anti-twist case occurr only in the northern patch ${\mathcal{U}_{+}}$ of the spindle. Expanding ${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}$ in powers of ${\rm q}$ gives
${\rm ind}_{\rm orb}{\left({L_{\widetilde{P}};\widehat{g}}\right)}$ $=\sum_{n\in\mathbb{Z}}e^{2{\rm i}{\epsilon}{k_{0}}{\left[{-n+{\left({r/2}\right)}\alpha_{3}+q_{G}\varphi_{G}}\right]}}\sum_{\ell\in\mathbb{N}}{\left({{\rm q}^{-\ell+{\left\lfloor{-\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}-{\rm q}^{\ell+1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}}\right)}~{},$ (3.78)
which, according to the rule (3.68), corresponds to the infinite product
$Z_{\rm ind}$ $=\prod_{n\in\mathbb{Z}}\prod_{\ell\in\mathbb{N}}\frac{-n+{\left({r-2}\right)}{\left({\alpha_{3}/2}\right)}+q_{G}\varphi_{G}+{\omega}{\left({\ell+1+{\left\lfloor{-\mathfrak{p}_{-}/{n_{-}}}\right\rfloor}}\right)}}{-n+{\left({r/2}\right)}\alpha_{3}+q_{G}\varphi_{G}+{\omega}{\left({-\ell+{\left\lfloor{-\mathfrak{p}_{+}/{n_{+}}}\right\rfloor}}\right)}}~{},$
$=\prod_{n\in\mathbb{Z}}\prod_{\ell\in\mathbb{N}}\frac{-n+{\omega}{\left({\ell+\frac{\mathfrak{b}+1-\mathfrak{c}}{2}}\right)}-r\,\gamma_{R}-q_{G}\gamma_{G}-\alpha_{3}}{-n+{\omega}{\left({\ell+\frac{\mathfrak{b}-1+\mathfrak{c}}{2}}\right)}+r\,\gamma_{R}+q_{G}\gamma_{G}}~{}.$ (3.79)
The latter perfectly reproduces the one-loop determinant of a chiral multiplet on anti-twisted $\mathbb{\Sigma}\times S^{1}$ in (3.2.1), already combined with the prescription (3.46), as anticipated.
3.3 Canonical and effective classical terms
3.3.1 Canonical terms
General supersymmetric gauge theories include Chern-Simons and Fayet-Iliopoulos terms. The canonical Chern-Simons action at level ${\rm k}$ for a non-Abelian vector multiplet in the adjoint representation of a gauge group $G$ is
$S_{\rm CS}=\frac{{\rm i}\,{\rm k}}{4\pi}\int{\rm d}^{3}x\,{\rm e}\,{\rm tr}{\left[{\epsilon^{\mu\nu\rho}{\left({{\mathcal{A}}_{\mu}\partial_{\nu}{\mathcal{A}}_{\rho}-\frac{2{\rm i}}{3}{\mathcal{A}}_{\mu}{\mathcal{A}}_{\nu}{\mathcal{A}}_{\rho}}\right)}+2{\rm i}D\sigma+2\widetilde{\lambda}\lambda}\right]}~{},$ ${\rm e}=\epsilon_{x\varphi\psi}~{},$ (3.80)
which, once evaluated on the BPS locus, contributes to the classical part of the partition function as follows:
$Z^{\rm CS}_{\rm class}|_{\rm BPS}=e^{-S_{\rm CS}}|_{\rm BPS}=e^{2\pi{\rm i}\,{\rm k}\gamma_{G}\mathfrak{f}_{G}}=e^{2\pi{\rm i}\,{\rm k}\,\gamma_{G}\,\mathfrak{m}/{\left({{n_{+}}{n_{-}}}\right)}}~{}.$ (3.81)
The latter can be generalized to a mixed Chern-Simons term at level ${\rm k}_{ij}$ that combines two vector multiplets in the adjoint of different gauge groups $G_{(i)}$ and $G_{(j)}$ with $i\neq j$:
$S_{\rm MCS}=\frac{{\rm i}\,{\rm k}_{ij}}{2\pi}\int{\rm d}^{3}x\,e\,{\left({\epsilon^{\mu\nu\rho}{\mathcal{A}}_{\mu}^{(i)}\partial_{\nu}{\mathcal{A}}_{\rho}^{(j)}+{\rm i}D^{(i)}\sigma^{(j)}+{\rm i}D^{(j)}\sigma^{(i)}+\widetilde{\lambda}^{(i)}\lambda^{(j)}+\widetilde{\lambda}^{(j)}\lambda^{(i)}}\right)}~{},$ (3.82)
yielding
$Z^{\rm MCS}_{\rm class}|_{\rm BPS}=e^{2\pi{\rm i}\,{\rm k}_{ij}{\left({\gamma_{G_{i}}\mathfrak{f}_{G_{j}}+\gamma_{G_{j}}\mathfrak{f}_{G_{i}}}\right)}}~{}.$ (3.83)
The gauge symmetry can also be mixed with the R-symmetry via an R-symmetry-gauge Chern-Simons term, whose action reads [34]
$S_{\rm RCS}=\frac{{\rm i}\,{\rm k}_{R}}{2\pi}\int{\rm d}^{3}x\,e\,{\left[{\epsilon^{\mu\nu\lambda}{\mathcal{A}}_{\mu}\partial_{\nu}{\left({A_{\lambda}-\frac{1}{2}V_{\lambda}}\right)}+{\rm i}D\sigma+\frac{{\rm i}\,\sigma}{4}{\left({R+2V_{\mu}V^{\mu}+2H^{2}}\right)}}\right]}~{},$ (3.84)
giving
$Z^{\rm RCS}_{\rm class}|_{\rm BPS}=e^{2\pi{\rm i}\,{\rm k}_{R}{\left({\gamma_{R}\mathfrak{f}_{G}+\gamma_{G}\mathfrak{f}_{R}}\right)}}~{},$ (3.85)
which depends on the twist through the R-symmetry fugacity $\gamma_{R}$ defined in (3.2.1) and the flux $\mathfrak{f}_{R}=\chi_{\pm}/2$. Finally, for any Abelian factor in a gauge group $G$, a classical contribution to the partition function can also descend from Fayet-Iliopoulos terms with parameter $\zeta_{\rm FI}$ and action
$S_{\rm FI}$ $=\frac{\zeta_{\rm FI}}{2\pi}\int{\rm d}^{3}x\,e{\left({D-{\mathcal{A}}_{\mu}V^{\mu}-\sigma H}\right)}~{}.$ (3.86)
In fact, thanks to the relation [34]
$V^{\mu}=-{\rm i}\epsilon^{\mu\nu\lambda}\partial_{\nu}\mathcal{C}_{\lambda}~{},$ (3.87)
the Fayet-Iliopoulos term can be interpreted as a mixed Chern-Simons term coupling fields in an Abelian vector multiplet to the background scalar $H$ and the graviphoton field $\mathcal{C}_{\mu}$. Hence,
$Z^{\rm FI}_{\rm class}|_{\rm BPS}=e^{2\pi{\rm i}\,\zeta_{\rm FI}{\left({\gamma_{\mathcal{C}}\mathfrak{f}_{G}+\gamma_{G}\mathfrak{f}_{\mathcal{C}}}\right)}}~{},$ (3.88)
where $\gamma_{\mathcal{C}}$ and $\mathfrak{f}_{\mathcal{C}}$ are the fugacity and the flux associated to $\mathcal{C}_{\mu}$, respectively.
3.3.2 Effective terms
In general, the classical contributions obtained from evaluating on the BPS locus the canonical actions above are not invariant under large gauge transformations along $S^{1}$. For example, under a large gauge transformation acting as $\gamma_{G}\to\gamma_{G}+\ell$, with $\ell\in\mathbb{N}$, the classical part of the canonical Chern-Simons partition function term behaves as follows:
$Z^{\rm CS}_{\rm class}|_{\rm BPS}=e^{2\pi{\rm i}\,{\rm k}\,\gamma_{G}\,\mathfrak{m}/{\left({{n_{+}}{n_{-}}}\right)}}\qquad\to\qquad\widetilde{Z}^{\rm CS}_{\rm class}|_{\rm BPS}=e^{2\pi{\rm i}\,{\rm k}\ell\mathfrak{m}/{\left({{n_{+}}{n_{-}}}\right)}}Z^{\rm CS}_{\rm class}|_{\rm BPS}~{}.$ (3.89)
In the case in which the spindle is a sphere, then ${n_{+}}={n_{-}}=1$ and the partition function exhibits invariance under large gauge transformations as $\widetilde{Z}^{\rm CS}_{\rm class}|_{\rm BPS}=Z^{\rm CS}_{\rm class}|_{\rm BPS}$. Conversely, for a general spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$, the parameters ${\left({{n_{+}},{n_{-}}}\right)}$ are arbitrary coprime integers and $\widetilde{Z}^{\rm CS}_{\rm class}|_{\rm BPS}\neq Z^{\rm CS}_{\rm class}|_{\rm BPS}$ unless either the Chern-Simons level ${\rm k}$ or $\mathfrak{m}$ is a multiple of ${\left({{n_{+}}{n_{-}}}\right)}$.
A similar circumstance arises in supersymmetric theories on the manifold with boundary $D^{2}\times S^{1}$, where $D^{2}$ represents a two-dimensional disk or hemisphere. The canonical Chern-Simons term lacks supersymmetry and gauge invariance on $D^{2}\times S^{1}$. Restoring these properties requires taking into account degrees of freedom localized at the boundary $\partial{\left({D^{2}\times S^{1}}\right)}=T^{2}$, with $T^{2}=S^{1}\times S^{1}$ being a two-dimensional torus. A convenient method for promptly deriving the correct expression for the effective Chern-Simons term on $D^{2}\times S^{1}$ was elucidated in [52], where it was demonstrated that the partition function for a Chern-Simons term at level ${\rm k}=\pm 1$ equals the collective partition function of two anomaly-free chiral multiplets ${\left({\phi_{(1)},\phi_{(2)}}\right)}$ coupled by a superpotential $W{\left({\phi_{(1)},\phi_{(2)}}\right)}=m_{0}\phi_{(1)}\phi_{(2)}$, where $m_{0}$ encodes a mass. Such a procedure allowed to show that the Chern-Simons partition function on $\partial{\left({D^{2}\times S^{1}}\right)}=T^{2}$ is given by the theta function $\theta{\left({{\bf x};{\bf q}}\right)}={\left({-\sqrt{\bf q}/{\bf x};{\bf q}}\right)}_{\infty}{\left({-\sqrt{\bf q}\,{\bf x};{\bf q}}\right)}_{\infty}$ depending on gauge and flavour fugacities ${\bf x}$, as well as on an equivariant parameter $\epsilon$ appearing in the fugacity ${\bf q}=e^{\epsilon}$ for the angular momentum on $D^{2}$. In particular, gauge invariance of such an effective Chern-Simons term is ensured by the analytic properties of the Theta function.
We can apply the same procedure to the case of $\mathbb{\Sigma}\times S^{1}$. Two chiral multiplets ${\left({\phi_{(1)},\phi_{(2)}}\right)}$ coupled by the aforementioned superpotential $W{\left({\phi_{(1)},\phi_{(2)}}\right)}=m_{0}\phi_{(1)}\phi_{(2)}$ have charges satisfying
$r^{(1)}+r^{(2)}=2~{},$ $q_{G}^{(1)}+q_{G}^{(2)}=0~{},$ (3.90)
which, in the case of the twisted spindle [^2] Similar relations hold in the case of the anti-twisted spindle $\mathbb{\Sigma}$., imply
$\mathfrak{b}^{(1)}+\mathfrak{b}^{(2)}=0~{},\qquad\mathfrak{c}^{(1)}+\mathfrak{c}^{(2)}=-\chi_{-}~{},$ $y_{(1)}y_{(2)}=1~{},$ (3.91)
where ${\left({\mathfrak{b}^{(i)},\mathfrak{c}^{(i)},y_{(i)}}\right)}$ are ${\left({\mathfrak{b},\mathfrak{c},y}\right)}$ evaluated at $r=r^{(i)}$ and $q_{G}=q_{G}^{(i)}$ with $i=1,2$. In particular, $y_{(1)}y_{(2)}=1$ follows from the constraint $\gamma_{R}-{\left({{\omega}\chi_{-}/4}\right)}\in\mathbb{Z}$. As in [52], we consider anomaly-free chiral multiplets, whose partition function is obtained from $Z^{\rm CM}_{\text{1-L}}$ by removing the phase factors and considering the q-Pochhammer part only. For instance, in the case of the topologically twisted $\mathbb{\Sigma}\times S^{1}$, we have
$Z^{(2)}$ $=\frac{1}{{\left({q^{{\left({1-\mathfrak{b}_{2}}\right)}/2}y_{(2)}^{-1};q}\right)}_{\mathfrak{b}_{2}}}=\frac{1}{{\left({q^{{\left({1+\mathfrak{b}_{1}}\right)}/2}y_{(1)};q}\right)}_{-\mathfrak{b}_{1}}}~{},$
$={\left({-y_{(1)}}\right)}^{\mathfrak{b}_{1}}{\left({q^{{\left({1-\mathfrak{b}_{1}}\right)}/2}y_{(1)}^{-1};q}\right)}_{\mathfrak{b}_{1}}={\left({-y_{(1)}}\right)}^{\mathfrak{b}_{1}}/{Z^{(1)}}~{},$ (3.92)
yielding
$Z^{(1)}Z^{(2)}={\left({-y_{(1)}}\right)}^{\mathfrak{b}_{1}}~{}.$ (3.93)
Then, according to [52] the partition function of the effective Chern-Simons term at level ${\rm k}=\pm 1$ has the form $Z^{\rm CS}_{\rm eff}=y^{\pm\mathfrak{b}}$. Since we have fixed only the sum of the R-charges ${\left({q_{G}^{(i)},r^{(i)}}\right)}$ but not their specific values, $Z^{\rm CS}_{\rm eff}$ contains information on all global and gauge symmetries affecting the chirals ${\left({\phi_{(1)},\phi_{(2)}}\right)}$. Indeed, if ${n_{+}}={n_{-}}=1$, the spindle $\mathbb{\Sigma}$ boils down to a two-sphere and
$Z^{\rm CS}_{\rm eff}|_{{n_{+}}={n_{-}}=1}=e^{2\pi{\rm i}q_{G}\gamma_{G}{\left({1-r-q_{G}\mathfrak{m}}\right)}}~{}.$ (3.94)
At arbitrary level ${\rm k}$ we have
$Z^{\rm CS}_{\rm eff}=y^{{\rm k}\,\mathfrak{b}}~{},$ (3.95)
More generally, an effective Chern-Simons term mixing two gauge groups ${\left({G_{i},G_{j}}\right)}$ reads
$Z^{\rm MCS}_{\rm eff}=y_{(i)}^{{\rm k}_{ij}\,\mathfrak{b}_{j}}y_{(j)}^{{\rm k}_{ij}\,\mathfrak{b}_{i}}~{}.$ (3.96)
The partition function (3.95) is manifestly invariant under large gauge transformations:
$Z^{\rm CS}_{\rm eff}|_{\rm BPS}\qquad\to\qquad\widetilde{Z}^{\rm CS}_{\rm eff}|_{\rm BPS}=e^{2\pi{\rm i}\,{\rm k}\ell\mathfrak{b}}Z^{\rm CS}_{\rm eff}|_{\rm BPS}=Z^{\rm CS}_{\rm eff}|_{\rm BPS}~{}.$ (3.97)
In presence of matter, the Chern-Simons level ${\rm k}$ is affected by non-trivial corrections: for instance, a chiral multiplet causes an half-integer shift ${\rm k}\to{\rm k}^{\prime}={\rm k}+{\left({1/2}\right)}$, as we see from (3.48).
Drawing an analogy from the case of $D^{2}\times S^{1}$, we conjecture that the modifications to the canonical Chern-Simons theory’s stem from degrees of freedom localized at the conical singularities of the spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{[{n_{+}},{n_{-}}]}$. We leave the proof of this hypothesis for future investigation; however, in Section 3.4 we assess the validity of (3.95) through the application of supersymmetric dualities. We use partition functions related to the topological twist, as the anti-twist case is similar.
3.4 Examples and dualities
3.4.1 Effective $U(1)$ Chern-Simons theory at level ${\rm k}$
As a first example let us calculate the effective partition function of an Abelian Chern-Simons theory with gauge group $U(1)$ at level ${\rm k}$. The corresponding contour integral reads
$Z=-\sum_{\mathfrak{m}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}y^{{\rm k}\mathfrak{b}}y^{\mathfrak{n}}z^{\mathfrak{b}}~{},$ (3.98)
where $y$ is the gauge fugacity, $\mathfrak{m}$ the label of the gauge flux $\mathfrak{f}_{G}=\mathfrak{m}/{\left({{n_{+}}{n_{-}}}\right)}$, while $z$ and $\mathfrak{n}$ are background flux and fugacity, respectively. The integer $\mathfrak{b}$ is given by
$\mathfrak{b}=\mathfrak{b}{\left({\mathfrak{m}}\right)}=1+{\left\lfloor{\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}+{\left\lfloor{-\frac{{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rfloor}=1+{\left\lfloor{\frac{\mathfrak{m}\,{\mathfrak{a}_{+}}}{{n_{+}}}}\right\rfloor}+{\left\lfloor{-\frac{\mathfrak{m}\,{\mathfrak{a}_{-}}}{{n_{-}}}}\right\rfloor}~{},$ ${n_{+}}{\mathfrak{a}_{-}}-{n_{-}}{\mathfrak{a}_{+}}=1~{}.$ (3.99)
We write $\mathfrak{m}\in\mathbb{Z}$ as
$\mathfrak{m}={n_{+}}{n_{-}}\mathfrak{m}^{\prime}+\mathfrak{r}~{},$ $\mathfrak{r}=0,\dots,{\left({{n_{+}}{n_{-}}-1}\right)}~{},$ (3.100)
with $\mathfrak{m}^{\prime}\in\mathbb{Z}$. The integral then becomes
$Z$ $=-\sum_{\mathfrak{m}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}y^{{\rm k}\mathfrak{b}+\mathfrak{n}}z^{\mathfrak{b}}~{},$
$=-\sum_{\mathfrak{r}=0}^{{n_{+}}{n_{-}}-1}\sum_{\mathfrak{m}^{\prime}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}y^{-{\rm k}\mathfrak{m}^{\prime}+{\rm k}\mathfrak{b}{\left({\mathfrak{r}}\right)}+\mathfrak{n}}z^{-\mathfrak{m}^{\prime}+\mathfrak{b}{\left({\mathfrak{r}}\right)}}~{},$
$=-\sum_{\mathfrak{r}=0}^{{n_{+}}{n_{-}}-1}\sum_{\mathfrak{m}^{\prime\prime}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}y}y^{{\rm k}\mathfrak{m}^{\prime\prime}+\mathfrak{n}}z^{\mathfrak{m}^{\prime\prime}}~{}.$ (3.101)
The latter is non-vanishing only if the contour $\mathscr{C}$ surrounding $y=0$ and if
${\rm k}\mathfrak{m}^{\prime\prime}+\mathfrak{n}=0~{},$ (3.102)
which is a Diophantine equation that can be solved by conveniently splitting $\mathfrak{n}$ as
$\mathfrak{n}={\rm k}\mathfrak{n}^{\prime}+{\left\llbracket{\mathfrak{n}}\right\rrbracket_{{\rm k}}}~{}.$ (3.103)
Indeed, plugging the expression above into the Diophantine equation gives
$\mathfrak{m}^{\prime\prime}=-\mathfrak{n}^{\prime}~{},$ ${\left\llbracket{\mathfrak{n}}\right\rrbracket_{{\rm k}}}=0~{},$ (3.104)
which yields the final result
$Z$ $={n_{+}}{n_{-}}\,z^{-\mathfrak{n}^{\prime}}\,\delta_{{\left\llbracket{\mathfrak{n}}\right\rrbracket_{{\rm k}}},0}={n_{+}}{n_{-}}\,z^{-{\left\lfloor{\mathfrak{n}/{\rm k}}\right\rfloor}}\,\delta_{{\left\llbracket{\mathfrak{n}}\right\rrbracket_{{\rm k}}},0}~{},$ (3.105)
meaning that $Z\neq 0$ only if ${\left\llbracket{\mathfrak{n}}\right\rrbracket_{{\rm k}}}=0$. Especially, if $\mathfrak{n}=0$, we find $Z={\left({{n_{+}}{n_{-}}}\right)}$, suggesting that the effective Chern-Simons theory on $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}},{n_{-}}}\right]}}$ has ${\left({{n_{+}}{n_{-}}}\right)}$ vacua. Technically, this degeneracy stems from the fact that for any $\overline{\mathfrak{b}}\in\mathbb{Z}$ there are ${\left({{n_{+}}{n_{-}}}\right)}$ values of $\mathfrak{m}\in\mathbb{Z}$ such that $\mathfrak{b}{\left({\mathfrak{m}}\right)}=\overline{\mathfrak{b}}$.
3.4.2 Effective $U(1)$ Chern-Simons theory coupled to a chiral multiplet
As another example we compute the partition function of a chiral multiplet coupled to an effective Chern-Simons term with gauge group $U(1)$ at level ${\rm k}=1$. We also add a mixed Chern-Simons term with flavour fugacity ${\left({-z}\right)}$ and effective flux $\widetilde{\mathfrak{b}}$. We again consider a countour ${\mathscr{C}}$ such that all contributions to the integral comes from the pole at $y=0$. The collective partition function is
$Z=\sum_{\mathfrak{m}\in\mathbb{Z}}\int_{\mathscr{C}}\frac{{\rm d}y}{2\pi{\rm i}\,y}\frac{{\left({-z}\right)}^{\mathfrak{b}}y^{\widetilde{\mathfrak{b}}}y^{\mathfrak{b}-1}}{{\left({y\,q^{{\left({1-\mathfrak{b}}\right)}/2};q}\right)}_{\mathfrak{b}}}~{},$ (3.106)
where the finite q-Pochhammer symbol corresponds to a chiral multiplet on topologically twisted $\mathbb{\Sigma}\times S^{1}$, while $y^{\mathfrak{b}-1}$ is the spindle counterpart of $e^{2\pi{\rm i}\gamma_{G}\mathfrak{m}}$, which is the Chern-Simons term on $S^{2}\times S^{1}$. It is convenient to expand the integrand as
$Z$ $=\sum_{\mathfrak{m}\in\mathbb{Z}}\int\frac{{\rm d}y}{2\pi{\rm i}y}\frac{{\left({-z}\right)}^{\mathfrak{b}}y^{\widetilde{\mathfrak{b}}+\mathfrak{b}-1}}{{\left({y\,q^{{\left({1-\mathfrak{b}}\right)}/2};q}\right)}_{\mathfrak{b}}}=\sum_{\mathfrak{m}\in\mathbb{Z}}\sum_{\ell\in\mathbb{N}}\int\frac{{\rm d}y}{2\pi{\rm i}y}{\left({-z}\right)}^{\mathfrak{b}}y^{\widetilde{\mathfrak{b}}+\mathfrak{b}-1+\ell}\frac{{\left({q^{\mathfrak{b}};q}\right)}_{\ell}}{{\left({q;q}\right)}_{\ell}}q^{\ell{\left({1-\mathfrak{b}}\right)}/2}~{},$ (3.107)
where we employed the $q$-binomial theorem
$\frac{1}{{\left({t;q}\right)}_{\mathfrak{b}}}=\frac{{\left({t\,q^{\mathfrak{b}};q}\right)}_{\infty}}{{\left({t;q}\right)}_{\infty}}=\sum_{\ell\in\mathbb{N}}\frac{{\left({q^{\mathfrak{b}};q}\right)}_{\ell}}{{\left({q;q}\right)}_{\ell}}t^{\ell}~{},$ (3.108)
valid for both finite and infinite q-Pochhammer symbols. The contour integral providing $Z$ is non-vanishing if
$\widetilde{\mathfrak{b}}+\mathfrak{b}-1+\ell=0~{},$ (3.109)
which is an integer-valued equation that can be solved by writing $\mathfrak{m}$ as
$\mathfrak{m}={n_{+}}{n_{-}}\mathfrak{m}^{\prime}+\mathfrak{r}~{},$ $\mathfrak{m}^{\prime}\in\mathbb{Z}\qquad\mathfrak{r}=0,\dots,{\left({{n_{+}}{n_{-}}-1}\right)}~{},$ (3.110)
yielding
$\widetilde{\mathfrak{b}}-\mathfrak{m}^{\prime}+\mathfrak{b}{\left({\mathfrak{r}}\right)}-1+\ell=0~{}.$ (3.111)
Inserting the latter in the expression for $Z$ gives
$Z$ $=-\sum_{\mathfrak{r}=0}^{{n_{+}}{n_{-}}-1}\sum_{\ell\in\mathbb{N}}{\left({-z}\right)}^{1-\widetilde{\mathfrak{b}}-\ell}\frac{{\left({q^{1-\widetilde{\mathfrak{b}}-\ell};q}\right)}_{\ell}}{{\left({q;q}\right)}_{\ell}}q^{\ell{\left({\mathfrak{n}+\ell}\right)}/2}~{},$
$=-{n_{+}}{n_{-}}\sum_{\ell\in\mathbb{N}}{\left({-z}\right)}^{1-\widetilde{\mathfrak{b}}-\ell}\frac{{\left({q^{\widetilde{\mathfrak{b}}};q}\right)}_{\ell}}{{\left({q;q}\right)}_{\ell}}q^{\ell{\left({\widetilde{\mathfrak{b}}+\ell}\right)}/2}{\left({-q^{1-\widetilde{\mathfrak{b}}-\ell}}\right)}^{\ell}q^{\ell{\left({\ell-1}\right)}/2}~{},$
$=-{n_{+}}{n_{-}}\sum_{\ell\in\mathbb{N}}{\left({-z}\right)}^{1-\widetilde{\mathfrak{b}}}\frac{{\left({q^{\widetilde{\mathfrak{b}}};q}\right)}_{\ell}}{{\left({q;q}\right)}_{\ell}}{\left({q^{{\left({1-\widetilde{\mathfrak{b}}}\right)}/2}z^{-1}}\right)}^{\ell}~{},$
$=-{n_{+}}{n_{-}}\frac{{\left({-z}\right)}^{1-\widetilde{\mathfrak{b}}}}{{\left({q^{{\left({1-\widetilde{\mathfrak{b}}}\right)}/2}z^{-1};q}\right)}_{\widetilde{\mathfrak{b}}}}~{},$
$={n_{+}}{n_{-}}\frac{z}{{\left({z\,q^{{\left({1-\widetilde{\mathfrak{b}}}\right)}/2};q}\right)}_{\widetilde{\mathfrak{b}}}}~{},$ (3.112)
where in the second, fourth and fifth equalities we exploited the identity
${\left({t;q}\right)}_{\ell}={\left({t^{-1}\,q^{1-\ell};q}\right)}_{\ell}{\left({-t\,q^{{\left({\ell-1}\right)}/2}}\right)}^{\ell}~{}.$ (3.113)
Therefore, we find that on $\mathbb{\Sigma}\times S^{1}$ the partition function of a chiral multiplet coupled to an effective Chern-Simons theory is dual to a free chiral multiplet, interpreted as describing a monopole operator. The overall factor ${\left({{n_{+}}{n_{-}}}\right)}$ originates from a sum over $\mathfrak{r}$ in the integer range going from 00 to ${\left({{n_{+}}{n_{-}}-1}\right)}$, spanning degenerate sectors of the original theory.
4 Localization on the branched lens space
4.1 BPS locus
We parametrize Abelian gauge or flavour fields on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ as follows:
${\mathcal{A}}={\mathcal{A}}_{\varphi}{\left({x}\right)}{\rm d}\varphi+\frac{h}{n}{\rm d}\psi~{},$ ${\mathfrak{f}}_{G}=\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{\mathcal{A}}=\frac{{\mathfrak{m}}}{{n_{+}}{n_{-}}}~{},\qquad\exp{\left({{\rm i}\oint_{S^{1}}{\mathcal{A}}}\right)}=e^{2\pi{\rm i}h/n}~{},$ (4.1)
with ${\mathfrak{m}}\in\mathbb{Z}$ labelling the flux through the spindle base and $h=0,\dots,{\left({n-1}\right)}$ the $\mathbb{Z}_{n}$ holonomy along the Hopf fiber. We again gauged away the component ${\mathcal{A}}_{x}{\left({x}\right)}$. As in the case of $\mathbb{\Sigma}\times S^{1}$, the behaviour of ${\mathcal{A}}_{\varphi}$ is
${\mathcal{A}}_{\varphi}{\left({0}\right)}=\frac{{\mathfrak{m}_{+}}}{{n_{+}}}~{},\qquad{\mathcal{A}}_{\varphi}{\left({\frac{\pi}{2}}\right)}=\frac{{\mathfrak{m}_{-}}}{{n_{-}}}~{},$ ${n_{+}}{\mathfrak{m}_{-}}-{n_{-}}{\mathfrak{m}_{+}}={\mathfrak{m}}~{},$ (4.2)
with $\mathfrak{m}_{\pm}=\mathfrak{m}\,\mathfrak{a}_{\pm}$ and ${n_{+}}{\mathfrak{a}_{-}}-{n_{-}}{\mathfrak{a}_{+}}=1$.
If we do not impose reality conditions on fields, the vector-multiplet BPS equations $\delta\lambda=\delta\widetilde{\lambda}=0$ require that on the supersymmetric locus the dynamical scalar $\sigma$ and the auxiliary scalar $D$ are
$\sigma|_{\rm BPS}=\sigma_{c}+\frac{{\rm i}}{b_{1}b_{2}}{\left[{-{n_{+}}b_{1}h+{\left({{n_{-}}b_{1}+{n_{+}}b_{2}}\right)}{\mathcal{A}}_{\varphi}{\left({x}\right)}}\right]}~{},$
$D|_{\rm BPS}=\frac{1}{b_{1}b_{2}f}{\left[{{\rm i}\,b_{1}b_{2}\sigma|_{\rm BPS}+{\left({{n_{+}}b_{2}\cot x-{n_{-}}b_{1}\tan x}\right)}{\mathcal{A}}_{\varphi}^{\prime}{\left({x}\right)}}\right]}~{},$ (4.3)
where $\sigma_{c}$ is a constant real mass in the cause of a weakly gauged theory with a non-dynamical background vector multiplet. Instead, $\sigma_{c}$ is integrated over in the path integral in the case of a dynamical vector multiplet.
The chiral multiplet BPS-equations $\delta\psi=\delta\widetilde{\psi}=0$, for generic values of flavour and R-symmetry fugacities require a trivial supersymmetric locus for matter fields:
$\phi|_{\rm BPS}=\widetilde{\phi}|_{\rm BPS}=F|_{\rm BPS}=\widetilde{F}|_{\rm BPS}=0~{}.$ (4.4)
4.2 One-loop determinants
We calculate the one-loop determinant for a chiral multiplet on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ via the same technique adopted in the case of $\mathbb{\Sigma}\times S^{1}$. Especially, the eigenfunction $B_{m_{\varphi},m_{\psi}}{\left({x}\right)}$ is non-singular at $x=0,\pi/2$ if the integers ${\left({m_{\varphi},m_{\psi}}\right)}$ do not violate the following bounds:
${n_{+}}m_{\varphi}\leq{\mathfrak{m}_{+}}~{},$ ${n_{-}}{\left({m_{\varphi}-nm_{\psi}+h}\right)}\leq{\mathfrak{m}_{-}}~{}.$ (4.5)
The latter are satisfied if
$m_{\varphi}$ $=-n\,\ell_{\varphi}-{\left\llbracket{h+{\left\llbracket{\Delta_{+}}\right\rrbracket_{n}}+j_{\psi}}\right\rrbracket_{n}}+{\left\lfloor{\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}~{},$
$m_{\psi}$ $=\ell_{\psi}-\ell_{\varphi}+{\left\lfloor{\frac{\Delta_{+}}{n}}\right\rfloor}+{\left\lfloor{\frac{h+{\left\llbracket{\Delta_{+}}\right\rrbracket_{n}}+j_{\psi}}{n}}\right\rfloor}~{},$
$\Delta_{\pm}$ =±$={\left\lfloor{\pm\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}-{\left\lfloor{\pm\frac{{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rfloor}~{},\qquad\ell_{\psi},\ell_{\varphi}\in\mathbb{N}~{},\qquad j_{\psi}=0,\dots,{\left({n-1}\right)}~{}.$ (4.6)
Correspondingly, the eigenfunction $\phi_{n_{\varphi},n_{\psi}}{\left({x}\right)}$ is non-singular at $x=0,\pi/2$ if the integer inequalities
${n_{+}}n_{\varphi}\geq{\mathfrak{m}_{+}}~{},$ ${n_{-}}{\left({n_{\varphi}-nn_{\psi}+h}\right)}\geq{\mathfrak{m}_{-}}~{},$ (4.7)
hold. They are fulfilled if
$n_{\varphi}$ $=n\,k_{\varphi}+t_{\varphi}-{\left\lfloor{-\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}~{},$
$n_{\psi}$ $=-k_{\psi}+k_{\varphi}+{\left\lfloor{-\frac{\Delta_{-}}{n}}\right\rfloor}+{\left\lfloor{\frac{h+{\left\llbracket{-\Delta_{-}}\right\rrbracket_{n}}+t_{\varphi}}{n}}\right\rfloor}~{},$
$\ell_{\psi},\ell_{\varphi}\in\mathbb{N}~{},\qquad t_{\varphi}=0,\dots,{\left({n-1}\right)}~{}.$ (4.8)
Altogether, the one-loop determinant generated by $\mathcal{B}_{m_{\varphi},m_{\psi}}$ and $\Phi_{n_{\varphi},n_{\psi}}$ is
$Z^{\mathcal{B},\Phi}_{\text{1-L}}$ $=\prod_{j=0}^{n-1}\prod_{k,\ell\in\mathbb{N}}\frac{\frac{1}{b}{\left({k-\frac{1}{n}{\left\lfloor{\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}+\frac{{\left\llbracket{h+{\left\llbracket{\Delta_{+}}\right\rrbracket_{n}}+j}\right\rrbracket_{n}}}{n}}\right)}+b{\left({\ell-\frac{1}{n}{\left\lfloor{\frac{{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rfloor}+\frac{j}{n}}\right)}+\frac{Q}{2n}+\frac{{\rm i}\,u}{n}}{\frac{1}{b}{\left({k-\frac{1}{n}{\left\lfloor{-\frac{{\mathfrak{m}_{+}}}{{n_{+}}}}\right\rfloor}+\frac{j}{n}}\right)}+b{\left({\ell-\frac{1}{n}{\left\lfloor{-\frac{{\mathfrak{m}_{-}}}{{n_{-}}}}\right\rfloor}+\frac{{\left\llbracket{h+{\left\llbracket{-\Delta_{-}}\right\rrbracket_{n}}+j}\right\rrbracket_{n}}}{n}}\right)}+\frac{Q}{2n}-\frac{{\rm i}\,u}{n}}~{},$
$b$ $=\sqrt{\frac{{n_{-}}b_{1}}{{n_{+}}b_{2}}}~{},\qquad Q={\left({\frac{b^{-1}}{{n_{+}}}+\frac{b}{{n_{-}}}}\right)}~{},\qquad u=\sqrt{\frac{b_{1}b_{2}}{{n_{+}}{n_{-}}}}\,\sigma_{c}-{\rm i}{\left({1-r}\right)}\frac{Q}{2}-{\rm i}\,b\,h~{}.$ (4.9)
Hence, on ${L_{b_{1},b_{2},{n_{+}},{n_{-}}}{\left({n,1}\right)}}$ the one-loop determinant for a chiral and a vector multiplet with weights $\rho\in\mathfrak{R}_{G}$ and $\alpha\in{\rm adj}\,G$ for a gauge group $G$ read
$Z^{\rm CM}_{\text{1-L}}=\prod_{\rho\in\mathfrak{R}_{G}}Z^{\mathcal{B},\Phi}_{\text{1-L}}{\left({\rho{\left({\mathfrak{m}_{\pm}}\right)},\rho{\left({u}\right)},\rho{\left({h}\right)}}\right)}~{},$ $Z^{\rm VM}_{\text{1-L}}=\prod_{\alpha\in{\rm adj}\,G}Z^{\mathcal{B},\Phi}_{\text{1-L}}{\left({\alpha{\left({\mathfrak{m}_{\pm}}\right)},\alpha{\left({u}\right)},\alpha{\left({h}\right)}}\right)}|_{r=2}~{}.$ (4.10)
These match the one-loop determinants on the squashed lens space computed in [31] upon setting $n_{\pm}=1$ and employing the function
$\langle x\rangle_{n}=\frac{1}{n}{\left({{\left\llbracket{x}\right\rrbracket_{n}}+\frac{1}{2}}\right)}-\frac{1}{2}~{}.$ (4.11)
Eventually, if $n_{\pm}=n=1$, the one-loop determinant above reduces to that of the squashed three-sphere. In this limit the one-loop determinant above also matches the determinant of fluctuations for matter fields computed in [32] on the resolved space of the branched sphere, once the identifications $b_{1}=q_{\rm there}$ and $b_{2}=p_{\rm there}$ are made.
5 Spindle from the squashed three-sphere
In this section we derive one-loop determinants for matter and gauge fields on the pure two-dimensional spindle starting from three dimensions. Indeed, by exploiting the fact that the spindle non-trivially embeds into the squashed three-sphere $S^{3}_{q,p}$, we reduce the one-loop determinant associated with an $\mathcal{N}=2$ chiral multiplet on $S^{3}_{q,p}$ to the one-loop determinant of an $\mathcal{N}=(2,2)$ chiral multiplet on $\mathbb{\Sigma}$. In turn, this will provide the one-loop determinant of a $\mathcal{N}=(2,2)$ vector multiplet.
5.1 $SL(2,\mathbb{Z})$ reduction
5.1.1 General Construction
Let the change of coordinates from $\theta\in{\left[{0,\pi/2}\right]}$ and $\phi_{1,2}\in{\left[{0,2\pi}\right]}$ to coordinates $\Theta\in{\left[{0,\pi}\right]}$ and $\sigma,y\in{\left[{0,2\pi}\right]}$ be
$\theta$ $=\Theta/2~{},$
$\phi_{1}$ $=t_{11}\,\sigma+t_{12}\,y~{},$
$\phi_{2}$ $=t_{21}\,\sigma+t_{22}\,y~{},$ (5.1)
with $t\in SL(2,\mathbb{Z})$. The inverse change of coordinates reads
$\Theta$ $=2\theta~{},$
$\sigma$ $=t_{22}\phi_{1}-t_{12}\phi_{2}~{},$
$y$ $=-t_{21}\phi_{1}+t_{11}\phi_{2}~{}.$ (5.2)
Also,
$\partial_{\theta}$ $=2\partial_{\Theta}~{},$
$\partial_{\phi_{1}}$ $=t_{22}\partial_{\sigma}-t_{21}\partial_{y}~{},$
$\partial_{\phi_{2}}$ $=-a_{12}\partial_{\sigma}+a_{11}\partial_{y}~{},$ (5.3)
In ${\left({\Theta,\sigma,y}\right)}$ coordinates the line element of $S^{3}_{q,p}$ reads
${\rm d}s^{\prime 2}$ $=c_{11}{\rm d}\Theta^{2}+c_{22}{\rm d}\sigma^{2}+h{\left({{\rm d}y+A}\right)}^{2}~{},$
$c_{11}$ $=\frac{1}{4}{\left({q^{2}\cos^{2}\frac{\Theta}{2}+p^{2}\sin^{2}\frac{\Theta}{2}}\right)}=f^{2}/4~{},$
$c_{22}$ $=\frac{p^{2}q^{2}\sin^{2}\Theta}{4{\left({p^{2}t_{22}^{2}\cos^{2}\frac{\Theta}{2}+q^{2}t_{12}^{2}\sin^{2}\frac{\Theta}{2}}\right)}}~{},$
$h$ $={\left({p^{2}t_{22}^{2}\cos^{2}\frac{\Theta}{2}+q^{2}t_{12}^{2}\sin^{2}\frac{\Theta}{2}}\right)}~{},$
$A$ $=\frac{p^{2}t_{21}t_{22}\cos^{2}\frac{\Theta}{2}+q^{2}t_{11}t_{12}\sin^{2}\frac{\Theta}{2}}{p^{2}t_{22}^{2}\cos^{2}\frac{\Theta}{2}+q^{2}t_{12}^{2}\sin^{2}\frac{\Theta}{2}}{\rm d}\sigma~{}.$ (5.4)
Near $\Theta=0$ the metric becomes
${\rm d}s^{2}_{N}$ $=\frac{q^{2}}{4}{\left({{\rm d}\Theta^{2}+\frac{\Theta^{2}}{t_{22}^{2}}{\rm d}\sigma^{2}}\right)}+p^{2}t_{22}^{2}{\left({{\rm d}\varphi+\frac{t_{21}}{t_{22}}{\rm d}\sigma}\right)}^{2}~{},$ (5.5)
while at $\Theta=\pi$ the metric reads
${\rm d}s^{2}_{S}$ $=\frac{p^{2}}{4}{\left({{\rm d}\Theta^{2}+\frac{{\left({\pi-\Theta}\right)}^{2}}{t_{12}^{2}}{\rm d}\sigma^{2}}\right)}+q^{2}t_{12}^{2}{\left({{\rm d}y+\frac{t_{11}}{t_{12}}{\rm d}\sigma}\right)}^{2}~{},$ (5.6)
implying
$t_{12}=n_{S}~{},$ $t_{22}=n_{N}~{}.$ (5.7)
The orbifold Euler-characteristic of the base is
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}\Theta{\rm d}z\sqrt{g}R=\frac{1}{|t_{12}|}+\frac{1}{|t_{22}|}=\frac{1}{n_{S}}+\frac{1}{n_{N}}=\chi_{\rm orb}~{}.$ (5.8)
Moreover,
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}A=\frac{t_{11}}{t_{12}}-\frac{t_{21}}{t_{22}}=\frac{\det t}{t_{12}t_{22}}=\frac{1}{n_{N}n_{S}}~{},$ (5.9)
where we used
$t_{11}t_{22}-t_{21}t_{12}=n_{N}t_{S}-n_{S}t_{N}=1~{},$ (5.10)
with the relabelling
$t_{11}=t_{S}~{},$ $t_{21}=t_{N}~{},$ (5.11)
giving
$t=\begin{pmatrix}t_{S}&n_{S}\\ t_{N}&n_{N}\end{pmatrix}~{},$ $\det t=n_{N}t_{S}-n_{S}t_{N}=1~{}.$ (5.14)
In ${\left({\sigma,y}\right)}$ coordinates, the phase of the Killing spinor $\zeta$ becomes
$\exp{\left({\frac{{\rm i}}{2}\alpha_{2}\phi_{1}+\frac{{\rm i}}{2}\alpha_{3}\phi_{2}}\right)}$ $=\exp{\left[{\frac{{\rm i}}{2}{\left({t_{11}\alpha_{2}+t_{21}\alpha_{3}}\right)}\sigma+\frac{{\rm i}}{2}{\left({t_{12}\alpha_{2}+t_{22}\alpha_{3}}\right)}y}\right]}~{}.$ (5.15)
We introduce
$2c_{R}$ $=t_{11}\alpha_{2}+t_{21}\alpha_{3}~{},$
$2\lambda_{R}$ $=t_{12}\alpha_{2}+t_{22}\alpha_{3}~{},$ (5.16)
with inverse relations
$\alpha_{2}$ $=2{\left({t_{22}c_{R}-t_{21}\lambda_{R}}\right)}~{},$
$\alpha_{3}$ $=2{\left({-t_{12}c_{R}+t_{11}\lambda_{R}}\right)}~{}.$ (5.17)
Invariance under $y\sim y+2\pi$ yields
$t_{12}\alpha_{2}+t_{22}\alpha_{3}=2\lambda_{R}~{},$ $\lambda_{R}\in\mathbb{Z}~{}.$ (5.18)
The background field $A^{C}$ is smooth on $S^{3}_{q,p}$ iff $\alpha_{2}=-1$ and $\alpha_{3}=-\mathfrak{s}_{w}$, hence
$-n_{S}-\mathfrak{s}_{w}n_{N}=2\lambda_{R}~{},$ $\mathfrak{s}_{w}=\pm 1~{}.$ (5.19)
If $\gcd{\left({n_{N},n_{S}}\right)}=1$, then both $n_{N}$ and $n_{S}$ are odd. If we allow for spinors that are anti-periodic along $y$, then either $n_{N}$ or $n_{S}$ can be even. If we denote by $c_{R}$ the coefficient of $\sigma$ in the phase of the Killing spinor $\zeta$, thus
$2c_{R}$ $=t_{11}\alpha_{2}+t_{21}\alpha_{3}=-t_{S}-\mathfrak{s}_{w}t_{N}~{},$
$2\lambda_{R}$ $=t_{12}\alpha_{2}+t_{22}\alpha_{3}=-n_{S}-\mathfrak{s}_{w}n_{N}~{},$ (5.20)
with inverse relations
$\alpha_{2}$ $=2{\left({t_{22}c_{R}-t_{21}\lambda_{R}}\right)}=2{\left({n_{N}c_{R}-t_{N}\lambda_{R}}\right)}~{},$
$\alpha_{3}$ $=2{\left({-t_{12}c_{R}+t_{11}\lambda_{R}}\right)}=2{\left({-n_{S}c_{R}+t_{S}\lambda_{R}}\right)}~{}.$ (5.21)
In ${\left({\sigma,y}\right)}$ coordinates, the phase of the Fourier modes on $S^{3}_{q,p}$ reads
$\exp{\left({{\rm i}m_{i}\phi_{1}+{\rm i}n_{i}\phi_{2}}\right)}$ $=\exp{\left[{{\rm i}{\left({t_{11}m_{i}+t_{21}n_{i}}\right)}\sigma+{\rm i}{\left({t_{12}m_{i}+t_{22}n_{i}}\right)}y}\right]}=\exp{\left[{{\rm i}j_{i}\sigma+{\rm i}\mathfrak{m}_{i}y}\right]}~{},$ (5.22)
with
$j_{i}$ $=t_{11}m_{i}+t_{21}n_{i}=t_{S}m_{i}+t_{N}n_{i}\in\mathbb{Z}~{},$
$\mathfrak{m}_{i}$ $=t_{12}m_{i}+t_{22}n_{i}=n_{S}m_{i}+n_{N}n_{i}\in\mathbb{Z}~{},$ (5.23)
and inverse relations
$m_{i}$ $=n_{N}j_{i}-t_{N}\mathfrak{m}_{i}~{},$
$n_{i}$ $=-n_{S}j_{i}+t_{S}\mathfrak{m}_{i}~{}.$ (5.24)
The Killing vector in ${\left({\sigma,y}\right)}$ coordinates is
$K$ $=\mathcal{N}p{\left({\frac{\mathfrak{s}_{w}}{q}\partial_{\phi_{1}}+\frac{1}{p}\partial_{\phi_{2}}}\right)}~{},$
$=\mathcal{N}p{\left[{\frac{\mathfrak{s}_{w}}{q}{\left({t_{22}\partial_{\sigma}-t_{21}\partial_{y}}\right)}+\frac{1}{p}{\left({-t_{12}\partial_{\sigma}+t_{11}\partial_{y}}\right)}}\right]}~{},$
$=\mathcal{N}p{\left[{{\left({\frac{\mathfrak{s}_{w}t_{22}}{q}-\frac{t_{12}}{p}}\right)}\partial_{\sigma}+{\left({\frac{t_{11}}{p}-\frac{\mathfrak{s}_{w}t_{21}}{q}}\right)}\partial_{y}}\right]}~{},$
$=\mathcal{N}p{\left({\frac{t_{11}}{p}-\frac{\mathfrak{s}_{w}t_{21}}{q}}\right)}{\left({{\rm i}\omega_{\rm 2d}\partial_{\sigma}+\partial_{y}}\right)}~{},$ (5.25)
where we defined the angular-momentum fugacity on the base $\mathbb{\Sigma}$ as
$\alpha_{\rm 2d}={\rm i}\omega_{\rm 2d}=\frac{\mathfrak{s}_{w}t_{22}p-t_{12}q}{t_{11}q-\mathfrak{s}_{w}t_{21}p}=\frac{\mathfrak{s}_{w}n_{N}p-n_{S}q}{t_{S}q-\mathfrak{s}_{w}t_{N}p}~{}.$ (5.26)
In the new variables the one-loop determinant becomes
$Z_{\text{1-L}}$ $=\frac{\prod_{\ell_{2},k_{2}\in\mathbb{N}}{\left({p\ell_{2}+qk_{2}-{\rm i}\mathfrak{s}_{w}pq\sigma_{c}-\frac{q_{R}^{B}}{2}{\left({p+q}\right)}}\right)}}{\prod_{\ell_{1},k_{1}\in\mathbb{N}}{\left({p\ell_{1}+qk_{1}+{\rm i}\mathfrak{s}_{w}pq\sigma_{c}+\frac{q_{R}^{\phi}}{2}{\left({p+q}\right)}}\right)}}~{},$
$=\frac{\prod_{m_{2},n_{2}}{\left[{-pm_{2}-\mathfrak{s}_{w}qn_{2}-{\rm i}\mathfrak{s}_{w}pq\sigma_{c}+\frac{q_{R}^{B}}{2}{\left({p\alpha_{2}+q\mathfrak{s}_{w}\alpha_{3}}\right)}}\right]}}{\prod_{m_{1},n_{1}}{\left[{pm_{1}+\mathfrak{s}_{w}qn_{1}+{\rm i}\mathfrak{s}_{w}pq\sigma_{c}-\frac{q_{R}^{\phi}}{2}{\left({p\alpha_{2}+q\mathfrak{s}_{w}\alpha_{3}}\right)}}\right]}}~{},$
$=\frac{\prod_{j_{2}}{\left[{-p{\left({n_{N}j_{2}-t_{N}\mathfrak{m}_{2}}\right)}-\mathfrak{s}_{w}q{\left({-n_{S}j_{2}+t_{S}\mathfrak{m}_{2}}\right)}-{\rm i}\mathfrak{s}_{w}pq\sigma_{c}+\frac{q_{R}^{B}}{2}{\left({p\alpha_{2}+q\mathfrak{s}_{w}\alpha_{3}}\right)}}\right]}}{\prod_{m_{1},n_{1}}{\left[{p{\left({n_{N}j_{1}-t_{N}\mathfrak{m}_{1}}\right)}+\mathfrak{s}_{w}q{\left({-n_{S}j_{1}+t_{S}\mathfrak{m}_{1}}\right)}+{\rm i}\mathfrak{s}_{w}pq\sigma_{c}-\frac{q_{R}^{\phi}}{2}{\left({p\alpha_{2}+q\mathfrak{s}_{w}\alpha_{3}}\right)}}\right]}}~{},$
$=\frac{\prod_{j_{2}}-\mathfrak{s}_{w}{\left({\mathfrak{s}_{w}pn_{N}-qn_{S}}\right)}j_{2}+\mathfrak{s}_{w}{\left({\mathfrak{s}_{w}pt_{N}-qt_{S}}\right)}\mathfrak{m}_{2}-{\rm i}\mathfrak{s}_{w}pq\sigma_{c}+q_{R}^{B}{\left[{{\left({pn_{N}-q\mathfrak{s}_{w}n_{S}}\right)}c_{R}+{\left({q\mathfrak{s}_{w}t_{S}-pt_{N}}\right)}\lambda_{R}}\right]}}{\prod_{j_{1}}\mathfrak{s}_{w}{\left({\mathfrak{s}_{w}pn_{N}-qn_{S}}\right)}j_{1}-\mathfrak{s}_{w}{\left({\mathfrak{s}_{w}pt_{N}-qt_{S}}\right)}\mathfrak{m}_{1}+{\rm i}\mathfrak{s}_{w}pq\sigma_{c}-q_{R}^{\phi}{\left[{{\left({pn_{N}-q\mathfrak{s}_{w}n_{S}}\right)}c_{R}+{\left({q\mathfrak{s}_{w}t_{S}-pt_{N}}\right)}\lambda_{R}}\right]}}~{},$
$=\frac{\prod_{j_{2}}{\left[{-\alpha_{\rm 2d}{\left({j_{2}-q_{R}^{B}c_{R}}\right)}-{\left({\mathfrak{m}_{2}-q_{R}^{B}\lambda_{R}}\right)}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}{\prod_{j_{1}}{\left[{\alpha_{\rm 2d}{\left({j_{1}-q_{R}^{\phi}c_{R}}\right)}+{\left({\mathfrak{m}_{1}-q_{R}^{\phi}\lambda_{R}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}~{},$ (5.27)
where we used
$\mathfrak{s}_{w}n_{N}p-n_{S}q=\alpha_{\rm 2d}{\left({t_{S}q-\mathfrak{s}_{w}t_{N}p}\right)}~{}.$ (5.28)
Altogether,
$Z_{\text{1-L}}$ $=\frac{\prod_{j_{2}}{\left[{-\alpha_{\rm 2d}{\left({j_{2}-q_{R}^{B}c_{R}}\right)}-{\left({\mathfrak{m}_{2}-q_{R}^{B}\lambda_{R}}\right)}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}{\prod_{j_{1}}{\left[{\alpha_{\rm 2d}{\left({j_{1}-q_{R}^{\phi}c_{R}}\right)}+{\left({\mathfrak{m}_{1}-q_{R}^{\phi}\lambda_{R}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}~{},$
$m_{i}$ $=n_{N}j_{i}-t_{N}\mathfrak{m}_{i}~{},\qquad n_{i}=-n_{S}j_{i}+t_{S}\mathfrak{m}_{i}~{},$
$\lambda_{R}$ $=-\frac{\mathfrak{s}_{w}n_{N}+n_{S}}{2}~{},\qquad c_{R}=-\frac{t_{S}+\mathfrak{s}_{w}t_{N}}{2}=\frac{\mathfrak{s}_{w}n_{N}-n_{S}}{4\,n_{N}n_{S}}+\frac{\lambda_{R}{\left({n_{N}t_{S}+n_{S}t_{N}}\right)}}{2\,n_{N}n_{S}}~{},$
$m_{1}$ $=\ell_{1}~{},\qquad n_{1}=\mathfrak{s}_{w}k_{1}~{},\qquad m_{2}=-\ell_{2}~{},\qquad n_{2}=-\mathfrak{s}_{w}k_{2}~{},\qquad\ell_{1,2},k_{1,2}\in\mathbb{N}~{}.$ (5.29)
5.1.2 Twisted Ellipsoid
The ellipsoid with a non-trivial R-symmetry flux is obtained by setting
$\mathfrak{s}_{w}=1~{},$ $n_{N}=n_{S}=1~{},\qquad t_{S}=1+t_{N}~{},$ (5.30)
implying
$m_{i}$ $=j_{i}-t_{N}\mathfrak{m}_{i}~{},\qquad n_{i}=-j_{i}+t_{N}\mathfrak{m}_{i}+\mathfrak{m}_{i}~{},$
$\lambda_{R}$ $=-1~{},\qquad c_{R}=-\frac{2t_{N}+1}{2}~{},$
$m_{1}$ $=\ell_{1}~{},\qquad n_{1}=k_{1}~{},\qquad m_{2}=-\ell_{2}~{},\qquad n_{2}=-k_{2}~{},\qquad\ell_{1,2},k_{1,2}\in\mathbb{N}~{},$ (5.31)
which give the intervals
$t_{N}\mathfrak{m}_{1}\leq j_{1}\leq t_{S}\mathfrak{m}_{1}=t_{N}\mathfrak{m}_{1}+\mathfrak{m}_{1}~{},$ $t_{N}\mathfrak{m}_{2}+\mathfrak{m}_{2}=t_{S}\mathfrak{m}_{2}\leq j_{2}\leq t_{N}\mathfrak{m}_{2}~{}.$ (5.32)
The one-loop determinant can be put in the form reported in [36] via the shifts
$j_{1}=i_{1}+\frac{\mathfrak{m}_{1}}{2}{\left({t_{N}+t_{S}}\right)}=i_{1}+\mathfrak{m}_{1}{\left({t_{N}+\frac{1}{2}}\right)}~{},$
$j_{2}=i_{2}+\frac{\mathfrak{m}_{2}}{2}{\left({t_{N}+t_{S}}\right)}=i_{2}+\mathfrak{m}_{2}{\left({t_{N}+\frac{1}{2}}\right)}~{},$ (5.33)
with $i_{1,2}$ being half-integers satisfying
$-\frac{\mathfrak{m}_{1}}{2}\leq i_{1}\leq\frac{\mathfrak{m}_{1}}{2}~{},$ $\frac{\mathfrak{m}_{2}}{2}\leq i_{2}\leq-\frac{\mathfrak{m}_{2}}{2}~{}.$ (5.34)
The integers $\mathfrak{m}_{1,2}$ are the fluxes assigned to $\phi,B$ on the ellipsoid, respectively [33] . Hence,
$\mathfrak{m}_{1}$ $=q_{G}\mathfrak{m}+q_{R}^{\phi}\mathfrak{f}_{R}=\mathfrak{m}+q_{R}^{\phi}\lambda_{R}=\mathfrak{m}-r=-\mathfrak{r}~{},$
$\mathfrak{m}_{2}$ $=q_{G}\mathfrak{m}+q_{R}^{B}\mathfrak{f}_{R}=\mathfrak{m}+q_{R}^{B}\lambda_{R}=\mathfrak{m}+2-r=2-\mathfrak{r}~{},$ (5.35)
giving
$\frac{\mathfrak{r}}{2}\leq i_{1}\leq-\frac{\mathfrak{r}}{2}~{},$ $-\frac{\mathfrak{r}-2}{2}\leq i_{2}\leq\frac{\mathfrak{r}-2}{2}~{}.$ (5.36)
Finally,
$Z_{\text{1-L}}|_{\mathfrak{r}\geq 2}$ $=\prod_{i_{2}={\left({2-\mathfrak{r}}\right)}/2}^{{\left({\mathfrak{r}-2}\right)}/2}{\left({-\alpha_{\rm 2d}i_{2}-\gamma_{G}}\right)}={\left({-}\right)}^{\mathfrak{r}-1}\prod_{i_{2}={\left({2-\mathfrak{r}}\right)}/2}^{{\left({\mathfrak{r}-2}\right)}/2}{\left({\alpha_{\rm 2d}i_{2}+\gamma_{G}}\right)}~{},$
$Z_{\text{1-L}}|_{\mathfrak{r}=1}$ $=1~{},$
$Z_{\text{1-L}}|_{\mathfrak{r}\leq 0}$ $=\prod_{i_{1}=\mathfrak{r}/2}^{-\mathfrak{r}/2}{\left({\alpha_{\rm 2d}i_{2}+\gamma_{G}}\right)}^{-1}~{},$ (5.37)
matching [36] . The gauge fugacity reads
$\gamma_{G}=\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({t_{N}+t_{S}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}+\mathfrak{m}~{}.$ (5.38)
The argument of the product in $Z_{\text{1-L}}$ is independent of the R-charge; namely, the R-symmetry fugacity $\gamma_{R}$ is vanishing. This is compatible with the 3d constraint ${\left({2\gamma_{R}}\right)}\in\mathbb{Z}$ valid for a $\mathcal{N}=2$ theory on $S^{2}\times S^{1}$ with R-symmetry flux through $S^{2}$.
5.1.3 Untwisted Ellipsoid
The ellipsoid with a no R-symmetry flux is obtained by setting
$\mathfrak{s}_{w}=-1~{},$ $n_{N}=n_{S}=1~{},\qquad t_{S}=1+t_{N}~{},$ (5.39)
leading to
$m_{i}$ $=n_{N}j_{i}-t_{N}\mathfrak{m}_{i}~{},\qquad n_{i}=-n_{S}j_{i}+t_{S}\mathfrak{m}_{i}~{},$
$\lambda_{R}$ $=0~{},\qquad c_{R}=-\frac{1}{2}~{},$
$m_{1}$ $=\ell_{1}~{},\qquad n_{1}=-k_{1}~{},\qquad m_{2}=-\ell_{2}~{},\qquad n_{2}=k_{2}~{},\qquad\ell_{1,2},k_{1,2}\in\mathbb{N}~{}.$ (5.40)
The integers $\mathfrak{m}_{1,2}$ are the fluxes assigned to $\phi,B$ on the ellipsoid, respectively [33] . Hence,
$\mathfrak{m}_{1}$ $=q_{G}\mathfrak{m}=\mathfrak{m}=\mathfrak{m}_{2}~{}.$ (5.41)
We then obtain the intervals
$j_{1}\geq t_{N}\mathfrak{m}~{},$ $j_{1}\geq t_{S}\mathfrak{m}~{},$
$j_{2}\leq t_{N}\mathfrak{m}~{},$ $j_{2}\leq t_{S}\mathfrak{m}~{}.$ (5.42)
The ranges above would introduce $|\mathfrak{m}|$ in the one-loop determinant. We can avoid that by picking up ranges from dual poles, which is the prescription employed in the index theorem computation. Therefore, if
$j_{1}\geq t_{N}\mathfrak{m}~{},$ $j_{2}\leq t_{S}\mathfrak{m}~{},$ (5.43)
and
$j_{1}=t_{N}\mathfrak{m}+i=\frac{\mathfrak{m}}{2}{\left({t_{N}+t_{S}}\right)}-\frac{\mathfrak{m}}{2}+i~{},$
$j_{2}=t_{S}\mathfrak{m}-i=\frac{\mathfrak{m}}{2}{\left({t_{N}+t_{S}}\right)}+\frac{\mathfrak{m}}{2}-i~{},\qquad i\in\mathbb{N}~{},$ (5.44)
we find
$Z_{\text{1-L}}=\prod_{i\in\mathbb{N}}\frac{i+1-\frac{r}{2}-\frac{\mathfrak{m}}{2}+{\rm i}\mathfrak{u}}{i+\frac{r}{2}-\frac{\mathfrak{m}}{2}-{\rm i}\mathfrak{u}}~{},$ (5.45)
matching [33]. The gauge fugacity is
$\gamma_{G}=\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({t_{N}+t_{S}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}+\mathfrak{m}=-{\rm i}\alpha_{\rm 2d}\mathfrak{u}~{},$ (5.46)
and R-symmetry fugacity
$\gamma_{R}=\frac{1}{2}\alpha_{\rm 2d}=\alpha_{\rm 2d}\frac{\chi{\left({S^{2}}\right)}}{4}~{},$ (5.47)
which is compatible with the 3d constraint valid for an anti-twisted $\mathcal{N}=2$ theory on $\mathbb{\Sigma}\times S^{1}$.
5.1.4 Twisted Spindle
The twisted spindle is obtained by setting
$\mathfrak{s}_{w}=1~{},$ (5.48)
with $n_{N}$ and $n_{S}$ being positive integers fulfilling $\gcd{\left({n_{N},n_{S}}\right)}=1$. Therefore,
$m_{i}$ $=n_{N}j_{i}-t_{N}\mathfrak{m}_{i}~{},\qquad n_{i}=-n_{S}j_{i}+t_{S}\mathfrak{m}_{i}~{},$
$\lambda_{R}$ $=-\frac{n_{N}+n_{S}}{2}~{},\qquad c_{R}=-\frac{t_{S}+t_{N}}{2}=\frac{1}{4}{\left({\frac{1}{n_{S}}-\frac{1}{n_{N}}}\right)}+\frac{\lambda_{R}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}~{},$
$m_{1}$ $=\ell_{1}~{},\qquad n_{1}=k_{1}~{},\qquad m_{2}=-\ell_{2}~{},\qquad n_{2}=-k_{2}~{},\qquad\ell_{1,2},k_{1,2}\in\mathbb{N}~{}.$ (5.49)
These correspond to the intervals
${\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}\leq j_{1}\leq{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}~{},$ ${\left\lceil{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rceil}\leq j_{2}\leq{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}~{}.$ (5.50)
By generalizing [33], we have
$\mathfrak{m}_{1}$ $=q_{G}\mathfrak{m}+q_{R}^{\phi}\mathfrak{f}_{R}=\mathfrak{m}+q_{R}^{\phi}\lambda_{R}=\mathfrak{m}-\frac{r}{2}{\left({n_{N}+n_{S}}\right)}~{},$
$\mathfrak{m}_{2}$ $=q_{G}\mathfrak{m}+q_{R}^{B}\mathfrak{f}_{R}=\mathfrak{m}+q_{R}^{B}\lambda_{R}=\mathfrak{m}+\frac{2-r}{2}{\left({n_{N}+n_{S}}\right)}~{},$ (5.51)
giving the partition function
$Z_{\text{1-L}}|_{\mathfrak{m}_{2}\leq 0}=\prod_{j_{2}={\left\lceil{{t_{S}\mathfrak{m}_{2}}/{n_{S}}}\right\rceil}}^{{\left\lfloor{{t_{N}\mathfrak{m}_{2}}/{n_{N}}}\right\rfloor}}{\left[{-\alpha_{\rm 2d}{\left({j_{2}-q_{R}^{B}c_{R}}\right)}-\mathfrak{m}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}}\right]}~{},$
$Z_{\text{1-L}}|_{1\leq{\left({r/2}\right)}{\left({n_{N}+n_{S}}\right)}-\mathfrak{m}\leq n_{N}+n_{S}-1}=1~{},$
$Z_{\text{1-L}}|_{\mathfrak{m}_{1}\geq 0}=\prod_{j_{1}={\left\lceil{{t_{N}\mathfrak{m}_{1}}/{n_{N}}}\right\rceil}}^{{\left\lfloor{{t_{S}\mathfrak{m}_{1}}/{n_{S}}}\right\rfloor}}{\left[{\alpha_{\rm 2d}{\left({j_{1}-q_{R}^{\phi}c_{R}}\right)}+\mathfrak{m}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}}\right]}^{-1}~{}.$ (5.52)
The same result can be rewritten in another form by noticing that
$j_{i}$ $=t_{N}n_{i}+t_{S}m_{i}~{},$
$=\frac{b_{i}}{2n_{N}n_{S}}+\frac{\mathfrak{m}_{i}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}\in\mathbb{Z}~{},$ (5.53)
where $b_{i}\in\mathbb{Z}$ is such that
$b_{i}=n_{S}m_{i}-n_{N}n_{i}~{},$ $n_{S}m_{i}+n_{N}n_{i}=\mathfrak{m}_{i}~{},$
$-\mathfrak{m}_{1}\leq b_{1}\leq\mathfrak{m}_{1}~{},$ $\mathfrak{m}_{2}\leq b_{2}\leq-\mathfrak{m}_{2}~{},$ (5.54)
yielding
$Z_{\text{1-L}}|_{\mathfrak{m}_{2}\leq 0}=\prod_{b_{2}=\mathfrak{m}_{2}}^{-\mathfrak{m}_{2}}{\left\{{-\alpha_{\rm 2d}{\left[{\frac{b_{2}}{2n_{N}n_{S}}-q_{R}^{B}{\left({\frac{n_{N}-n_{S}}{4\,n_{N}n_{S}}}\right)}}\right]}-\gamma_{G}}\right\}}~{},$
$Z_{\text{1-L}}|_{1\leq{\left({r/2}\right)}{\left({n_{N}+n_{S}}\right)}-\mathfrak{m}\leq n_{N}+n_{S}-1}=1~{},$
$Z_{\text{1-L}}|_{\mathfrak{m}_{1}\geq 0}=\prod_{b_{1}=-\mathfrak{m}_{1}}^{\mathfrak{m}_{1}}{\left\{{\alpha_{\rm 2d}{\left[{\frac{b_{1}}{2n_{N}n_{S}}-q_{R}^{\phi}{\left({\frac{n_{N}-n_{S}}{4\,n_{N}n_{S}}}\right)}}\right]}+\gamma_{G}}\right\}}^{-1}~{},$ (5.55)
where the gauge fugacity is
$\gamma_{G}=\mathfrak{m}+\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}~{},$ (5.56)
while the R-symmetry fugacity reads
$\gamma_{R}=\alpha_{\rm 2d}{\left({\frac{n_{N}-n_{S}}{4\,n_{N}n_{S}}}\right)}=\frac{\alpha_{\rm 2d}}{4}\xi~{},$ $\xi=\frac{1}{n_{S}}-\frac{1}{n_{N}}~{},$ (5.57)
where the latter is consistent with the two-dimensional reduction of the three-dimensional constraint for a $\mathcal{N}=2$ theory on $\mathbb{\Sigma}\times S^{1}$ with R-symmetry twist on $\mathbb{\Sigma}$.
In fact, if we introduce $i_{2}$ by
$j_{2}=-i_{2}+{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}~{},$ $0\leq i_{2}\leq{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}-{\left\lceil{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rceil}~{},$ (5.58)
we have
$Z_{\text{1-L}}|_{\mathfrak{m}_{2}\leq 0}$ $=\prod_{j_{2}={\left\lceil{{t_{S}\mathfrak{m}_{2}}/{n_{S}}}\right\rceil}}^{{\left\lfloor{{t_{N}\mathfrak{m}_{2}}/{n_{N}}}\right\rfloor}}{\left[{-\alpha_{\rm 2d}{\left({j_{2}-q_{R}^{B}c_{R}}\right)}-\mathfrak{m}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}}\right]}~{},$
$=\prod_{i_{2}=0}^{{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}-{\left\lceil{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rceil}}{\left[{-\alpha_{\rm 2d}{\left({-i_{2}+{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}-q_{R}^{B}c_{R}}\right)}-\mathfrak{m}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}}\right]}~{},$
$=\prod_{i_{2}=0}^{{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}-{\left\lceil{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rceil}}{\left({\alpha_{\rm 2d}i_{2}+x_{2}}\right)}~{},$
with
$x_{2}$ $=\alpha_{\rm 2d}{\left({-{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}+q_{R}^{B}c_{R}}\right)}-\mathfrak{m}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}~{},$
$=\alpha_{\rm 2d}{\left\{{-{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}+q_{R}^{B}{\left[{\frac{1}{4}{\left({\frac{1}{n_{S}}-\frac{1}{n_{N}}}\right)}+\frac{\lambda_{R}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}}\right]}}\right\}}-\gamma_{G}+\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}~{},$
$=\alpha_{\rm 2d}{\left[{-{\left({\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}-\frac{\mu_{2,N}}{n_{N}}}\right)}+q_{R}^{B}\frac{\xi}{4}}\right]}-\gamma_{G}+\alpha_{\rm 2d}\frac{\mathfrak{m}_{2}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}~{},$
$=\alpha_{\rm 2d}{\left({\frac{\mathfrak{m}_{2}}{2n_{N}n_{S}}+\frac{\mu_{2,N}}{n_{N}}+q_{R}^{B}\frac{\xi}{4}}\right)}-\gamma_{G}~{},$ (5.60)
giving
$Z_{\text{1-L}}|_{\mathfrak{m}_{2}\leq 0}$ $=\prod_{i_{2}=0}^{{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}-{\left\lceil{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rceil}}{\left[{\alpha_{\rm 2d}{\left({i_{2}+\frac{\mathfrak{m}_{2}}{2n_{N}n_{S}}+\frac{\mu_{2,N}}{n_{N}}+q_{R}^{B}\frac{\xi}{4}}\right)}-\gamma_{G}}\right]}~{}.$ (5.61)
Analogously, if
$j_{1}=-i_{1}+{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}~{},$ $0\leq i_{1}\leq{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}-{\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}~{},$ (5.62)
then,
$Z_{\text{1-L}}|_{\mathfrak{m}_{1}\geq 0}$ $=\prod_{j_{1}={\left\lceil{{t_{N}\mathfrak{m}_{1}}/{n_{N}}}\right\rceil}}^{{\left\lfloor{{t_{S}\mathfrak{m}_{1}}/{n_{S}}}\right\rfloor}}{\left[{\alpha_{\rm 2d}{\left({j_{1}-q_{R}^{\phi}c_{R}}\right)}+\mathfrak{m}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}}\right]}^{-1}~{},$
$=\prod_{i_{1}=0}^{{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}-{\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}}{\left[{\alpha_{\rm 2d}{\left({-i_{1}+{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}-q_{R}^{\phi}c_{R}}\right)}+\mathfrak{m}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}}\right]}^{-1}~{},$
$=\prod_{i_{1}=0}^{{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}-{\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}}{\left[{-\alpha_{\rm 2d}i_{1}+x_{1}}\right]}^{-1}~{},$
with
$x_{1}$ $=\alpha_{\rm 2d}{\left({{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}-q_{R}^{\phi}c_{R}}\right)}+\mathfrak{m}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}~{},$
$=\alpha_{\rm 2d}{\left\{{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}-\frac{\mu_{1,S}}{n_{S}}-q_{R}^{\phi}{\left[{\frac{\xi}{4}+\frac{\lambda_{R}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}}\right]}}\right\}}+\gamma_{G}-\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}~{},$
$=\alpha_{\rm 2d}{\left\{{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}-\frac{\mu_{1,S}}{n_{S}}-q_{R}^{\phi}\frac{\xi}{4}}\right\}}+\gamma_{G}-\alpha_{\rm 2d}\frac{\mathfrak{m}_{1}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}~{},$
$=\alpha_{\rm 2d}{\left({\frac{\mathfrak{m}_{1}}{2n_{N}n_{S}}-\frac{\mu_{1,S}}{n_{S}}-q_{R}^{\phi}\frac{\xi}{4}}\right)}+\gamma_{G}~{},$ (5.64)
yielding, all in all,
$Z_{\text{1-L}}|_{\mathfrak{m}_{2}\leq 0}=\prod_{i_{2}=0}^{{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}-{\left\lceil{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rceil}}{\left[{\alpha_{\rm 2d}{\left({i_{2}+\frac{\mathfrak{m}_{2}}{2n_{N}n_{S}}+\frac{\mu_{2,N}}{n_{N}}+q_{R}^{B}\frac{\xi}{4}}\right)}-\gamma_{G}}\right]}~{},$
$Z_{\text{1-L}}|_{1\leq{\left({r/2}\right)}{\left({n_{N}+n_{S}}\right)}-\mathfrak{m}\leq n_{N}+n_{S}-1}=1~{},$
$Z_{\text{1-L}}|_{\mathfrak{m}_{1}\geq 0}=\prod_{i_{1}=0}^{{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}-{\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}}{\left[{\alpha_{\rm 2d}{\left({-i_{1}+\frac{\mathfrak{m}_{1}}{2n_{N}n_{S}}-\frac{\mu_{1,S}}{n_{S}}-q_{R}^{\phi}\frac{\xi}{4}}\right)}+\gamma_{G}}\right]}^{-1}~{},$ (5.65)
with
$\mu_{2,N}={\rm mod}{\left({t_{N}\mathfrak{m}_{2},n_{N}}\right)}~{},$ $\mu_{1,S}={\rm mod}{\left({t_{S}\mathfrak{m}_{1},n_{S}}\right)}~{},$ (5.66)
where ${\rm mod}{\left({n,d}\right)}$ is the remainder of the integer division $n/d$. The effective gauge fugacity is
$\gamma_{G}=\mathfrak{m}+\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-t_{N}p}~{},$ (5.67)
and the effective R-symmetry fugacity is
$\gamma_{R}=\frac{\alpha_{\rm 2d}}{4}\xi~{},$ $\xi=\frac{1}{n_{S}}-\frac{1}{n_{N}}=\frac{\alpha_{\rm 2d}}{4}{\left({\frac{1}{n_{S}}-\frac{1}{n_{N}}}\right)}~{},$ (5.68)
which is the two-dimensional version of the 3d constraint for a twisted $\mathcal{N}=2$ theory on $\mathbb{\Sigma}\times S^{1}$.
The partition function can be written in a form resembling (3.44) of [53]:
$Z_{\text{1-L}}|_{\mathfrak{m}_{2}\leq 0}=\prod_{i_{2}=0}^{{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}+{\left\lfloor{-\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rfloor}}{\left[{\alpha_{\rm 2d}{\left({i_{2}+\frac{\mathfrak{m}_{2}}{2n_{N}n_{S}}+\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}-{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}+q_{R}^{B}\frac{\xi}{4}}\right)}-\gamma_{G}}\right]}~{},$
$Z_{\text{1-L}}|_{1\leq{\left({r/2}\right)}{\left({n_{N}+n_{S}}\right)}-\mathfrak{m}\leq n_{N}+n_{S}-1}=1~{},$
$Z_{\text{1-L}}|_{\mathfrak{m}_{1}\geq 0}=\prod_{i_{1}=0}^{{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}+{\left\lfloor{-\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rfloor}}{\left[{\alpha_{\rm 2d}{\left({-i_{1}+\frac{\mathfrak{m}_{1}}{2n_{N}n_{S}}+{\left\lfloor{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rfloor}-\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}-q_{R}^{\phi}\frac{\xi}{4}}\right)}+\gamma_{G}}\right]}^{-1}~{},$ (5.69)
with the half-fluxes $\mathfrak{m}_{i}/{\left({n_{N}n_{S}}\right)}$ playing the role of the half-flux $s$, while $t_{N,S}\mathfrak{m}_{i}/n_{N,S}$ play that of the vortices contributions $\eta_{N,S}$.
5.1.5 Antitwisted Spindle
The antitwisted spindle is obtained by setting
$\mathfrak{s}_{w}=-1~{},$ (5.70)
with $n_{N}$ and $n_{S}$ being positive integers satisfying $\gcd{\left({n_{N},n_{S}}\right)}=1$. Thus,
$m_{i}$ $=n_{N}j_{i}-t_{N}\mathfrak{m}_{i}~{},\qquad n_{i}=-n_{S}j_{i}+t_{S}\mathfrak{m}_{i}~{},$
$\lambda_{R}$ $=\frac{n_{N}-n_{S}}{2}~{},\qquad c_{R}=\frac{-t_{S}+t_{N}}{2}=-\frac{n_{N}+n_{S}}{4\,n_{N}n_{S}}+\frac{\lambda_{R}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}~{},$
$m_{1}$ $=\ell_{1}~{},\qquad n_{1}=-k_{1}~{},\qquad m_{2}=-\ell_{2}~{},\qquad n_{2}=k_{2}~{},\qquad\ell_{1,2},k_{1,2}\in\mathbb{N}~{}.$ (5.71)
In particular,
$m_{1}=n_{N}j_{1}-t_{N}\mathfrak{m}_{1}\geq 0~{},$ $n_{1}=-n_{S}j_{1}+t_{S}\mathfrak{m}_{1}\leq 0~{},$
$m_{2}=n_{N}j_{2}-t_{N}\mathfrak{m}_{2}\leq 0~{},$ $n_{2}=-n_{S}j_{2}+t_{S}\mathfrak{m}_{2}\geq 0~{},$ (5.72)
that is
$j_{1}\geq{\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}~{},$ $j_{1}\geq{\left\lceil{\frac{t_{S}\mathfrak{m}_{1}}{n_{S}}}\right\rceil}~{},$
$j_{2}\leq{\left\lfloor{\frac{t_{N}\mathfrak{m}_{2}}{n_{N}}}\right\rfloor}~{},$ $j_{2}\leq{\left\lfloor{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rfloor}~{}.$ (5.73)
As in the case of the untwisted ellipsoid, we pick up ranges compatible with those emerging from the index computation, namely
$j_{1}\geq{\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}~{},$ $j_{2}\leq{\left\lfloor{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rfloor}~{}.$ (5.74)
Moreover, the prescription employed in [33] suggests that
$\mathfrak{m}_{1}$ $=n_{N}n_{S}{\left({q_{G}\mathfrak{f}_{G}+q_{R}^{\phi}\mathfrak{f}_{R}}\right)}=\mathfrak{m}+q_{R}^{\phi}\lambda_{R}=\mathfrak{m}+\frac{r}{2}{\left({n_{N}-n_{S}}\right)}~{},$
$\mathfrak{m}_{2}$ $=n_{N}n_{S}{\left({q_{G}\mathfrak{f}_{G}+q_{R}^{B}\mathfrak{f}_{R}}\right)}=\mathfrak{m}+q_{R}^{B}\lambda_{R}=\mathfrak{m}+\frac{r-2}{2}{\left({n_{N}-n_{S}}\right)}~{}.$ (5.75)
Eventually,
$Z_{\text{1-L}}$ $=\frac{\prod_{j_{2}={\left\lfloor{{t_{S}\mathfrak{m}_{2}}/{n_{S}}}\right\rfloor}}^{-\infty}{\left[{-\alpha_{\rm 2d}{\left({j_{2}-q_{R}^{B}c_{R}}\right)}-\mathfrak{m}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}{\prod_{j_{1}={\left\lceil{{t_{N}\mathfrak{m}_{1}}/{n_{N}}}\right\rceil}}^{+\infty}{\left[{\alpha_{\rm 2d}{\left({j_{1}-q_{R}^{\phi}c_{R}}\right)}+\mathfrak{m}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}\right]}}~{},$
$=\prod_{i\in\mathbb{N}}\frac{\alpha_{\rm 2d}{\left({i-{\left\lfloor{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rfloor}+q_{R}^{B}c_{R}}\right)}-\mathfrak{m}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}{\alpha_{\rm 2d}{\left({i+{\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}-q_{R}^{\phi}c_{R}}\right)}+\mathfrak{m}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}}~{}.$ (5.76)
We can rewrite the result by noticing that
$t_{N}\mathfrak{m}_{1}=n_{N}p_{1,N}+\mu_{1,N}~{},$ $p_{1,N}\in\mathbb{Z}~{},\qquad\mu_{1,N}\in{\left\{{0,\dots,n_{N}-1}\right\}}~{},$ (5.77)
giving
${\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}$ $=p_{1,N}+{\left\lceil{\frac{\mu_{1,N}}{n_{N}}}\right\rceil}=\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}-\frac{\mu_{1,N}}{n_{N}}+{\left\lceil{\frac{\mu_{1,N}}{n_{N}}}\right\rceil}~{},$ (5.78)
where ${\left\lceil{{\mu_{1,N}}/{n_{N}}}\right\rceil}=0$ if $\mu_{1,N}=0$ and ${\left\lceil{{\mu_{1,N}}/{n_{N}}}\right\rceil}=1$ if $\mu_{1,N}=1,\dots,n_{N}-1$. Furthermore,
$t_{S}\mathfrak{m}_{2}=n_{S}p_{2,S}+\mu_{2,S}~{},$ $p_{2,S}\in\mathbb{Z}~{},\qquad\mu_{2,S}\in{\left\{{0,\dots,n_{S}-1}\right\}}~{},$ (5.79)
implying
${\left\lfloor{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rfloor}$ $=p_{2,S}=\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}-\frac{\mu_{2,S}}{n_{S}}~{}.$ (5.80)
Hence, if
$Z_{\text{1-L}}$ $=\prod_{i\in\mathbb{N}}\frac{\alpha_{\rm 2d}i+x_{\rm num}}{\alpha_{\rm 2d}i+x_{\rm den}}~{},$ (5.81)
then,
$x_{\rm num}$ $=\alpha_{\rm 2d}{\left({-{\left\lfloor{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rfloor}+q_{R}^{B}c_{R}}\right)}-\mathfrak{m}-\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}~{},$
$=\alpha_{\rm 2d}{\left\{{-\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}+\frac{\mu_{2,S}}{n_{S}}+q_{R}^{B}{\left[{-\frac{\chi}{4}+\frac{\lambda_{R}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}}\right]}}\right\}}-\gamma_{G}+\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}~{},$
$=\alpha_{\rm 2d}{\left\{{-\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}+\frac{\mu_{2,S}}{n_{S}}-q_{R}^{B}\frac{\chi}{4}+\frac{\mathfrak{m}_{2}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}}\right\}}-\gamma_{G}~{},$
$=\alpha_{\rm 2d}{\left({-\frac{\mathfrak{m}_{2}}{2n_{N}n_{S}}+\frac{\mu_{2,S}}{n_{S}}-q_{R}^{B}\frac{\chi}{4}}\right)}-\gamma_{G}~{},$ (5.82)
as well as
$x_{\rm den}$ $=\alpha_{\rm 2d}{\left({{\left\lceil{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rceil}-q_{R}^{\phi}c_{R}}\right)}+\mathfrak{m}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q-\mathfrak{s}_{w}t_{N}p}~{},$
$=\alpha_{\rm 2d}{\left[{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}-\frac{\mu_{1,N}}{n_{N}}+{\left\lceil{\frac{\mu_{1,N}}{n_{N}}}\right\rceil}-q_{R}^{\phi}{\left({-\frac{\chi}{4}+\frac{\lambda_{R}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}}\right)}}\right]}+\gamma_{G}-\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}~{},$
$=\alpha_{\rm 2d}{\left[{\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}-\frac{\mu_{1,N}}{n_{N}}+{\left\lceil{\frac{\mu_{1,N}}{n_{N}}}\right\rceil}+q_{R}^{\phi}\frac{\chi}{4}-\frac{\mathfrak{m}_{1}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}}\right]}+\gamma_{G}~{},$
$=\alpha_{\rm 2d}{\left({q_{R}^{\phi}\frac{\chi}{4}-\frac{\mathfrak{m}_{1}}{2n_{N}n_{S}}-\frac{\mu_{1,N}}{n_{N}}+{\left\lceil{\frac{\mu_{1,N}}{n_{N}}}\right\rceil}}\right)}+\gamma_{G}~{}.$ (5.83)
Consequently,
$Z_{\text{1-L}}$ $=\prod_{i\in\mathbb{N}}\frac{\alpha_{\rm 2d}{\left({i-q_{R}^{B}\frac{\chi}{4}-\frac{\mathfrak{m}_{2}}{2n_{N}n_{S}}+\frac{\mu_{2,S}}{n_{S}}}\right)}-\gamma_{G}}{\alpha_{\rm 2d}{\left({i+q_{R}^{\phi}\frac{\chi}{4}-\frac{\mathfrak{m}_{1}}{2n_{N}n_{S}}-\frac{\mu_{1,N}}{n_{N}}+{\left\lceil{\frac{\mu_{1,N}}{n_{N}}}\right\rceil}}\right)}+\gamma_{G}}~{},$
$=\prod_{i\in\mathbb{N}}\frac{\alpha_{\rm 2d}{\left[{i-q_{R}^{B}\frac{\chi}{4}-\frac{\mathfrak{m}_{2}}{2n_{N}n_{S}}+\frac{{\rm mod}{\left({t_{S}\mathfrak{m}_{2},n_{S}}\right)}}{n_{S}}}\right]}-\gamma_{G}}{\alpha_{\rm 2d}{\left[{i+q_{R}^{\phi}\frac{\chi}{4}-\frac{\mathfrak{m}_{1}}{2n_{N}n_{S}}+\frac{{\rm mod}{\left({-t_{N}\mathfrak{m}_{1},n_{N}}\right)}}{n_{N}}}\right]}+\gamma_{G}}~{},$ (5.84)
where we used
${\left\lceil{\frac{x}{y}}\right\rceil}-\frac{x}{y}=\frac{{\rm mod}{\left({-x,y}\right)}}{y}~{},$ (5.85)
where ${\rm mod}{\left({n,d}\right)}$ is the remainder of the integer division $n/d$. The effective gauge fugacity is
$\gamma_{G}=\alpha_{\rm 2d}\frac{\mathfrak{m}}{2}{\left({\frac{t_{S}}{n_{S}}+\frac{t_{N}}{n_{N}}}\right)}+\frac{{\rm i}pq\sigma_{c}}{t_{S}q+t_{N}p}+\mathfrak{m}~{},$ (5.86)
while the effective R-symmetry fugacity reads
$\gamma_{R}=\alpha_{\rm 2d}\frac{\chi}{4}=\frac{\alpha_{\rm 2d}}{4}{\left({\frac{1}{n_{S}}+\frac{1}{n_{N}}}\right)}~{},$ (5.87)
which is the two-dimensional version of the 3d constraint for an antitwisted $\mathcal{N}=2$ theory on $\mathbb{\Sigma}\times S^{1}$.
The partition function can be written in a form resembling (3.44) of [53]:
$Z_{\text{1-L}}$ $=\prod_{i\in\mathbb{N}}\frac{\alpha_{\rm 2d}{\left({i-q_{R}^{B}\frac{\chi}{4}-\frac{\mathfrak{m}_{2}}{2n_{N}n_{S}}+\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}-{\left\lfloor{\frac{t_{S}\mathfrak{m}_{2}}{n_{S}}}\right\rfloor}}\right)}-\gamma_{G}}{\alpha_{\rm 2d}{\left({i+q_{R}^{\phi}\frac{\chi}{4}-\frac{\mathfrak{m}_{1}}{2n_{N}n_{S}}-\frac{t_{N}\mathfrak{m}_{1}}{n_{N}}-{\left\lfloor{\frac{-t_{N}\mathfrak{m}_{1}}{n_{N}}}\right\rfloor}}\right)}+\gamma_{G}}~{},$ (5.88)
where in the numerator the hemi-flux ${\mathfrak{m}_{2}}/{\left({2n_{N}n_{S}}\right)}$ corresponds to the hemi-flux $s$, while ${t_{S}\mathfrak{m}_{2}}/{n_{S}}$ appears just like the vortex contribution $\eta_{S}$. The same applies to the denominator, $S\to N$ provided.
6 Conclusions
In this paper we considered three-dimensional ${\cal N}=2$ supersymmetric field theories compactified on a large class of three-dimensional orbifolds. Our results include the calculation of the partition function of gauge theories on $\mathbb{\Sigma}\times S^{1}$ in the presence of both anti-twist and twist for the R-symmetry background gauge field, that were anticipated in [1]. We computed the one-loop determinants via the eigenvalue pairing method as well as by means of the equivariant orbifold index theorem. We also computed both the canonical expressions and the effective corrections to Chern-Simons theories and Fayet-Iliopoulos terms, which were crucial for testing our partition functions by means of non-perturbative dualities. Furthermore, we derived by supersymmetric localization the refined partition function on the branched lens space, which is a generalization of the usual lens space by the inclusion of conical singularities, thus generalizing the partition function on the so-called branched three-sphere. We also discussed a method for computing the one-loop determinants for two-dimensional quantum field theories on the spindle, again with both twist and anti-twist, extending an approach proposed in [33]. The derivation of this result exploited the interpretation of the squashed three-sphere as a smooth non-trivial circle orbi-bundle over a spindle.
[…]
On the field theory side, further investigation of supersymmetric dualities is interesting to gain a deeper comprehension of the quantum moduli space pertaining to SQFTs on orbifolds. For instance, in the study of gauge theories on manifolds, the phenomenon of partition functions factorizing into holomorphic blocks is well-known [58, 52, 44, 59]. Extending such results to gauge theories on orbifolds characterized by conical singularities, as those explored in this article, is an intriguing avenue of research. Another interesting outlook lies in generalizing the findings established for twisted gauge theories on non-compact manifolds [60, 61] to cover the case of non-compact orbifolds. Moreover, gauge theory partition functions on the two-sphere are known to provide the instanton corrections to the Kähler moduli space of Calabi-Yau manifolds, which directly correlate with Gromov-Witten invariants, as shown in [62]. Extending this understanding to incorporate two-dimensional gauge theories defined on spindles presents an intriguing area for exploration. Relatedly, it would be interesting to explore possible applications of our spindle index in the context of enumerative geometry, possibly generalizing the results of [63, 64].
It would be important to generalize our results to higher-dimensional models, including either $\mathcal{N}=1$ or $\mathcal{N}=2$ four-dimensional gauge theories on the direct product of a spindle and a Riemann surface, on $\mathbb{\Sigma}\times\mathbb{\Sigma}$, $\mathbb{WCP}^{2}_{{\left[{n_{1},n_{2},n_{3}}\right]}}$ (about which some results have been anticipated in [27]), or more general toric four-dimensional orbifolds. This could be done by exploiting the cohomological formalism developed e.g. in [65]. Orbifold partition functions of five-dimensional supersymmetric theories would allow to make contact with the supergravity solutions describing D4 branes wrapped on various orbifolds constructed in [66, 67, 68, 69, 70, 71, 72] and should also be relevant for six-dimensional SCFTs [73] compactified on orbifolds. Finallly, it is natural to wonder whether enlarging to the realm of orbifolds the equivariant localization program pursued in [74, 75, 76] would allow to uncover interesting aspects of (K-theoretic) Donaldson-Thomas and Gromov-Witten invariants associated to toric geometries.
Acknowledgements
The authors acknowledge partial support by the INFN. We thank L. Cassia, P. Ferrero, and F. Nieri for useful conversations. We also thank A. Zaffaroni for comments on a draft of this paper.
Appendix A Notation, Conventions and Identities
A.1 Spinors and Clifford algebra
We employ the spinor conventions of [77]: especially, indices of two-component spinors in the 2 representation of $SU(2)$ are raised and lowered by acting on the left with the anti-symmetric tensors $\varepsilon^{\alpha\beta}$ and $\varepsilon_{\alpha\beta}$, respectively, where $\varepsilon^{12}=\varepsilon_{21}=+1$. Given two spinors $\chi$ and $\eta$, their $SU(2)$-invariant contraction is defined as follows:
${\left({\chi\eta}\right)}=\chi^{\alpha}\eta_{\alpha}=-\chi_{\alpha}\eta^{\alpha}=\varepsilon^{\alpha\beta}\chi_{\beta}\eta_{\alpha}=\varepsilon_{\alpha\beta}\chi^{\alpha}\eta^{\beta}~{}.$ (A.1)
Let $\chi_{i}$, with $i=1,2,3$, be three commuting or anti-commuting spinors with components $\chi_{i,\alpha}$. Then, the Fierz identity
$\chi_{1,\alpha}(\chi_{2}\chi_{3})+(\chi_{1}\chi_{2})\chi_{3,\alpha}+\chi_{1,\beta}\chi_{2,\alpha}\chi_{3}^{\beta}=0~{}.$ (A.2)
holds for any $\alpha=1,2$. In the case of four commuting spinors, namely $\chi_{j}$ with $j=1,\dots,4$, the identity above implies
$(\chi_{1}\chi_{2})(\chi_{3}\chi_{4})=(\chi_{1}\chi_{3})(\chi_{2}\chi_{4})+(\chi_{1}\chi_{4})(\chi_{3}\chi_{2})~{}.$ (A.3)
The Euclidean gamma-matrices are proportional to the Pauli matrices $\sigma^{i}$ with $i=1,2,3$:
$\gamma_{1}=\sigma^{2}~{},\qquad\gamma_{2}=\sigma^{3}~{},\qquad\gamma_{3}=\sigma^{1}~{},$ $\gamma^{i}=\delta^{ij}\gamma_{j}~{}.$ (A.4)
We sometimes used the skew-symmetric product of two gamma matrices,
$\gamma_{ij}={\frac{1}{2}}{\left({\gamma_{i}\gamma_{j}-\gamma_{j}\gamma_{i}}\right)}={\rm i}\,\varepsilon_{ijk}\gamma^{k}~{},$ (A.5)
where the three-dimensional Levi-Civita tensor $\varepsilon_{ijk}$ satisfies
$\varepsilon^{ijm}\varepsilon_{k\ell m}={\delta^{i}}_{k}{\delta^{j}}_{\ell}-{\delta^{i}}_{\ell}{\delta^{j}}_{k}~{},\qquad\varepsilon^{i\ell m}\varepsilon_{k\ell m}=2{\delta^{i}}_{k}~{},$ $\varepsilon^{123}=\varepsilon_{123}=+1~{}.$ (A.6)
The following identities hold:
$\gamma_{i}\gamma_{j}$ $=\delta_{ij}+{\rm i}\,\varepsilon_{ijk}\gamma^{k}~{},$
$\gamma_{k}\gamma^{ki}$ $=\gamma^{ik}\gamma_{k}=2\,\gamma^{i}~{},$
${\gamma^{i}}_{k}\gamma^{kj}$ $=2\,\delta^{ij}+\gamma^{ij}~{},$ (A.7)
as well as the completeness relation
${{\left({\gamma^{j}}\right)}_{\alpha}}^{\beta}{{\left({\gamma_{j}}\right)}_{\gamma}}^{\delta}=2\,{\delta_{\alpha}}^{\delta}{\delta_{\gamma}}^{\beta}-{\delta_{\alpha}}^{\beta}{\delta_{\gamma}}^{\delta}~{},$ (A.8)
implying
$(\chi_{1}\gamma^{j}\chi_{2})(\chi_{3}\gamma_{j}\chi_{4})=2(\chi_{1}\chi_{4})(\chi_{3}\chi_{2})-(\chi_{1}\chi_{2})(\chi_{3}\chi_{4})~{},$ (A.9)
with $\chi_{1,2,3,4}$ being four commuting spinors.
A.2 Differential geometry
The spin connection $\omega_{\mu\,ab}$ used in the main text is defined by the relations
$\omega_{\mu\,ab}={e_{a}}^{\nu}\nabla_{\mu}e_{b\nu}=-e_{b\nu}\nabla_{\mu}{e_{a}}^{\nu}~{},$ ${\rm d}e_{a}+{\omega_{a}}^{b}\wedge e_{b}=0~{}.$ (A.10)
The spin connection allows for defining the covariant derivative of a spinor $\Psi$,
$\nabla_{\mu}\Psi=\partial_{\mu}\Psi+{\frac{1}{4}}\omega_{\mu\,ab}\gamma^{ab}\Psi=\partial_{\mu}\Psi+{\frac{{\rm i}}{4}}\omega_{\mu\,ab}\varepsilon^{abc}\gamma_{c}\Psi~{}.$ (A.11)
The Riemann tensor $R_{\mu\nu ab}$, the Ricci tensor $R_{\mu\nu}$ and the Ricci scalar $R$ read
$R_{\mu\nu ab}=\partial_{\mu}\omega_{\nu\,ab}-\partial_{\nu}\omega_{\mu\,ab}+{\omega_{\mu\,a}}^{c}\omega_{\nu\,cb}-{\omega_{\nu\,a}}^{c}\omega_{\mu\,cb}~{},\qquad R_{\mu\nu}={R^{\rho}}_{\mu\rho\nu}~{},\qquad R={R^{\mu}}_{\mu}~{}.$ (A.12)
The Lie derivatives along a vector $X^{\mu}$ of a scalar $\Phi$ and of a 1-form $A_{\mu}$ respectively are
$\mathcal{L}_{X}\Phi=X^{\nu}\partial_{\nu}\Phi~{},$ ${\left({\mathcal{L}_{X}A}\right)}_{\mu}=X^{\nu}\partial_{\nu}A_{\mu}+A_{\nu}\partial_{\mu}X^{\nu}~{},$ (A.13)
while the Lie derivative along a vector $X^{\mu}$ of a spinor $\Psi$ reads
$\mathcal{L}_{X}\Psi=X^{\mu}\nabla_{\mu}\Psi+{\frac{1}{4}}{\left({\nabla_{\mu}X_{\nu}}\right)}\gamma^{\mu\nu}\Psi=X^{\mu}\nabla_{\mu}\Psi+{\frac{{\rm i}}{4}}\varepsilon^{\mu\nu\lambda}{\left({\nabla_{\mu}X_{\nu}}\right)}\gamma_{\lambda}\Psi~{}.$ (A.14)
Appendix B Special functions
B.1 Zeta-functions and gamma-functions with multiple arguments
The multiple zeta-function is defined as
$\zeta_{N}(s,w|\vec{a}_{N})=\sum_{\vec{m}_{N}\in\mathbb{N}^{N}}{\frac{1}{{\left({w+\vec{a}_{N}\cdot\vec{m}_{N}}\right)}^{s}}}~{},$ ${\rm Re}(s)>N~{},$ (B.1)
which is related to the multiple gamma-function by
$\Gamma_{N}(w|\vec{a}_{N})=\exp{{\left[{\partial_{s}\zeta_{N}(s,w|\vec{a}_{N})|_{s=0}}\right]}}=\prod_{\vec{m}_{N}\in\mathbb{N}^{N}}{\frac{1}{w+\vec{a}_{N}\cdot\vec{m}_{N}}}~{}.$ (B.2)
If ${\rm Im}(\alpha_{i})>0$ with $i=1,\dots,N$, the identity
$\Gamma_{N+1}(w|1,\vec{a}_{N})\Gamma_{N+1}(1-w|1,-\vec{a}_{N})=e^{-{\rm i}\pi\zeta_{N+1}(0,w|1,\vec{\alpha}_{N})}\prod_{\vec{m}_{N}\in\mathbb{N}^{N}}{\frac{1}{1-e^{2\pi{\rm i}(w+\vec{a}_{N}\cdot\vec{m}_{N})}}}~{},$ (B.3)
holds. A special value of the multiple zeta-function employed in the main text is
$\zeta_{2}(0,w|a_{1},a_{2})={\frac{1}{12\,a_{1}a_{2}}}{\left[{6w^{2}-6(a_{1}+a_{2})w+a_{1}^{2}+a_{2}^{2}+3a_{1}a_{2}}\right]}~{}.$ (B.4)
Appendix C CKSE
Let us begin by recalling that a Killing spinor $\zeta$ is also a conformal Killing-spinor: if $\zeta$ fulfils the KSE with background fields $(A_{\mu},V_{\mu},H)$, then it exists a background field $A^{C}_{\mu}$ such that $\zeta$ satisfies the following CKSE:
${\left({\nabla_{\mu}-{\rm i}A^{C}_{\mu}}\right)}\zeta={\frac{1}{3}}\gamma_{\mu}\gamma^{\nu}{\left({\nabla_{\nu}-{\rm i}A^{C}_{\nu}}\right)}\zeta~{},$ (C.1)
where
$A^{C}_{\mu}=A_{\mu}-{\frac{3}{2}}V_{\mu}~{}.$ (C.2)
Indeed, by using
$\gamma_{\mu}\gamma_{\nu}=g_{\mu\nu}+{\rm i}\varepsilon_{\mu\nu\rho}\gamma^{\rho}~{},$ $\gamma^{\nu}{\left({\nabla_{\nu}-{\rm i}A_{\nu}}\right)}\zeta=-{\frac{3H}{2}}\zeta~{},$ (C.3)
we can rewrite the KSE for $\zeta$ as follows:
${\left({\nabla_{\mu}-{\rm i}A_{\mu}}\right)}\zeta=$ $-{\frac{H}{2}}\gamma_{\mu}\zeta-{\rm i}V_{\mu}\zeta-{\frac{1}{2}}\varepsilon_{\mu\nu\rho}V^{\nu}\gamma^{\rho}\zeta~{},$
$=$ ${\frac{1}{3}}\gamma_{\mu}\gamma^{\nu}{\left({\nabla_{\nu}-{\rm i}A_{\nu}}\right)}\zeta-{\rm i}V_{\mu}\zeta+{\frac{{\rm i}}{2}}V^{\nu}\gamma_{\mu}\gamma_{\nu}\zeta-{\frac{{\rm i}}{2}}V_{\mu}\zeta~{},$ (C.4)
namely
${\left({\nabla_{\mu}-{\rm i}A^{C}_{\mu}}\right)}\zeta={\frac{1}{3}}\gamma_{\mu}\gamma^{\nu}{\left({\nabla_{\nu}-{\rm i}A^{C}_{\nu}}\right)}\zeta~{},$ $A^{C}_{\mu}=A_{\mu}-{\frac{3}{2}}V_{\mu}~{}.$ (C.5)
The corresponding equation for ${\widetilde{\zeta}}$ is readily found by applying the maps $A_{\mu}\to-A_{\mu}$ and $V_{\mu}\to-V_{\mu}$, yielding $A^{C}_{\mu}\to-A^{C}_{\mu}$ and
${\left({\nabla_{\mu}+{\rm i}A^{C}_{\mu}}\right)}{\widetilde{\zeta}}={\frac{1}{3}}\gamma_{\mu}\gamma^{\nu}{\left({\nabla_{\nu}+{\rm i}A^{C}_{\nu}}\right)}{\widetilde{\zeta}}~{}.$ (C.6)
Now, let ${\mathcal{D}}_{\mu}$ be the $R$-symmetry covariant derivative at $V_{\mu}=0$, namely
${\mathcal{D}}_{\mu}=\nabla_{\mu}-{\rm i}q_{R}A_{\mu}~{}.$ (C.7)
Then, the Killing spinors $(\zeta,{\widetilde{\zeta}})$ of $R$-charges $(+1,-1)$ satisfy the following Killing-spinor equations (KSEs):
${\mathcal{D}}_{\mu}\zeta=(\nabla_{\mu}-{\rm i}A_{\mu})\zeta=-{\frac{H}{2}}\gamma_{\mu}\zeta-{\rm i}V_{\mu}\zeta-{\frac{1}{2}}\varepsilon_{\mu\nu\rho}V^{\nu}\gamma^{\rho}\zeta~{},$
${\mathcal{D}}_{\mu}{\widetilde{\zeta}}=(\nabla_{\mu}+{\rm i}A_{\mu}){\widetilde{\zeta}}=-{\frac{H}{2}}\gamma_{\mu}{\widetilde{\zeta}}+{\rm i}V_{\mu}{\widetilde{\zeta}}+{\frac{1}{2}}\varepsilon_{\mu\nu\rho}V^{\nu}\gamma^{\rho}{\widetilde{\zeta}}~{}.$
The latter imply the relations
$\gamma^{\mu}{\mathcal{D}}_{\mu}\zeta=\gamma^{\mu}{\left({\nabla_{\mu}-{\rm i}A_{\mu}}\right)}\zeta=-{\frac{3H}{2}}\zeta~{},$ $\gamma^{\mu}{\mathcal{D}}_{\mu}{\widetilde{\zeta}}=\gamma^{\mu}{\left({\nabla_{\mu}+{\rm i}A_{\mu}}\right)}{\widetilde{\zeta}}=-{\frac{3H}{2}}{\widetilde{\zeta}}~{},$ (C.8)
which are also useful to connect the KSE to the corresponding conformal Killing-spinor equations (CKSEs).
Appendix D Alternative representation of the supersymmetric background
The supersymmetry conditions can also be rewritten in terms of two R-symmetry neutral objects $(\xi,\Phi)$ defined as follows:
$\xi_{\mu}={v}^{-1}K_{\mu}~{},\qquad\Phi_{\mu\nu}=\varepsilon_{\mu\nu\rho}\xi^{\rho}={\rm i}P_{[\mu}\widetilde{P}_{\nu]}~{},$ (D.1)
satisfying
$\iota_{\xi}\xi=1~{},$ $g_{\mu\nu}=\xi_{\mu}\xi_{\nu}-{\Phi_{\mu}}^{\rho}\Phi_{\rho\nu}=\xi_{\mu}\xi_{\nu}-P_{(\mu}\widetilde{P}_{\nu)}~{},$ (D.2)
as well as
$\Phi_{\mu\nu}\xi^{\nu}=0~{},\qquad\Phi_{\mu\nu}P^{\nu}=-{\rm i}P^{\mu}~{},\qquad\Phi_{\mu\nu}\widetilde{P}^{\nu}=+{\rm i}\widetilde{P}^{\mu}~{}.$ (D.3)
When ${\widetilde{\zeta}}=\zeta^{c}$ coincides with the charge conjugate of the spinor $\zeta$, the one-form $\xi$ and the two-form $\Phi$ are both real and define an almost contact metric structure [34]. However, one can use these variables to write the background fields even if the background is not “real”, as in our set-up. The derivatives of the bilinears are
$\nabla_{\mu}\log{v}=\Phi_{\mu\rho}V^{\rho}~{},$ $\nabla_{\mu}\xi_{\nu}={\rm i}H\Phi_{\mu\nu}-\varepsilon_{\mu\nu\lambda}V^{\lambda}-\Phi_{\mu\rho}V^{\rho}\xi_{\nu}~{},$ (D.4)
implying
$V^{\lambda}$ $=\nabla_{\mu}\Phi^{\mu\lambda}+{\left({{\rm i}H-{\frac{1}{2}}\xi_{\rho}\nabla_{\mu}\Phi^{\mu\rho}}\right)}\xi^{\lambda}=\Phi^{\lambda\mu}\nabla_{\mu}\log{v}+{\left({{\rm i}H+{\frac{1}{2}}\xi_{\rho}\nabla_{\mu}\Phi^{\mu\rho}}\right)}\xi^{\lambda}~{},$ (D.5)
and
$A_{\mu}$ $={\frac{{\rm i}}{4}}\widetilde{P}^{\nu}\nabla_{\mu}P_{\nu}+{\frac{{\rm i}H}{2}}\xi_{\mu}+V_{\mu}~{}.$ (D.6)
The background field $A^{C}$ appearing in the conformal Killing spinor equation is independent of $H$ and reads
$A^{C}_{\mu}$ $=A_{\mu}-{\frac{3}{2}}V_{\mu}~{},$
$={\frac{{\rm i}}{4}}\widetilde{P}^{\nu}\nabla_{\mu}P_{\nu}+{\frac{{\rm i}H}{2}}\xi_{\mu}-{\frac{1}{2}}V_{\mu}~{},$
$={\frac{{\rm i}}{4}}\widetilde{P}^{\nu}\nabla_{\mu}P_{\nu}-{\frac{1}{2}}\nabla_{\rho}{\Phi^{\rho}}_{\mu}+{\frac{1}{4}}{\left({\xi_{\rho}\nabla_{\lambda}\Phi^{\lambda\rho}}\right)}\xi_{\mu}~{},$
$={\frac{{\rm i}}{4}}\widetilde{P}^{\nu}\nabla_{\mu}P_{\nu}-{\frac{1}{2}}{\Phi_{\mu}}^{\lambda}\nabla_{\lambda}\log{v}-{\frac{1}{4}}{\left({\xi_{\rho}\nabla_{\lambda}\Phi^{\lambda\rho}}\right)}\xi_{\mu}~{}.$ (D.7)
Appendix E Supersymmetry and cohomology
E.1 Vector multiplet
Supersymmetry transformations.
Given a gauge group $G$, a non-Abelian $\mathcal{N}=2$ vector multiplet in three dimensions contains a gauge field ${\mathcal{A}}_{\mu}$, a scalar field $\sigma$, the two-component spinors $\lambda_{\alpha}$, ${\widetilde{\lambda}}_{\alpha}$ and an auxiliary field $D$, all transforming in the adjoint representation of $G$. The fields $({\mathcal{A}}_{\mu},\sigma,\lambda,{\widetilde{\lambda}},D)$ respectively have $R$-charges $(0,0,1,-1,0)$. The corresponding supersymmetry transformations are
$\delta{\mathcal{A}}_{\mu}$ $=-{\rm i}\zeta\gamma_{\mu}{\widetilde{\lambda}}-{\rm i}{\widetilde{\zeta}}\gamma_{\mu}\lambda~{},$
$\delta\sigma$ $=-\zeta{\widetilde{\lambda}}+{\widetilde{\zeta}}\lambda~{},$
$\delta\lambda$ $=-{\frac{{\rm i}}{2}}\varepsilon^{\mu\nu\rho}\gamma_{\rho}\zeta{\mathcal{F}}_{\mu\nu}+{\rm i}\zeta(D+\sigma H)-\gamma^{\mu}\zeta({\rm i}{\rm D}_{\mu}\sigma-V_{\mu}\sigma)~{},$
$\delta{\widetilde{\lambda}}$ $=-{\frac{{\rm i}}{2}}\varepsilon^{\mu\nu\rho}\gamma_{\rho}{\widetilde{\zeta}}{\mathcal{F}}_{\mu\nu}-{\rm i}{\widetilde{\zeta}}(D+\sigma H)+\gamma^{\mu}{\widetilde{\zeta}}({\rm i}{\rm D}_{\mu}\sigma+V_{\mu}\sigma)~{},$
$\delta D$ $={\rm D}_{\mu}{\left({\zeta\gamma^{\mu}{\widetilde{\lambda}}-{\widetilde{\zeta}}\gamma^{\mu}\lambda}\right)}-{\rm i}V_{\mu}{\left({\zeta\gamma^{\mu}{\widetilde{\lambda}}+{\widetilde{\zeta}}\gamma^{\mu}\lambda}\right)}-{\left[{\sigma,\zeta{\widetilde{\lambda}}+{\widetilde{\zeta}}\lambda}\right]}-H{\left({\zeta{\widetilde{\lambda}}-{\widetilde{\zeta}}\lambda}\right)}~{},$ (E.1)
where ${\mathcal{F}}_{\mu\nu}=\partial_{\mu}{\mathcal{A}}_{\nu}-\partial_{\nu}{\mathcal{A}}_{\mu}-{\rm i}{\left[{{\mathcal{A}}_{\mu},{\mathcal{A}}_{\nu}}\right]}$, while the gauge-covariant derivative ${\rm D}_{\mu}$ acts on a field $\Psi$ of $R$-charge $q_{R}$, central charge $q_{Z}$ and in the representation ${\mathcal{R}}$ of the gauge group $G$ as follows:
${\rm D}_{\mu}\Psi={\left[{\nabla_{\mu}-{\rm i}q_{R}{\left({A_{\mu}-{\frac{1}{2}}V_{\mu}}\right)}-{\rm i}q_{Z}{\mathcal{C}}_{\mu}-{\rm i}{\mathcal{A}}_{\mu}\circ_{\mathcal{R}}}\right]}\Psi~{},$ (E.2)
with, for instance,
${\mathcal{A}}\circ_{\rm adjoint}\Psi={\left[{{\mathcal{A}},\Psi}\right]}~{},\qquad{\mathcal{A}}\circ_{\rm fundamental}\Psi={\mathcal{A}}\Psi~{},\qquad{\mathcal{A}}\circ_{\rm anti-fundamental}\Psi=-\Psi{\mathcal{A}}~{}.$ (E.3)
The fields belonging to the vector multiplet have vanishing central charge, hence ${\mathcal{C}}_{\mu}$ appears in (E.1) only implicitly, via $V_{\mu}$.
Cohomological complex.
The supersymmetry variation $\delta=\delta_{\zeta}+\delta_{\widetilde{\zeta}}$ is an equivariant differential. Indeed, the supersymmetry transformations for the vector multiplet can be rewritten as a cohomological complex containing the gauge field ${\mathcal{A}}$, the scalar $\sigma$, the Grassmann-odd 0-form $\chi={\rm i}({\widetilde{\zeta}}\lambda+\zeta{\widetilde{\lambda}})/(2{v})$ and their $\delta$-differentials:
$\delta{\mathcal{A}}$ $=\Lambda~{},$ $\delta\Lambda=-2{\rm i}{\left({L_{K}+{\mathcal{G}}_{\Phi_{G}}}\right)}{\mathcal{A}}~{},$
$\delta\sigma$ $=-({\rm i}/{v})\iota_{K}\Lambda~{},$ $\delta{\Phi_{G}}=0~{},$
$\delta\chi$ $=\Delta~{},$ $\delta\Delta=-2{\rm i}{\left({L_{K}+{\mathcal{G}}_{\Phi_{G}}}\right)}\chi~{},$ (E.4)
with $L_{K}$ being the covariant Lie derivative
$L_{K}$ $={\mathcal{L}}_{K}-{\rm i}q_{R}{\left({\iota_{K}A-{\frac{1}{2}}\iota_{K}V}\right)}-{\rm i}q_{Z}(\iota_{K}{\mathcal{C}})+{v}(q_{Z}-q_{R}H)~{},$
$={\mathcal{L}}_{K}-{\rm i}q_{R}\Phi_{R}-{\rm i}q_{Z}(\iota_{K}{\mathcal{C}}+{\rm i}{v})~{}.$ (E.5)
In fact, $L_{K}={\mathcal{L}}_{K}$ in (E.1) as all fields in (E.1) have vanishing central charge $q_{Z}$ and $R$-charge $q_{R}$. The cohomological complex of the vector multiplet is independent of the $U(1)_{R}$ bundle. In (E.1) also appears ${\mathcal{G}}_{\Phi_{G}}$, which is a gauge transformation with parameter ${\Phi_{G}}$, respectively acting on ${\mathcal{A}}$ and on a field $X\neq{\mathcal{A}}$ in the representation ${\mathcal{R}}$ of the gauge group as
${\mathcal{G}}_{\Phi_{G}}{\mathcal{A}}=-{\rm d}_{A}{\Phi_{G}}=-{\rm d}{\Phi_{G}}+{\rm i}{\left[{{\mathcal{A}},{\Phi_{G}}}\right]}~{},$ ${\mathcal{G}}_{\Phi_{G}}X=-{\rm i}{\Phi_{G}}\circ_{\mathcal{R}}X~{}.$ (E.6)
Explicitly, the Grassmann-odd 1-form $\Lambda_{\mu}$, the Grassmann-odd 0-form $\chi$ and the Grassmann-even 0-forms ${\Phi_{G}},\Delta$ are defined by
$\Lambda_{\mu}$ $=-{\rm i}{\left({\zeta\gamma_{\mu}{\widetilde{\lambda}}+{\widetilde{\zeta}}\gamma_{\mu}\lambda}\right)}~{},$
$\chi$ $={\frac{{\rm i}}{2{v}}}{\left({{\widetilde{\zeta}}\lambda+\zeta{\widetilde{\lambda}}}\right)}~{},$
${\Phi_{G}}$ $=\iota_{K}{\mathcal{A}}-{\rm i}{v}\sigma~{},$
$\Delta$ $=D+{\frac{1}{{v}}}\iota_{K}{\left({\star{\mathcal{F}}}\right)}+{\frac{{\rm i}}{{v}}}\sigma{\left[{\iota_{K}V-{\rm i}{v}H}\right]}=D+{\frac{{\rm i}}{2}}P^{\mu}\widetilde{P}^{\nu}{\mathcal{F}}_{\mu\nu}+{\frac{{\rm i}}{{v}}}\sigma{\left[{\iota_{K}V-{\rm i}{v}H}\right]}~{},$ (E.7)
with the map from $(\Lambda,\chi)$ to $(\lambda,{\widetilde{\lambda}})$ being
$\lambda={\rm i}{\left({{\frac{1}{2{v}^{2}}}\iota_{K}\Lambda+\chi}\right)}\zeta-{\frac{{\rm i}}{2{v}}}{\left({\iota_{P}\Lambda}\right)}{\widetilde{\zeta}}~{},$ ${\widetilde{\lambda}}={\rm i}{\left({{\frac{1}{2{v}^{2}}}\iota_{K}\Lambda-\chi}\right)}{\widetilde{\zeta}}-{\frac{{\rm i}}{2{v}}}{\left({\iota_{\widetilde{\rm P}}\Lambda}\right)}\zeta~{}.$ (E.8)
E.2 Chiral multiplet
Supersymmetry transformations.
A three-dimensional $\mathcal{N}=2$ chiral multiplet of $R$-charge $r$, central charge $z$ and in a representation ${\mathcal{R}}$ of a gauge group $G$ contains a complex scalar field $\phi$, a two-component spinor $\psi_{\alpha}$ and an auxiliary field $F$. The fields $(\phi,\psi,F)$ respectively have $R$-charges $(r,r-1,r-2)$ and central charge $z$. The corresponding supersymmetry transformations are
$\delta\phi$ $=\sqrt{2}\zeta\psi~{},$
$\delta\psi$ $=\sqrt{2}\zeta F+{\rm i}\sqrt{2}{\left({\sigma+rH-z}\right)}{\widetilde{\zeta}}\phi-{\rm i}\sqrt{2}\gamma^{\mu}{\widetilde{\zeta}}{\rm D}_{\mu}\phi~{},$
$\delta F$ $={\rm i}\sqrt{2}{\left[{z-\sigma-(r-2)H}\right]}{\widetilde{\zeta}}\psi-{\rm i}\sqrt{2}{\rm D}_{\mu}{\left({{\widetilde{\zeta}}\gamma^{\mu}\psi}\right)}+2{\rm i}({\widetilde{\zeta}}{\widetilde{\lambda}})\phi~{}.$ (E.9)
Analogously, a three-dimensional $\mathcal{N}=2$ anti-chiral multiplet of $R$-charge $-r$, central charge $-z$ and in the conjugate representation $\overline{\mathcal{R}}$ of a gauge group $G$ contains a complex scalar field ${\widetilde{\phi}}$, a two-component spinor ${\widetilde{\psi}}_{\alpha}$ and an auxiliary field ${\widetilde{F}}$. The fields $({\widetilde{\phi}},{\widetilde{\psi}},{\widetilde{F}})$ respectively have $R$-charges $(-r,-r+1,-r+2)$ and central charge $-z$. In this case, the supersymmetry transformations read
$\delta{\widetilde{\phi}}$ $=-\sqrt{2}{\widetilde{\zeta}}{\widetilde{\psi}}~{},$
$\delta{\widetilde{\psi}}$ $=\sqrt{2}{\widetilde{\zeta}}{\widetilde{F}}-{\rm i}\sqrt{2}{\widetilde{\phi}}{\left({\sigma+rH-z}\right)}\zeta+{\rm i}\sqrt{2}\gamma^{\mu}\zeta{\rm D}_{\mu}{\widetilde{\phi}}~{},$
$\delta{\widetilde{F}}$ $={\rm i}\sqrt{2}(\zeta{\widetilde{\psi}}){\left[{z-\sigma-(r-2)H}\right]}-{\rm i}\sqrt{2}{\rm D}_{\mu}{\left({\zeta\gamma^{\mu}{\widetilde{\psi}}}\right)}{\color[rgb]{1,0,0}+}2{\rm i}{\widetilde{\phi}}(\zeta\lambda)~{}.$ (E.10)
Cohomological complex.
Similarly to the case of the vector multiplet, the supersymmetry variation $\delta=\delta_{\zeta}+\delta_{\widetilde{\zeta}}$ acts on the fields of the chiral multiplet as an equivariant differential, which induces the following cohomological complex:
$\delta\phi=C~{},$ $\delta C=$ $-2{\rm i}{\left({L_{K}+{\mathcal{G}}_{\Phi_{G}}}\right)}\phi~{},$
$\delta B=\Theta~{},$ $\delta\Theta=$ $-2{\rm i}{\left({L_{K}+{\mathcal{G}}_{\Phi_{G}}}\right)}B~{},$ (E.11)
where the Grassmann-odd 0-forms $B,C$ and the Grassmann-even 0-form $\Theta$ are defined by
$B$ $=-{\frac{{\widetilde{\zeta}}\psi}{\sqrt{2}\,{v}}}~{},$
$C$ $=\sqrt{2}(\zeta\psi)~{},$
$\Theta$ $=F+{\rm i}{\left[{{\mathcal{L}}_{\widetilde{\rm P}}-{\rm i}r{\left({\iota_{\widetilde{\rm P}}A-{\frac{1}{2}}\iota_{\widetilde{\rm P}}V}\right)}-{\rm i}z(\iota_{\widetilde{\rm P}}{\mathcal{C}})-{\rm i}(\iota_{\widetilde{\rm P}}{\mathcal{A}})}\right]}\phi~{}.$ (E.12)
The inverse map from the Grassmann-odd scalars $(B,C)$ to the spinor $\psi$ is
$\psi=\sqrt{2}B\zeta+{\frac{C}{\sqrt{2}\,{v}}}{\widetilde{\zeta}}~{}.$ (E.13)
For completeness, we report the intermediate form of the supersymmetry transformations, the one obtained before introducing $\Theta$:
$\delta\phi=$ $C~{},$
$\delta C=$ $-2{\rm i}{\left[{{\mathcal{L}}_{K}-{\rm i}r{\left({\iota_{K}A-{\frac{1}{2}}\iota_{K}V}\right)}+{v}(z-rH)-{\rm i}z(\iota_{K}{\mathcal{C}})-{\rm i}{\Phi_{G}}}\right]}\phi~{},$
$\delta B=$ $F+{\rm i}{\left[{{\mathcal{L}}_{\widetilde{\rm P}}-{\rm i}r{\left({\iota_{\widetilde{\rm P}}A-{\frac{1}{2}}\iota_{\widetilde{\rm P}}V}\right)}-{\rm i}z(\iota_{\widetilde{\rm P}}{\mathcal{C}})-{\rm i}(\iota_{\widetilde{\rm P}}{\mathcal{A}})}\right]}\phi~{},$
$\delta F=$ $-2{\rm i}{\left[{{\mathcal{L}}_{K}-{\rm i}(r-2){\left({\iota_{K}A-{\frac{1}{2}}\iota_{K}V}\right)}-{\rm i}z\iota_{K}{\mathcal{C}}+{v}{\left[{z-{\left({r-2}\right)}H}\right]}-{\rm i}{\Phi_{G}}}\right]}B$
$-{\rm i}{\left[{{\mathcal{L}}_{\widetilde{\rm P}}-{\rm i}r{\left({\iota_{\widetilde{\rm P}}A-{\frac{1}{2}}\iota_{\widetilde{\rm P}}V}\right)}-{\rm i}z(\iota_{\widetilde{\rm P}}{\mathcal{C}})-{\rm i}(\iota_{\widetilde{\rm P}}{\mathcal{A}})}\right]}C-(\iota_{\widetilde{\rm P}}\Lambda)\phi~{}.$ (E.14)
Analogously, the cohomological complex of the anti-chiral multiplet reads
$\delta{\widetilde{\phi}}={\widetilde{C}}~{},$ $\delta{\widetilde{C}}=$ $-2{\rm i}{\left({L_{K}+{\mathcal{G}}_{\Phi_{G}}}\right)}{\widetilde{\phi}}~{},$
$\delta{\widetilde{B}}={\widetilde{\Theta}}~{},$ $\delta{\widetilde{\Theta}}=$ $-2{\rm i}{\left({L_{K}+{\mathcal{G}}_{\Phi_{G}}}\right)}{\widetilde{B}}~{},$ (E.15)
where the Grassmann-odd 0-forms ${\widetilde{B}},{\widetilde{C}}$ and the Grassmann-even 0-form ${\widetilde{\Theta}}$ are defined by
${\widetilde{B}}$ $={\frac{\zeta{\widetilde{\psi}}}{\sqrt{2}\,{v}}}~{},$
${\widetilde{C}}$ $=-\sqrt{2}({\widetilde{\zeta}}{\widetilde{\psi}})~{},$
${\widetilde{\Theta}}$ $={\widetilde{F}}+{\rm i}{\left[{{\mathcal{L}}_{P}+{\rm i}r{\left({\iota_{P}A-{\frac{1}{2}}\iota_{P}V}\right)}+{\rm i}z(\iota_{P}{\mathcal{C}})}\right]}{\widetilde{\phi}}-{\widetilde{\phi}}(\iota_{P}{\mathcal{A}})~{},$ (E.16)
with the inverse map from the Grassmann-odd scalars $({\widetilde{B}},{\widetilde{C}})$ to the spinor ${\widetilde{\psi}}$ being
${\widetilde{\psi}}=\sqrt{2}{\widetilde{B}}{\widetilde{\zeta}}+{\frac{{\widetilde{C}}}{\sqrt{2}\,{v}}}\zeta~{}.$ (E.17)
The intermediate step in the derivation of the cohomological complex of the anti-chiral multiplet is given by the transformations
$\delta{\widetilde{\phi}}=$ ${\widetilde{C}}~{},$
$\delta{\widetilde{C}}=$ $-2{\rm i}{\left[{{\mathcal{L}}_{K}+{\rm i}r{\left({\iota_{K}A-{\frac{1}{2}}\iota_{K}V}\right)}+{v}(-z+rH)+{\rm i}z(\iota_{K}{\mathcal{C}})}\right]}{\widetilde{\phi}}+2{\widetilde{\phi}}{\Phi_{G}}~{},$
$\delta{\widetilde{B}}=$ ${\widetilde{F}}+{\rm i}{\left[{{\mathcal{L}}_{P}+{\rm i}r{\left({\iota_{P}A-{\frac{1}{2}}\iota_{P}V}\right)}+{\rm i}z(\iota_{P}{\mathcal{C}})}\right]}{\widetilde{\phi}}-{\widetilde{\phi}}(\iota_{P}{\mathcal{A}})~{},$
$\delta{\widetilde{F}}=$ $-2{\rm i}{\left[{{\mathcal{L}}_{K}+{\rm i}(r-2){\left({\iota_{K}A-{\frac{1}{2}}\iota_{K}V}\right)}+{\rm i}z(\iota_{K}{\mathcal{C}})+{v}{\left[{-z-{\left({-r+2}\right)}H}\right]}}\right]}{\widetilde{B}}+2{\widetilde{B}}{\Phi_{G}}$
$-{\rm i}{\left[{{\mathcal{L}}_{P}+{\rm i}r{\left({\iota_{P}A-{\frac{1}{2}}\iota_{P}V}\right)}+{\rm i}z(\iota_{P}{\mathcal{C}})}\right]}{\widetilde{C}}+{\widetilde{C}}(\iota_{P}{\mathcal{A}})+{\widetilde{\phi}}(\iota_{P}\Lambda)~{}.$ (E.18)